### v1

# Sensor-fusion based Prognostics Framework for Complex Engineering Systems Exhibiting Multiple Failure Modes

Abstract
Complex engineering systems are often subject to multiple failure modes. Developing a remaining useful life (RUL) prediction model that does not consider the failure mode causing degradation is likely to result in inaccurate predictions. However, distinguishing between causes of failure without manually inspecting the system is nontrivial. This challenge is increased when the causes of historically observed failures are unknown. Sensors, which are useful for monitoring the state-of-health of systems, can also be used for distinguishing between multiple failure modes as the presence of multiple failure modes results in discriminatory behavior of the sensor signals. When systems are equipped with multiple sensors, some sensors may exhibit behavior correlated with degradation, while other sensors do not. Furthermore, which sensors exhibit this behavior may differ for each failure mode. In this paper, we present a simultaneous clustering and sensor selection approach for unlabeled training datasets of systems exhibiting multiple failure modes. The cluster assignments and the selected sensors are then utilized in real-time to first diagnose the active failure mode and then to predict the system RUL. We validate the complete pipeline of the methodology using a simulated dataset of systems exhibiting two failure modes and on a turbofan degradation dataset from NASA.
Index Terms:
Multiple failure modes, Sensor-fusion, prognostics, Sensor Selection, Adaptive Sparse Group Lasso, Mixture of Gaussian regressions, Functional Principal Component Analysis
publicationid: pubid: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
I Introduction
Prognostic models aim to estimate the remaining useful life (RUL) of critical components and systems. They are important because they drive condition-based maintenance decisions that aim to increase equipment availability, reduce downtime, and prevent unexpected failures. Generally, prognostic models can be classified into two categories: (1) data-driven models that focus on modeling sensor data and assume that the data is directly (or indirectly) correlated with the underlying physical degradation process, and (2) physics-based models that rely on first principles and require a thorough understanding of the underlying physics driving the degradation process. While physics-based approaches are known for their high accuracy, data-driven approaches are growing in popularity, especially for modeling complex systems. Several reasons are driving this trend. First, complex systems often exhibit complex physics of failure that are difficult to characterize using first principles. The second reason is the rise of Machine Learning and Artificial Intelligence models, which when trained properly can be very powerful prognostic tools. However, many data-driven prognostics models assume that complex systems exhibit a single predominant failure mode that can be used to accurately predict the system’s RUL. In practice, complex engineering systems often exhibit multiple failure modes and thus require more sophisticated modeling approaches to account for the different ways/modes a system can fail.
In this paper, we focus on developing a prognostic methodology for complex systems monitored by multiple sensors that exhibit multiple failure modes. We design an integrated framework comprised of (1) a sensor selection algorithm, (2) a dimensionality reduction and data fusion methodology, and (3) a RUL prediction model. Our methodology relies on a historical dataset to simultaneously diagnose observed failure events and select the subset of the most informative sensors associated with each failure mode. One of the unique aspects of our methodology relates to the sensor selection model, which uses an unsupervised approach to identify the dominant failure modes. Unlike most of the existing works, we assume that the labels of the failure events are unknown a priori. This assumption introduces some unique challenges around model estimation. These challenges are not trivial, especially when they are considered in conjunction with sensor selection.
Our methodology is comprised of two parts, an offline component and an online prognostic model. In the offline part, we develop a novel algorithm based on the Expectation-Maximization (EM) algorithm that simultaneously labels the failures in the historical training set and selects the optimal sensors for each failure mode. In the online part, we utilize real-time data to perform online diagnosis of the predominant failure mode and RUL prediction. During the online learning phase, Multivariate Functional Principle Component Analysis (MFPCA) is utilized for feature extraction and dimensionality reduction. We then use the nearest neighbor classifier to diagnose the active failure mode in the real-time signals before fitting a weighted functional regression model using the extracted features. The main contributions of this paper are threefold:
  • First, we propose a feature extraction methodology comprised of fusing signals from multiple sensors originating from a set of systems exhibiting multiple failure modes.
  • Second, we propose an optimization methodology for simultaneous failure mode labeling and sensor selection for any given unlabeled historical dataset. We call this as the “offline sensor selection” step.
  • Finally, we provide a two-step sequential approach to leverage this knowledge of failure labels and informative sensors to (1) perform active failure mode diagnosis and (2) subsequent RUL prediction on systems degrading in real-time. The paper mentions this as the “online diagnostics and prognostics” step.
The remainder of the paper is outlined as follows. In Section II, we review the relevant literature associated with this research area. In Section III, we outline our proposed multi-failure mode prognostics framework. In Section IV, we present our offline sensor selection scheme that utilizes an EM algorithm. In Section V, we present our online diagnostics and prognostics method. It discusses a strategy to fuse informative sensors for real-time diagnosis of failure modes and subsequent prediction of RUL. Our methodology is demonstrated through two case studies – one using a simulated dataset in section VI and the other NASA turbofan engine dataset in section VII. We conclude by summarizing our framework and key findings in Section VIII.
II Literature Review
Data-driven prognostic models utilize a variety of approaches. A significant number of these rely on Artificial Intelligence tools, such as neural networks to predict component lifetime [1, 2]. Some approaches have focused on directly modeling degradation-based signals that are derived from raw sensor data. Examples include random coefficient models [3], relevance vector machines [4], stochastic processes such as the Brownian motion process [5], the Gamma process [6], and hidden Markov models [7], and similarity-based prognostic models [8]. Most of these models have been developed for modeling a single sensor or degradation-based signal. Extending these approaches to applications involving multiple sensors is not straightforward.
Fusion algorithms have been leveraged to model multivariate sensor data. According to [9], fusion is usually performed at the decision level or the data level. Decision-based fusion involves combining the results of multiple prediction models in an ensemble fashion [10]. Conversely, data fusion methodologies focus on aggregating data from multiple sensors to develop a single (univariate) health index, which can then be modeled using the conventional methods discussed above. Most of the works in this area are characterized by their data fusion approach. Examples include principal component analysis (PCA) [11, 12, 13], maximum likelihood [14], deep learning [15], and state space models [16, 17, 18]. Most data fusion methodologies tend to aggregate all the sensor data. In practice, some sensors may not be correlated with the underlying failure process. Thus, integrating signals from these sensors into a data fusion algorithm introduces noise and may impact RUL prediction accuracy. To address this issue, several works have incorporated sensor selection algorithms. For example, [14] and [9] employed adaptive lasso to penalize non-informative sensors. [13] combined location-scale regression with group regularization technique to perform sensor selection. While these techniques can select an optimal subset of sensors for monitoring, they still assume that system failure is caused by a single predominant failure mode.
Most complex engineering systems can fail due to multiple failure modes. Competing risk models have been a popular approach for modeling multiple failure modes. Traditionally, competing risk models rely on statistical approaches. For example, [19] uses the Cox proportional hazard model (PHM) [20] with the Weibull baseline hazard rate to develop a mixture of Weibull PHM. For an extensive review of statistical competing risk models, we refer the reader to [21]. Deep learning has also been used to develop competing risk models. These models include convolutional neural network (CNN) [22], recurrent neural networks [23], and LSTM [24]. Recently, [25] developed a deep learning methodology to model complex survival problems, such as recurrent hospitalizations due to kidney failure or pancreas graft. For more examples, we refer the reader to [26], which provides a review of machine learning methodologies for survival analysis. Competing risk models often output survival/hazard rates that are sometimes difficult to interpret in engineering applications. Consequently, many researchers have adopted methods that explicitly model sensor data to predict RUL.
The literature that focuses explicitly on RUL prediction of systems with multiple failure models can be loosely divided into two groups. The first group uses deep learning approaches while the second group utilizes more statistical methods. Deep learning models such as [27] and [28] develop feature extraction methodologies based on convolutional neural networks (CNN) for modeling systems with multiple failure modes. [29] developed a Gaussian model whose mean is a Bayesian deep neural network that maps sensor data to the RUL. In [30], the authors used a support vector machine to diagnose the active failure mode using features extracted from a deep belief network. Following diagnosis, a particle filter is employed to predict RUL. [31] trained an LSTM-based RUL prediction model for all present failure modes. The choice of prediction model is based on a deep CNN-based classifier trained with physics-informed failure labels. While deep learning approaches have been successful at predicting RUL, their parameters are often difficult to interpret. Additionally, they require significantly larger datasets for training compared to statistical methods. In [32], the authors derived a composite health index for each failure mode and used a random coefficients-based polynomial degradation model to predict the system’s RUL. In their model training/estimation, the authors assumed that the labels of the failure modes were given. In fact, this assumption has been widely used in the literature and is often supported by visual inspection of the data [27, 30, 32] or available domain knowledge [31]. [33] assumes that a subset of the data is labeled and utilizes a similarity-based approach for classifying the rest of the unlabeled failure modes in the training set.
Our methodology differs from previous works in that we consider a scenario where the failure labels are completely unknown. Furthermore, we assume that for each failure mode, there exists a (not necessarily disjoint) subset of informative sensors that are useful for prognostics. Our objective is to design an unsupervised prognostic methodology that synthesizes signals from multiple sensors in a systematic fashion. We assume the availability of a historical repository of system failures whose degradation was monitored from the moment an incipient fault occurred up to system failure using multiple sensors. The methodology assumes that all systems fail independently of each other. In addition, we assume that only one type of failure mode can occur in the entire lifetime of the system. While the failure labels are unknown, the total number of potential failure modes for the system is assumed to be known.
III Overview of the Prognostics Framework
This section describes our multiple failure mode prognostics framework, which is summarized in the flowchart in Figure 1. This framework is comprised of two parts:
Offline Sensor Selection: The first part is performed offline and primarily focuses on utilizing a training set of historically observed degradation signals to perform sensor selection. Sensor selection is important because not all sensors are informative for prognostics, and the inclusion of non-informative sensors in a prognostics algorithm can compromise the accuracy of RUL predictions. When performing sensor selection, it is computationally convenient to represent each sensor by a parsimonious set of features that capture the system dynamics. Therefore we further divide this step into the following substeps:
  1. 1.
    Feature Extraction: It involves a variant of the traditional Functional Principal Component Analysis (FPCA) framework to account for multiple failure modes in the unlabeled data.
  2. 2.
    Optimization: This involves formulating an optimization model to simultaneously label the failure modes and select the informative sensors for each failure mode.
Feature Extraction: There are many ways to extract parsimonious features from complex and multivariate data. In our framework, we choose to use a Functional Principal Component Analysis (FPCA) framework due to its flexibility and extensibility to various types of functional data, allowing us to efficiently capture the underlying patterns and variability in complex, high-dimensional datasets. FPCA models the variability of temporal data observations around a mean function. However, the presence of multiple failure modes can cause heterogeneity in the magnitude and/or sign of the signals’ rate of change, compromising the effectiveness of FPCA as a prognostic tool. Therefore, the first step of our offline sensor selection algorithm is to extract features using a Covariate-Adjusted FPCA (CA-FPCA) approach [34].
CA-FPCA is an extension of conventional FPCA that allows modeling functional data while accounting for the effects of explanatory variables, known as covariates. In our context, the covariates are used to represent the failure modes. One of the unique contributions of our approach is the assumption that the failure modes are not known a priori. However, CA-FPCA assumes that covariates are observable. We note that the way each sensor responds to different failure modes can be highly variable. For example, a sensor may be more sensitive to one failure mode than another resulting in heterogeneity in the rate of change of its corresponding sensor signals. Furthermore, a sensor signal may increase in magnitude in response to one failure mode, and decrease for another. To address the lack of observable covariates, we first perform regular FPCA on the signals from each sensor. Then, we perform K-means clustering on the FPC scores for each sensor individually, using the resulting cluster labels as the covariate values to perform CA-FPCA.
Optimization: After feature extraction the sensor signals are transformed into parsimonious CA-FPC-scores, which serve as predictors for a Mixture of Gaussian (MGR) model used to model the natural logarithm of the time-to-failure, $\ln{TTF}$, as a function of the CA-FPC-scores. This model is motivated by the assumption that the functional relationship between the $\ln{TTF}$ and the CA-FPC scores is dependent on the latent failure mode. We fit the parameters of the MGR model using the Expectation-Maximization (EM) algorithm. In the E-step, we obtain a soft clustering of the ((((CA-FPC-scores, $\ln{TTF})$ tuples into their respective failure modes. In the M-step, we incorporate the Adaptive Sparse Group Lasso (ASGL) penalty into the regression model to remove groups of features associated with non-informative sensors. From now on, we refer to this model as MGR-ASGL. The output of the offline sensor sensor selection step of our methodology are (1) failure mode labels for each sample in the training set and (2) a subset of informative sensors for each failure mode.
Online diagnosis and Prognostics: The second part of our methodology consists of an online step used to process and analyze data from units operating in the field. The online step consists of diagnosing the active failure mode of the fielded unit and then predicting the RUL of the fielded unit. In the diagnostic step, features are extracted from the real-time data generated by all selected sensors using Multivariate FPCA (MFPCA), which is an extension of FPCA to multivariate functional data. We use MFPCA to generate a parsimonious set of fused signal features from multivariate sensor signals; hereafter referred to as MFPC-scores. Then, we classify the active failure mode of the fielded unit by assigning to the unit the failure mode most frequently represented by its K-nearest neighbors. Once this step is complete, we recalculate the MFPC scores using only the informative sensors associated with the diagnosed failure mode of the unit. In the prognostic step, we fit a functional regression model using the MFPC scores that predict the RUL of the unit. In the following sections, we formalize our prognostics framework.
Figure 1: Flowchart of the multiple failure mode prognostics framework
IV Offline Sensor Selection Algorithm
We consider a scenario where a system is monitored by $P$ sensors. We assume there exists a historical dataset comprised of $P$-dimensional degradation-based sensor signals from monitoring $N$ similar systems. We denote $s_{i,p}(t)$ as the signal data from sensor $p\in\{1,...,P\}$ for system $i\in\{1,...,N\}$ observed at time $t\in[0,T]$, where $T=min(\{TTF_{i}\}_{i=1}^{N})$. By truncating all signals to the minimum failure time, we ensure that all signals in the training set have the same time domain and that they are all included when modeling. For this framework, we assume that the total number of failure modes, $K$ is known, but that the actual failure modes in the training dataset are unknown. This poses a challenge for CA-FPCA, which requires observed covariates for computation. To determine these covariates, we perform a sensor-wise clustering of the training samples. For each sensor, we perform FPCA for feature extraction and then use K-means clustering with $K$ clusters to label the signals from that sensor. Once the sensor-wise clustering is complete, the next step involves feature extraction by applying CA-FPCA and then regressing these features against the time-to-failure to identify the most informative subset of sensors for each failure mode. The next two subsections discuss details about CA-FPCA and the penalized Mixture of Gaussian regression model (MGR-ASGL).
IV-A Feature Extraction: CA-FPCA
CA-FPCA extends conventional FPCA to handle complex data structures where the dynamics of time-varying signals are affected by one or more additional factors, i.e. covariates. In [34], the covariates are assumed to be continuous variables, and thus, mean and covariance estimation requires smoothing over both time and the covariate space. In this paper, we use the sensor-wise cluster labels as covariates. Let $\mathcal{I}_{p}^{k}$ denote the set of indices for sensor $p$ signals assigned to cluster $k,k\in\{1,2,...,K\}$. The CA-FPCA problem was reduced to performing FPCA on each cluster. For a given system $i$, signals from cluster $k$ of sensor $p$ are modeled as:
$s_{i,p}(t)=v_{i,p}^{k}(t)+\epsilon_{i,p}(t),i\in\mathcal{I}_{k}^{p}\vspace*{-1em}$ (1)
where $v_{i,p}^{k}(t)$ is a smooth random function, and $\epsilon_{i,p}(t)$ are assumed to be independently and identically distributed (i.i.d.) errors with mean zero and variance $\varsigma_{p}^{2}$. $v_{i,p}^{k}(t)$ and $\epsilon_{i,p}(t)$ are assumed to be independent of each other. The mean and covariance functions of $v_{i,p}^{k}(t)$ are given by $\mu_{p}^{k}(t)$ and $\mathbb{C}_{p}^{k}(t,t^{\prime})$, respectively. Using Mercer’s theorem [35], the covariance can be decomposed as follows, $\mathbb{C}_{p}^{k}(t,t^{\prime})=\sum_{m=1}^{\infty}\lambda_{m,p}^{k}\phi_{m,p}^{k}(t)\phi_{m,p}^{k}(t^{\prime})$, where $\{\phi_{m,p}^{k}(t)\}_{m=1}^{\infty}$ are the orthogonal eigenfunctions of $\mathbb{C}_{p}^{k}(t,t^{\prime})$ and $\lambda_{1,p}^{k}\geq...\geq\lambda_{m,p}^{k}\geq...$ are the corresponding eigenvalues. By projecting the mean-subtracted signal data onto the eigenfunctions, we can represent the signals by equation (2).
$s_{i,p}(t)=\mu_{p}^{k}(t)+\sum_{m=1}^{\infty}\xi_{i,m,p}^{k}\phi_{m,p}^{k}(t)+\epsilon_{i,p}(t)\vspace*{-1em}$ (2)
where $\xi_{i,m,p}^{k}$ is the $m$th CA-FPC-score with mean 0 and variance $\lambda_{m,p}^{k}$. The CA-FPC scores are computed using the PACE algorithm in [36]. Since the eigenvalues decrease as $m\to\infty$, we can obtain a low-dimensional representation of $s_{i,p}^{k}(t)$ using the first $q_{p}$ CA-FPC-scores.
To ensure all features are scaled equally, we use the standardized versions of the CA-FPC-scores: $x_{i,m}^{p}=\frac{\sum_{k=1}^{K}\xi_{i,m,p}^{k}\mathit{\mathbb{I}}(i\in\mathcal{I}_{p}^{k})}{\sqrt{\frac{1}{N-1}\sum_{k=1}^{r_{p}^{*}}N_{p}^{k}\lambda_{m,p}^{k}}}$, where $N_{p}^{k}$ is the number of signals assigned to cluster $k$ for sensor $p$. We extract $\boldsymbol{x}_{i,p}=(x_{i,1}^{p},...,x_{i,q_{p}}^{p})^{T}$ for sensor $p$ of system $i$,. We then utilize this $\ boldsymbol{x}_{i,p}$’s as predictors for fitting the MGR-ASGL model.
IV-B Optimization of the MGR-ASGL model
To perform sensor selection, we seek to model the distribution of $\ln{TTF}$ conditioned on the observed set of features. In what follows, we note that values of $\ln{TTF}$ are standardized. Our modeling approach assumes that the failure times are conditionally independent given the observed features. Let $y_{i}$ denote the random variable corresponding to the $\ln{TTF}$ of the system $i$. We consider $\boldsymbol{Z}_{i}$ to denote a “one-hot” random vector of length $K$, where $\boldsymbol{Z}_{i}[k]=1$ if system $i$ failed due to failure mode $k$, and zero otherwise. For mathematical brevity, we let $\boldsymbol{Y}=\left(y_{1},\ldots,y_{N}\right)$ denote the observed $\ln{TTF}$ and $\mathbb{Z}=\left(\boldsymbol{Z}_{1},\boldsymbol{Z}_{2},\ldots,\boldsymbol{Z}_{N}\right)$ denote the set of random vectors representing the failure modes for each system. In the following subsections, we first assign probability distributions to these random variables to formalize the MGR model. Second, we construct the log-likelihood function for the MGR model. Third, we provide theoretical intuition on the use of the EM algorithm for model fitting. Fourth, we describe the ASGL penalty. Finally, we provide details for implementing the EM algorithm.
IV-B1 Mixture of Gaussian Regression (MGR) model
We assume that the distribution of $y_{i}$ conditioned on the observed features $\boldsymbol{x}_{i,1},\ldots,\boldsymbol{x}_{i,P}$} and $\boldsymbol{Z}_{i}[k]=1$ is Gaussian. If we let $\pi_{1},\ldots,\pi_{K}$, where $\sum_{k=1}^{K}\pi_{k}=1$, denote the mixing coefficients, which represent the probability that $\boldsymbol{Z}_{i}[k]=1$ ($\forall k=\{1,\ldots,K\}$), then equation (3) gives the marginal distribution of $y_{i}$ given only $\{\boldsymbol{x}_{i,1},\ldots,\boldsymbol{x}_{i,p}\}$. This marginal distribution is the MGR model.
$\begin{split}f_{\vartheta}(y_{i}\mid\boldsymbol{x}_{i,1},&\ldots,\boldsymbol{x}_{i,P})=\sum_{k=1}^{K}\pi_{k}\frac{1}{\sqrt{2\pi}\sigma_{k}}\times\\ &\ \exp\left(-\frac{1}{2\sigma_{k}^{2}}\left(y_{i}-\beta_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\beta}_{p,k}\right)^{2}\right)\end{split}$ (3)
IV-B2 Log-Likelihood
[…]
Definition 1.
The log-likelihood function for the MGR model is given by equation (4). This function is called as the ”incomplete-data log-likelihood” (IDLL), since we do not know the cause of the observed failure in the data.
$\begin{split}\ell(\boldsymbol{\Theta};&\boldsymbol{Y})=\sum_{i=1}^{N}\ln\sum_{k=1}^{K}\pi_{k}\times\\ &\ \frac{\rho_{k}}{\sqrt{2\pi}}\exp\left(-\frac{1}{2}\left(\rho_{k}y_{i}-\psi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\psi}_{p,k}\right)^{2}\right)\end{split}\vspace*{-1em}$ (4)
Definition 2.
The “complete-data log-likelihood” (CDLL) given by equation (5) assumes knowledge of the underlying causes of the observed failure in the data.
$\begin{split}\ell_{C}(\boldsymbol{\Theta};&\boldsymbol{Y},\mathbb{Z})=\sum_{i=1}^{N}\sum_{k=1}^{K}\boldsymbol{Z}_{i}[k]\ln\pi_{k}\left(\frac{\rho_{k}}{\sqrt{2\pi}}\right)\\ &\exp\left(-\frac{1}{2}\left(\rho_{k}y_{i}-\psi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\psi}_{p,k}\right)^{2}\right)\end{split}\vspace*{-1em}$ (5)
The IDLL given in equation (4) is difficult to optimize globally [38]. Instead, we search for a local optimum by using EM algorithm, an iterative algorithm suitable for fitting probability models with latent variables. To derive the steps of the EM algorithm, we note that since $\ln(.)$ is a concave function We use Jensen’s inequality to form a lower bound on the IDLL. This lower bound is the expectation of the CDLL given in equation (5) and shown in Lemma 1.
Input : Number of systems $N$, Number of sensors $P$, Unlabeled data from $P$ sensors of all $N$ systems, Maximum iterations $iter_{max}$ for EM algorithm, threshold $\delta$ for regression coefficient’s $l_{2}$-norm
Output : Vector $\boldsymbol{Z}$ indicating failure labels, Set of informative sensors $\mathcal{P}_{k}$ for each failure mode $k$
/* Feature Extraction using CA-FPCA from all $P$ sensors */
for $p=1$ to $P$ do
       $\mathcal{I}_{k}^{p}\leftarrow$ K-means clustering on FPC-scores;
       for $k=1$ to $K$ do
             $\xi_{i,m,p}(k)\leftarrow$ Use $\mathcal{I}_{k}^{p}$ to obtain CA-FPC-scores (Eq. 2);
       end for
      $x_{i,p}\leftarrow$ Obtain standardized CA-FPC-scores;
       $q_{p}\leftarrow$ Number of CA-FPC-scores retained for sensor $p$;
end for
/* MGR-ASGL based Expectation-Maximization (EM) algorithm */
Initialize $\mathcal{P}_{k}\leftarrow\emptyset$ for all $k$;
Initialize $\Psi^{0}\leftarrow 0,\Pi^{0}\sim\texttt{Uniform}(K),\boldsymbol{\rho}^{0}\leftarrow\texttt{randint}$;
$iter\leftarrow 0$;
while $iter\leq iter_{max}$ do
       $\gamma_{i}^{iter}\leftarrow$ Use $x_{i,p}$ in Eq. (7) for all $i$ ;
        // E-step
       $\Pi^{iter}\leftarrow$ Update using Eq. (9) ;
        // M-step for $\Pi$
       $\boldsymbol{\rho}^{iter},\Psi^{iter}\leftarrow$ Solve Eq. (8) using $\sum_{p=1}^{P}\sqrt{q_{p}}$ and $\gamma_{i}^{iter}$;
       $iter\leftarrow iter+1$;
end while
$\gamma_{i}^{iter_{max}}\leftarrow$ Final responsibilities from Eq. (7);
$\boldsymbol{Z}_{i}[\operatorname*{arg\,max}_{k}{\gamma_{i}^{iter_{max}}}]\leftarrow 1$ for all $i$;
$\boldsymbol{Z}_{i}[\sim\operatorname*{arg\,max}_{k}{\gamma_{i}^{iter_{max}}}]\leftarrow 0$ for all $i$;
foreach $p$ and $k$ do
       if $\lVert\psi^{iter_{max}}_{p,k}\rVert_{2}\geq\delta$ then
             Append $p$ to $\mathcal{P}_{k}$ ;
              // Significant sensors for failure mode $k$
       end if
end foreach
Algorithm 1 Offline Sensor Selection Algorithm
Lemma 1.
The expectation of the CDLL w.r.t distribution $g$ is a lower bound for the IDLL i.e.,
$\mathbb{E}_{\boldsymbol{g}}\left[\ell_{C}(\boldsymbol{\Theta};\boldsymbol{Y},\mathbb{Z})\right]\leq\ell(\boldsymbol{\Theta};\boldsymbol{Y})$
Proof.
Let $g$ be an arbitrary distribution on $\boldsymbol{Z}_{i}[k]$. Then we can re-write equation (4) as follows:
$\ell(\boldsymbol{\Theta};\boldsymbol{Y})=\sum_{i=1}^{N}\ln\left(\sum_{k=1}^{K}\frac{g\left(\boldsymbol{Z}_{i}[k]\right)p\left(y_{i},\boldsymbol{Z}_{i}[k];\boldsymbol{\Theta}\right)}{g\left(\boldsymbol{Z}_{i}[k]\right)}\right)\vspace*{-1em}$
where,
$\begin{split}&p\left(y_{i},\boldsymbol{Z}_{i}[k];\boldsymbol{\Theta}\right)\\ &=\pi_{k}\cdot\left(\frac{\rho_{k}}{\sqrt{2\pi}}\exp\left(-\frac{1}{2}\left(\rho_{k}y_{i}-\psi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\psi}_{p,k}\right)^{2}\right)\right)\end{split}$
Since $\ln(.)$ is concave, from Jensen’s inequality we can write that,
$\ell(\boldsymbol{\Theta};\boldsymbol{Y})\geq\sum_{i=1}^{N}\sum_{k=1}^{K}g\left(\boldsymbol{Z}_{i}[k]\right)\ln{\frac{p\left(y_{i},\boldsymbol{Z}_{i}[k];\boldsymbol{\Theta}\right)}{g\left(\boldsymbol{Z}_{i}[k]\right)}}\vspace*{-1em}$
and thus,
$\ell(\boldsymbol{\Theta};\boldsymbol{Y})\geq\mathbb{E}_{\boldsymbol{g}}\left[\ell_{C}(\boldsymbol{\Theta};\boldsymbol{Y},\mathbb{Z})\right]-\sum_{i=1}^{N}\sum_{k=1}^{K}g\left(\boldsymbol{Z}_{i}[k]\right)\ln{g\left(\boldsymbol{Z}_{i}[k]\right)}\vspace*{-1em}$
The second term on the right side of the inequality is the negative entropy of the $g$ distribution. Since it is positive, the inequality holds. ∎
For every iteration of the EM algorithm, the E-step calculates the conditional expectation of the CDLL given the observed failure times, which we denote as $\mathbb{E}_{\boldsymbol{\Theta}_{\text{old}}}[\ell_{C}(\boldsymbol{\Theta};\boldsymbol{Y},\mathbb{Z})|\boldsymbol{Y}]$. In the M-step, we find the parameters that maximize the conditional expectation (see equation (8)). The parameter estimation equations of the EM algorithm are provided in III.B.5.
IV-B3 Adaptive Sparse Group Lasso (ASGL) penalty
The inclusion of features from non-informative sensors can negatively impact the accuracy of the RUL prediction. To address this issue, we penalize the negative-scaled conditional expectation of the CDLL function using the Adaptive Sparse Group Lasso (ASGL) penalty to obtain the following minimization problem given by 6.
$\begin{split}\min_{\boldsymbol{\Theta}}-\frac{1}{N}\mathbb{E}_{\boldsymbol{\Theta}_{\text{old}}}[\ell_{C}(\boldsymbol{\Theta};&\boldsymbol{Y},\mathbb{Z})|\boldsymbol{Y}]+\lambda\sum_{k=1}^{K}\pi_{k}\sum_{p=1}^{P}\sqrt{q_{p}}\times\\ &\left[(1-\alpha)\left\|\boldsymbol{\psi}_{p,k}\right\|_{2}+\alpha\left\|\boldsymbol{\psi}_{p,k}\right\|_{1}\right]\end{split}$ (6)
ASGL is an extension of lasso [39] that ensures the coefficients for all features of a non-informative group (sensor) are reduced to zero. In the context of regularization, adaptive refers to the penalty accounting for the number of CA-FPC-scores retained for each sensor i.e., $q_{p}$, and the mixing coefficient for each failure mode i.e., $\pi_{k}$. While the inclusion of $q_{p}$ ensures a sensor is not retained simply by having more coefficients, $\pi_{k}$ accounts for the class imbalance. $\lambda>0$ and $\alpha\in[0,1]$ are the hyperparameters for this model.
IV-B4 Expectation Maximization algorithm
We now describe the EM algorithm used to find a local optimum to the IDLL described. Given an initial estimate of the model parameters $\boldsymbol{\Theta}_{old}$, we perform the E-step and M-step as follows:
E-step: The E-step consists of computing the expectation of CDLL in (6) conditioned on $\boldsymbol{Y}$. Let $\mathcal{Q}\left(\boldsymbol{\Theta}\mid\boldsymbol{\Theta}_{\text{old}}\right)=-\mathbb{E}_{\boldsymbol{\Theta}_{\text{old}}}\left[\ell_{C}(\boldsymbol{\Theta};\boldsymbol{Y},\mathbb{Z})\mid\boldsymbol{Y}\right]$ denote the negative expectation of the CDLL. In order to compute $\mathcal{Q}\left(\boldsymbol{\Theta}\mid\boldsymbol{\Theta}_{\text{old}}\right)$, we need to compute $\gamma_{i,k}=$ $\mathbb{E}_{\boldsymbol{\Theta}_{\text{old }}}\left[Z_{i}[k]\mid\boldsymbol{Y}\right]$, which represents the responsibility of failure mode $k$ to system $i$ and is given by equation (7). For any system $i$, we denote the vector of the responsibilites $\gamma_{i,k}$ across all failure modes as $\boldsymbol{\gamma}_{i}$ such that, $\boldsymbol{\gamma}_{i}=\{\gamma_{i,1},...,\gamma_{i,K}\}$.
$\footnotesize\gamma_{i,k}=\frac{\pi_{k}^{\text{old }}\rho_{k}^{\text{old }}\times\exp\left(-\frac{1}{2}\left(\rho_{k}^{\text{old }}y_{i}-\psi_{0,k}^{\text{old }}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\psi}_{p,k}^{\text{old }}\right)^{2}\right)}{\sum_{l=1}^{K}\pi_{l}^{\text{old }}\rho_{l}^{\text{old }}\exp\left(-\frac{1}{2}\left(\rho_{l}^{\text{old }}y_{i}-\psi_{0,l}^{\text{old }}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\psi}_{p,l}^{\text{old }}\right)^{2}\right)}\vspace*{-0.5em}$ (7)
M-Step: During the M-step, we update the parameters by solving the optimization problem in (8) with respect to $\boldsymbol{\Theta}$.
$\begin{split}\min_{\boldsymbol{\Theta}}\frac{1}{N}Q(\boldsymbol{\Theta}\mid&\boldsymbol{\Theta}_{\text{old}})+\lambda\sum_{k=1}^{K}\pi_{k}\sum_{p=1}^{P}\sqrt{q_{p}}\times\\ &\left[(1-\alpha)\left\|\boldsymbol{\psi}_{p,k}\right\|_{2}+\alpha\left\|\boldsymbol{\psi}_{p,k}\right\|_{1}\right]\end{split}\vspace*{-1em}$ (8)
Improvement with respect to $\boldsymbol{\Pi}$: To update the mixing coefficients, we set $\rho_{k}=\rho_{k}^{\text{old}},\boldsymbol{\psi}_{0,k}=\boldsymbol{\psi}_{0,k}^{\text{old }},\boldsymbol{\psi}_{p,k}=\boldsymbol{\psi}_{p,k}^{\text{old }}$. and solve optimization problem (8). We utilize the update strategy from [37]. We start by identifying a feasible solution $\overline{\boldsymbol{\Pi}}^{\text{new}}=\left(\frac{1}{N}\sum_{i=1}^{N}\gamma_{i,1},\ldots,\frac{1}{N}\sum_{i=1}^{N}\gamma_{i,K}\right)^{T}$ and then update the mixing coefficients as described by equation (9), where $t^{\text{new }}\in(0,1]$ is chosen to be the largest value in $\left\{0.1^{k};k=0,1,2,\ldots\right\}$ such that the objective function does not increase.
$\boldsymbol{\Pi}^{\text{new}}=\boldsymbol{\Pi}^{\text{old}}+t^{\text{new}}\left(\overline{\boldsymbol{\Pi}}^{\text{new }}-\boldsymbol{\Pi}^{\text{old}}\right)\vspace*{-0.5em}$ (9)
Improvement with respect to $\boldsymbol{\rho}$ and $\boldsymbol{\Psi}$: For this, we utilize the updated mixing coefficients $\Pi^{\text{new }}$ of the previous step. Since the $\rho_{k}$’s and $\boldsymbol{\psi}_{p,k}$’s are independent for each failure mode, the expression in (8) decouples into $K$ distinct optimization problems. Once the estimators $\rho_{k}^{\text{new }},\psi_{0,k}^{\text{new }}$, and $\boldsymbol{\psi}_{p,k}^{\text{new }}$ are determined, they are used to update the cluster responsibilities in the E-step.
After the EM algorithm is completed, we obtain optimal $\boldsymbol{\Pi}$, $\boldsymbol{\rho}$, and, $\boldsymbol{\Psi}$. We substitute these optimal values to equation (7) to obtain the responsibilities $\boldsymbol{\gamma}_{i}$’s. For each pair of sensor data and failure time, we assign the failure mode with the largest responsibility. This results in class labels for the training set. To perform sensor selection for failure mode $k$, we compute $\left\|\boldsymbol{\psi}_{p,k}\right\|_{2}$, the $l_{2}$-norm of the regression coefficients for sensor $p$ and failure mode $k$. We deem a sensor informative for failure mode $k$ if the $l_{2}$-norm of its regression coefficients is greater than 0.001. The detailed step-wise implementation of the offline sensor selection step is given in algorithm 1.
V Online diagnosis and Prognostics
This section details the steps of the online part of our multi-failure mode prognostics framework. In this part, a fielded unit is degrading in real-time, and our goal is to predict the time remaining before the unit fails. We begin by showing how MFPCA is used to fuse the signals from multiple sensors into a parsimonious set of features. Then, we demonstrate how we utilize MFPCA to i) diagnose the active failure mode of the degrading unit, in real-time, and ii) predict its RUL.
Input : Smoothed signals from training & test units, Failure labels of training data from Algorithm 1, Set of informative sensors $\mathcal{P}_{k}$ from Algorithm 1
Output : Active failure mode in the test unit $k^{*}$, Remaining Useful Life (RUL) of the test unit $RUL_{test}$
/* Diagnostic Model Training */
$\zeta_{(i),h}\leftarrow$ MFPCA on training units (Eq. 10);
Train a KNN classifier using $\zeta_{(i),h}$;
/* Diagnostic Model Testing */
$\zeta_{test,h}\leftarrow$ MFPCA on test unit;
$k^{*}\leftarrow$ Mode of labels in KNN of $\zeta_{test,h}$;
/* Prognostic Model Training */
$\zeta^{k^{*}}_{(i),h}\leftarrow$ MFPCA: selected sensors (Eq. 14);
$c_{0}^{k^{*}},\dots,c_{h}^{k^{*}}\leftarrow$ Obtain $\operatorname*{arg\,min}$ in Eq. 17;
/* Prognostic Model Testing */
$\zeta^{k^{*}}_{test,h}\leftarrow$ MFPCA on test unit (Eq. 14);
$RUL_{test}\leftarrow$ Equation 18;
Algorithm 2 Online Diagnosis and Prognostics
V-A Sensor Fusion: MFPCA
Without loss of generality, let us redefine the sensor index as $\{1,2,\ldots,\mathcal{P}\}$, where $\mathcal{P}$ is number of sensors we seek to fuse. Now assume that we have monitored an online unit (operating in the field) for a duration given by the time interval $[0,t^{*}]$. Since we are interested in predicting the remaining lifetime, we only use signals from units with failure times greater than $t^{*}$. We define this set as the “$t^{*}-$training” set, whose indices we denote by the set $\mathcal{I}_{t^{*}}$, which has cardinality $N_{t^{*}}$. Furthermore, we smooth the degradation signals using the ’rloess’ method as described in [40]. This method uses weighted least squares with a 2nd-order polynomial for smoothing. The weights come from a bisquare weight function requiring a specified bandwidth parameter. This method also identifies outliers and assigns them weights of zero to improve signal mean estimation. To perform MFPCA, we model the $i$th smoothed multivariate signal evaluated at time $t$ as $\boldsymbol{s}_{i}(t)=\boldsymbol{u}_{i}(t)+\boldsymbol{\epsilon}_{i}(t)$, where $\boldsymbol{u}_{i}(t)\in\mathcal{R}^{\mathcal{P}}$ is a random multivariate function with mean $\boldsymbol{\mu}(t)$ and covariance function $\mathcal{C}(t,t^{\prime})$, $\boldsymbol{\epsilon}_{i}(t)\in\mathcal{R}^{\mathcal{P}}$, and $t\in[0,t^{*}]$ are i.i.d. zero-mean error vectors with covariance $\Sigma_{\boldsymbol{\epsilon}}$. We assume $\boldsymbol{u}_{i}(t)$ and $\boldsymbol{\epsilon}_{i}(t)$ are independent of each other. Using the first $H$ multivariate eigenfunctions, we can represent the multivariate signal by equation 10 where the MFPC-scores $\zeta$’s are computed according to equation 11.
$\boldsymbol{s}_{i}(t)=\boldsymbol{\mu}(t)+\sum_{h=1}^{H}\zeta_{i,h}\boldsymbol{\varpi}_{h}(t)+\boldsymbol{\epsilon}_{i}(t),\vspace*{-1.2em}$ (10)
[…]
V-B Failure Mode Diagnosis: K-nearest Neighbors
In order to diagnose the failure mode of a degrading unit at time $t^{*}$, we apply MFPCA on the $t^{*}$ training dataset. Next, $K$ nearest neighbors (KNN) is applied to the extracted MFPC scores to identify the active failure mode. Here, the distance metric of the KNN classifier is defined using the Euclidean distance between the MFPC scores. Given this diagnosis, we now develop a model for predicting the unit RUL.
V-C RUL Prediction: Weighted Functional Regression
To predict RUL, we use a weighted time-varying functional regression model that maps multivariate signal features to the $\ln{TTF}$. First, we discuss the basics of a time-varying functional regression and then motivate the case for a weighted version of the model
V-C1 Time-Varying Functional Regression
When predicting RUL, we are often faced with updating the relationship between time-to-failure and the current values of the extracted signal features (FPC scores). The predictor trajectories up to the current time are represented by time-varying FPC scores, which are continuously updated as time progresses and are considered to be time-varying predictor variables for the functional regression model. This is referred to as time-varying functional regression. It was first proposed by [41] and has been used for prognostics in [42, 13].
In our problem setting, let $k^{*}$ denote the diagnosed failure mode and $\mathcal{I}_{t^{*}}^{k^{*}}\subset\mathcal{I}_{t^{*}}$ denote the set of indices for training units that failed due to failure mode $k^{*}$, but survived up to time $t^{*}$. Furthermore, let $\mathcal{P}_{k^{*}}\subset\{1,2,...,\mathcal{P}\}$ denote the set of sensors selected (identified as most informative) for failure mode $k^{*}$. To predict RUL, we fit a functional regression model using least squares as shown in equation (12), where $y_{i}=\ln{TTF_{i}}$, $\varphi_{0}^{k^{*}}$ is the bias term, $\boldsymbol{\varphi}^{k^{*}}(t)\in\mathcal{R}^{|\mathcal{P}_{k^{*}}|}$ is the multivariate coefficient function, and $\boldsymbol{s}_{i}^{k^{*}}(t)$ is the multivariate signal consisting only of sensors significant to failure mode $k^{*}$.
[…]
We expand the coefficient function using the eigenfunctions obtained from applying MFPCA on the training set $\mathcal{I}_{t^{*}}^{k^{*}}$. The regression coefficient function can now be expanded as shown in equation (13), where $H_{k^{*}}$ denotes the number of MFPC-scores that were retained for failure mode $k^{*}$.
$\boldsymbol{\varphi}(t)=\sum_{h=1}^{H_{k^{*}}}c_{h}^{k^{*}}\boldsymbol{\varpi}_{k}^{k^{*}}(t)\vspace*{-0.5em}$ (13)
Likewise, we expand the multivariate signals using the eigenfunctions as shown in equation (14).
$\boldsymbol{s}_{i}^{k^{*}}(t)=\boldsymbol{\mu}^{k^{*}}(t)+\sum_{h=1}^{H_{k^{*}}}\zeta_{i,h}^{k^{*}}\boldsymbol{\varpi}_{h}^{k^{*}}(t)\vspace*{-0.5em}$ (14)
Since the eigenfunctions are orthogonal, substituting equations (13) and (14) into equation (12) yields the following optimization problem,
$\min_{c_{0}^{k^{*}},c_{1}^{k^{*}},\ldots,c_{H_{k^{*}}}^{k^{*}}}\sum_{i\in\mathcal{I}_{t^{*}}^{k^{*}}}\left(y_{i}-c_{0}^{k^{*}}-\sum_{h=1}^{H_{k^{*}}}\zeta_{i,h}^{k^{*}}c_{h}^{k^{*}}\right)^{2}\vspace*{-1em}$ (15)
where,
$c_{0}^{k^{*}}=\varphi_{0}^{k^{*}}+\int_{0}^{t^{*}}\sum_{h=1}^{H_{k^{*}}}c_{h}^{k^{*}}\boldsymbol{\varpi}_{h}^{k^{*}}(t)^{T}\boldsymbol{\mu}^{k^{*}}(t)\vspace*{-0.5em}$ (16)
Weighted Version: We note that the offline sensor selection algorithm may generate some outliers, which can impact the accuracy of the regression problem. These outliers become more apparent after identifying the active failure mode and the MFPC scores to estimate the time-varying functional regression model. To address this issue, we use weighted least squares to estimate the time-varying functional regression mode. To derive the weights, we first measure the Euclidean distance between the MFPC score corresponding to a unit $i$ and the centroid of the cluster of the active failure mode. We then use the reciprocal of that distance to define the corresponding weight for the unit, $w_{i}$, of that unit. If the unit’s failure is misclassified, $w_{i}$ would be large, and thus, unit $i$ is assigned a smaller weight in the regression problem. Next, We regress the $\ln{TTF}$ on the MFPC scores using a weighted linear regression shown in equation (17). The solution to this problem can be found using weighted least squares regression [43]. For variable selection, we include the lasso penalty in the optimization problem.
$\begin{split}\min_{c_{0}^{k^{*}},c_{1}^{k^{*}},\ldots,c_{H_{k^{*}}}^{k^{*}}}\sum_{i\in\mathcal{I}_{t^{*}}^{k^{*}}}&\left(\sqrt{w_{i}}y_{i}-\sqrt{w_{i}}c_{0}^{k^{*}}-\sum_{h=1}^{H_{k^{*}}}\sqrt{w_{i}}\zeta_{i,h}^{k^{*}}c_{h}^{k^{*}}\right)^{2}\\ &+\lambda\sum_{h=1}^{H_{k^{*}}}\lVert c_{h}^{k^{*}}\rVert\end{split}\vspace*{-0.5em}$ (17)
V-C2 RUL prediction
As we have shown, MFPCA enables the transformation of the functional regression problem to a multiple linear regression problem, which enables us to solve the problem only by estimating $H_{k^{*}}+1$ parameters. After fitting the regression coefficients, we extract the MFPC scores for the test signal $\zeta_{test,1}^{k^{*}},\ldots,\zeta_{test,H_{k^{*}}}^{k^{*}}$. These scores are then used to predict the RUL at time $t^{*}$ using equation (18).
${RUL}_{test}=\exp\left(c_{0}^{k^{*}}+\sum_{k=1}^{H_{k^{*}}}\zeta_{test,h}^{k^{*}}c_{k}^{k^{*}}\right)-t^{*}$ (18)
VI Case Study 1: Simulated Dataset
We present a case study using a simulated data set to evaluate the performance of our methodology in terms of its ability to cluster the training data, effectiveness in selecting informative sensors for each failure mode, and accuracy in predicting the RUL. Furthermore, we test the robustness of our methodology to varying degrees of signal noise. In section VI-A, we describe the simulation model. In section VI-B, we display the simulated data and evaluate the performance of the MGR-ASGL model at clustering and sensor selection. Finally, we evaluate the prediction accuracy of our online prognostics approach in section VI-C.
VI-A Simulation Model
We utilize a modified version of the simulation model used in [13] to generate degradation signals for a set of engineering systems. We assume that each system is monitored by 20 sensors and exhibits one of two possible failure modes that we refer to as failure mode 1 (FM 1) and failure mode 2 (FM 2). We generate data for 200 systems for each failure mode. Out of the 200 systems, we use 160 systems for model training and 40 systems for testing model performance. The simulated data consists of the sensor signals for each system and the corresponding failure time, which is utilized later as the ground truth to evaluate prediction accuracy.
The signals of the underlying degradation process for system $i$ exhibiting failure mode $k$ are generated according to equation (19). The signal is parameterized by the random coefficient $\theta_{i}^{k}\sim N\left(\mu_{k},0.1^{2}\right)$, which is used to control the degradation rate. We let $\mu_{1}=1$, and $\mu_{2}=0.8$ be the degradation rates for FM 1 and FM 2, respectively. The TTF is computed as the first time the degradation signal $s_{i}^{k}(t)$ crosses a soft failure threshold $D_{k}$. We set the failure thresholds for FM 1 and FM 2 as $D_{1}=2$ and $D_{2}=1.5$, respectively.
$s_{i}^{k}(t)=-\frac{\theta_{i}^{k}}{\ln t}$ (19)
This paragraph describes the procedure for simulating degradation signals for each sensor. Here we consider a practical scenario where only a subset of sensors is correlated with the underlying degradation process associated with a given failure mode. These sensors are the informative sensors and their signal trends reflect the severity and progression of the degradation process. Note that sensors that do not exhibit any consistent trends leading up to a failure event are deemed noninformative. We define $\mathcal{J}_{k}$ as the index set of informative sensors for failure mode $k$, and $\mathcal{N}_{k}$ as the set of noninformative sensors such that $\mathcal{N}_{k}\cap\mathcal{J}_{k}=\{\}$ (null set). Out of the 20 sensors, we set $\mathcal{J}_{1}=\{5,12,16,19\}$ as the set of informative sensors for FM 1 and $\mathcal{J}_{2}=\{3,7,9,19\}$ as the set of informative sensors for FM 2. The $p^{th}$ sensor signal for the $i^{th}$ system exhibiting failure mode $k$ is generated using equation (20).
$s_{i,p}^{k}(t)=-\frac{\theta_{i,p}^{k}}{\ln t}+\epsilon_{i,p}^{k}(t)$ (20)
In this equation, $\epsilon_{i,p}^{k}(t)\sim N\left(0,\left(\sigma_{k}^{p}\right)^{2}\right)$ is a white noise process parameterized by $\sigma_{k}^{p}=\mu_{k}/SNR_{p}^{k}$, where $SNR_{p}^{k}$ is the signal-to-noise for sensor $p$ and failure mode $k$. To account for the correlation between the signal trends of the informative sensors and the underlying degradation process, we generate $\theta_{i,p}^{k}$ from the following conditional distribution $\theta_{i,p}^{k}\mid\theta_{i}^{k}\sim N\left(\mu_{k}\left(1-\sqrt{1-\rho_{p}^{k}}\right),(0.1)^{2}\right)$ such that the correlation, $\rho_{p}^{k}$, is uniformly sampled from the interval $[0.80,0.99]$. The process for generating signals for the noninformative sensors is similar to that of the informative sensors, except the interval from which we sample $\rho_{p}^{k}$ is $[0.1,0.6]$. Finally, we randomly assign a sign to $\theta_{i,p}^{k}$ to allow for signals that increase or decrease in response to degradation.
In this simulation, we seek to test the robustness of our methodology to various levels of signal noise. Therefore, we generate three datasets, each with a unique interval from which to sample $SNR_{p}^{k}$. For $p\in\mathcal{J}_{k}$, $SNR_{p}^{k}$ is sampled uniformly from $[2,5]$, $[5,8]$, and $[8,11]$ for datasets 1, 2, and 3, respectively. For $p\in\mathcal{N}_{k}$, $SNR_{p}^{k}$ is sampled uniformly from $[1,3]$ for all three datasets. Figure 2 displays samples of our sensor signals in a 3$\times$4 grid. Each row displays signals from four sensors (1, 3, 5, 19) for a particular signal-to-noise ratio. Sensor 1 is not informative for either failure mode so no discernible trend is present. Sensors 3 and 5 are informative for FM 2 and FM 1, respectively. For these sensors, one signal trends downward due to degradation while the other signal appears to resemble random noise. Finally, for sensor 19, the signal decreases in response to both failure modes.
Figure 2: Case study 1 data: Signals in sensors 1, 3, 5, 19 for both failure modes 1 and 2 under three different SNRs i.e., [2, 5], [5, 8], and [8, 11]
VI-B Failure Mode Labeling and Sensor Selection
We analyze the ability of the MGR-ASGL model to accurately cluster the degradation signals into one of the two possible failure modes along with selecting the subset of informative sensors. We start by utilizing CA-FPCA as mentioned in IV-A to extract features for each sensor separately. To determine the number of CA-FPC scores to retain, we use the 95% fraction of variance explained (FVE) criterion. For each cluster, we select the first $q_{p}^{k}$ CA-FPC-scores whose eigenvalues sum to at least 0.95. Since the number of CA FPC scores to retain can vary for each cluster, we retain the maximum of $q_{p}^{k}$ over both failure modes. For the remainder of this paper, we use 95% FVE criterion for determining the number of FPC scores, CA-FPC scores, and MFPC scores to retain. The CA-FPC scores serve as inputs to the MGR-ASGL model. Since this model has two tuning parameters, we utilize 5-fold cross-validation with $\lambda\in\{0.0050,0.0258,0.0466,...,0.3792,0.4000\}$ and $\alpha\in\{0.00,0.25,0.50,0.75,1.00\}$. We use minimum mean-squared error (minMSE) criterion, which selects the tuning parameters that minimize the mean-squared prediction error averaged over all folds. Table I displays the tuning parameter selection using both criteria over all three datasets. For each dataset and criterion, we fit the MGR-ASGL model using these tuning parameters.
TABLE I: Tuning parameter selection for case study 1
$SNR_{p}^{k}$ [2,5] [5,8] [8,11]
$\lambda$ 0.0466 0.0258 0.0258
$\alpha$ 1 0 0.25
We then evaluate the performance of the MGR-ASGL model on the three datasets. While clustering is an unsupervised learning task, we know the ground truth regarding the failure modes. Therefore, we evaluate the ability of the clustering algorithm to match the ground truth. Furthermore, we have predetermined which sensors are most informative and we would like to determine whether or not the MGR-ASGL selects these sensors. The results of this performance evaluation are shown in Table II.
TABLE II: MGR-ASGL performance for case study 1
$SNR_{p}^{k}$ FM Accuracy Sensors Selected
[2, 5] 1 0.800 1:3, 5, 7:9, 11:16, 18:20
2 0.875 1, 3:4, 6:10, 13:14, 17:20
[5, 8] 1 0.506 1:7,9:12,14:20
2 0.669 1:20
[8, 11] 1 0.825 1,3:20
2 0.788 1:3,5:20
For all scenarios, the initial failure mode labels were determined randomly. Table II shows that for the lowest and the highest signal-to-noise ratios, the clustering algorithm was capable of achieving at worst 78.8% accuracy for a particular failure mode. The exception is the dataset [5,8], where the algorithm has low accuracy at clustering. This is due to the tuning parameter combination resulting in minimal coefficient shrinkage that enabled the nonsignificant sensors to influence cluster assignments. A slightly larger $\lambda$ value would decrease this error. For all three models, several sensors that are deemed ”noninformative” were selected. This can be attributed to some of these sensors having high values of $\rho_{p}^{k}$ relative to the specified range of [0.1,0.6]. However, simply reporting the sensors selected does not account for how informative they are. Table III reports the four sensors with the highest $l_{2}$-norm of the regression coefficients associated with that sensor. Except for the [5,8] dataset, at least three of the four sensors with the largest $l_{2}$-norm for each failure mode are elements of the set of informative sensors. This indicates that the MGR-ASGL model results in the informative sensors having larger regression coefficients (in magnitude) than the noninformative sensors. Given the selected sensors, we want to test the robustness of the online portion of the methodology to predict remaining life.
TABLE III: $l_{2}$-norms of four most informative sensors for FM1 & FM2
$SNR_{p}^{k}$ FM 1 FM 2
Sensor ID $l_{2}$-norm Sensor ID $l_{2}$-norm
5 1.809 9 1.667
[2, 5] 12 0.912 7 0.849
16 0.604 19 0.303
7 0.269 6 0.218
5 1.137 9 1.153
[5, 8] 12 0.641 17 0.547
17 0.500 19 0.537
18 0.499 4 0.504
12 1.982 9 2.150
[8, 11] 5 1.422 3 0.559
16 1.016 19 0.540
18 0.394 17 0.428
VI-C RUL Prediction
Each of the datasets consist of 80 test systems (40 from each failure mode). For each test system, we observe degradation up until the following life percentiles (in retrospect): 10%, 20%, 30%, 40%, 50%, 60%, 70%, 80%, and 90%. We smooth the training and the test signals using a bandwidth parameter of 0.5. Next, we perform MFPCA selecting the number using only the training systems that survived up to the observation time $t^{*}$ corresponding to each life percentile. For a given test signal, we calculate its MFPC scores. The active failure mode is diagnosed by being assigned to the most frequent failure mode present in its closest $0.1N_{t^{*}}$ neighbors. Once the active failure mode is identified, the MFPC scores of the test signals are updated and recomputed using the subset of training signals associated with the active failure mode. Next, we fit the weighted regression model and predict the RUL using the scores of the test signal. Prediction accurately is calculated retrospectively. We compute the “Estimated Life” as the current operating time plus the predicted RUL. A relative error is then computed using equation (21).
[…]
The performance of our methodology on predicting RUL is shown in Figure 3. In this figure, the relative errors have a generally decreasing trend for all datasets, with the [8,11] dataset tending to have smaller relative errors on average. Furthermore, the variability appears to be much larger when making predictions early in the system’s life rather than later. Interestingly, the relative error increased for the [8,11] dataset when making predictions at 90% of the system’s life. This increase may be attributed to the increased influence of incorrectly clustered training samples affecting the RUL prediction. Since this dataset has less noise, the influence of clustering may be more pronounced compared to the noisier datasets, especially if training samples have been removed for failing before 90% of the monitored system’s lifetime.
VII Case Study 2: NASA Turbofan Engine Dataset
Figure 4: The system diagram used in Case Study 2 adapted from [44]
To further validate our methodology, we evaluate its performance on one of the Commercial Modular Aero-Propulsion System Simulation (C-MAPSS) datasets [44]. C-MAPSS is a simulation module that generates turbofan engine degradation datasets. Figure 4 displays a diagram of this turbofan engine. In this paper, we utilize C-MAPSS-3, which consists of both a training and a test set. Both sets contain 100 observed engine failures caused by either fan degradation or high-pressure compressor degradation. Furthermore, both sets contain data from 21 sensors, which are listed in Table IV. The signals for these sensors are displayed in Figure 5. These sensors monitor the systems from a good-as-new state to engine failure. However, sensors 1, 5, 6, 10, 16, 18, and 19 contain little to no information and are thus removed from the analysis. While the dataset provides the potential causes of failure, it does not provide the failure labels. Furthermore, the test set consists of systems whose degradation is only partially observed along with their corresponding failure times. Therefore, our goal is to use our methodology to predict the remaining life of each testing unit following the final cycle at which their degradation was observed.
TABLE IV: Sensors used in case study 2 (C-MAPPS-3 [44])
Sensor ID Sensor Name Functionality
1 T2 Total temperature at fan inlet
2 T24 Total temperature at LPC outlet
3 T30 Total temperature at HPC outlet
4 T50 Total temperature at LPT outlet
5 P2 Pressure at fan inlet
6 P15 Total pressure in bypass-duct
7 P30 Total pressure at HPC outlet
8 Nf Physical fan speed
9 Nc Physical core speed
10 epr Engine pressure ratio (P50/P2)
11 Ps30 Static pressure at HPC outlet
12 phi Ratio of fuel flow to Ps30
13 NRf Corrected fan speed
14 Nrc Corrected core speed
15 BPR Bypass ratio
16 farB Burner fuel-air ratio
17 htBleed Bleed enthalpy
18 Nfdmd Demanded fan speed
19 PCNfRdmd Demanded corrected fan speed
20 W31 HPT coolant bleed
21 W32 LPT coolant bleed
Figure 5: Raw sensor data used for offline step in case study 2. The data is inherently unlabeled i.e., failure modes are not known.
Our first step is to perform traditional FPCA on the signals for each sensor. We obtain initial failure mode labels for the training set by applying K-means clustering to all FPC-scores. Then, for each sensor, we apply K-means clustering to the FPC-scores associated with that sensor to obtain the covariate information needed to perform CA-FPCA. Following feature extraction, we use 3-fold cross-validation to determine $(\lambda,\alpha)$. The range of values for these tuning parameters are identical to that of the simulation case study. The combination of tuning parameters that minimize the mean squared error across the three folds is $(\lambda_{MSE},\alpha_{MSE})=(0.1089,1)$. Table V displays the $l_{2}$-norms of the regression coefficients for each sensor. The sensors selected are those with positive $l_{2}$-norms.
TABLE V: Sensor Selection for Case Study 2
Sensor ID FM 1 FM2 Sensor ID FM 1 FM 2
2 0.266 0 12 0 1.148
3 0.584 0 13 0.90 0
4 0.897 0.033 14 0.148 0.119
7 0 0.877 15 0.501 0
8 0.117 0 17 0.230 0.042
9 0.545 0.291 20 0 0.573
11 0 0 21 0.116 0.318
In addition to performing sensor selection, the MGR-ASGL model provides cluster labels for each observed failure. These labels are utilized for the online prediction portion of the methodology. For the online prediction step, we first smooth the degradation signals using 5-fold cross-validation to select the bandwidth parameter. After smoothing, we perform MFPCA on signals from the sensors selected across both failure modes to build our classifier. Then, we classify the failure mode based on the $0.1N_{t^{*}}$ nearest neighbors. Given this classification, we perform MFPCA on signals from the sensors selected for the classified failure mode. Finally, we perform Weighted Time-varying Functional Regression to regress $lnTTF$ on the MFPC-scores using Leave-one-out cross-validation to select the tuning parameter.
Benchmarking: To analyze our methodology, we compare its performance to that of [32] (Chehade et al. 2018), whose methodology assumes knowledge of the failure modes but does not perform sensor selection (outside of removing the blank sensors), and [33] (Li et al. 2023), whose methodology assumes knowledge of a fraction of the training set and does perform sensor selection. Figure 6 displays the relative prediction error averaged over all test systems with RUL at most 100, 80, 60, 40, and 20 cycles. From this figure, our methodology becomes more accurate as the RUL decreases. Furthermore, our methodology is more accurate than both when making longer predictions. For test units whose RUL is at most 20, our accuracy is between that of [32] and [33]. Only at RUL at most 40, does our methodology perform worse than both. To further analyze our methodology, we refer to Figure 7, which displays boxplots showing the distribution of the relative prediction error for test units with remaining life in the following intervals: $[0,25],[26,50],[51,75],[76,100],[101,125],[126,150]$. The results indicate that our methodology achieves a lower median relative error than [32] in all intervals while maintaining comparable variability. However, the variability for systems with RUL between 25 and 50 is higher for our methodology than for [32]. This indicates a targeted area for improvement of our methodology. However, the general trend is that sensor selection resulted in a more accurate RUL prediction than no sensor selection. Furthermore, we were capable of achieving more accurate RUL prediction than both methods from earlier in the systems’ lives despite not having information about the actual failure mode.
[…]
VIII Conclusion
In this paper, we present a methodology for predicting the RUL of systems with multiple failure modes using multiple sensors. This methodology consists of two phases: an offline phase and an online phase. Assuming that failure modes are mutually exclusive, the offline phase consists of a framework for diagnosing the observed failures in the training set and selecting an optimal subset of sensors for monitoring. In this framework, we utilize CA-FPCA to capture the multiple failure modes exhibited by the sensor signals enabling the extraction of informative features. These features are used by our proposed EM algorithm to train the MGR-ASGL model which simultaneously diagnoses the training set failures and selects an optimal subset of informative sensors for each failure mode. In the online phase, we use MFPCA to fuse all the signals of the selected sensors into a parsimonious set of features to diagnose the failure mode for an actively degrading system. Then MFPCA is re-used to fuse the signals from informative sensors of the diagnosed failure mode for RUL prediction. RUL prediction is done using a time-varying weighted functional regression model.
To assess the performance of our methodology, we implemented it on two datasets – a) a simulated dataset and b) the NASA turbofan engine degradation dataset. For the simulated dataset, we tested the robustness of our methodology to various degrees of signal noise. We demonstrated that with a sufficiently high $\lambda$ value, the clustering matches the actual clusters at an accuracy of about 80%. Furthermore, the most significant sensors occupied 3 of the top 4 slots with regard to $l_{2}$-norm of regression coefficients. Finally, the methodology showed high accuracy with an average relative prediction error of no larger than 3, irrespective of the percentage of lifetime the prediction was performed.
In our real-world data case study, results show that our model makes early predictions more accurately than the two benchmarks while achieving comparable accuracy in later predictions. To achieve this, we identified 10 sensors to monitor the first failure mode while identifying 8 sensors to monitor the second failure mode. Our methodology removed one sensor completely. The region in which our methodology performed worst relative to the two baselines was when the RUL was between 25 and 50. A possibility for improvement would be to relax the assumption of a linear relationship between the $lnTTF$ and the sensor signals. However, embedding a variable selection technique within a nonlinear regression model is an open topic in the literature. Therefore, the extension of this work to a nonlinear setting is a topic for future work.
Acknowledgments
This effort is supported by NASA under grant number 80NSSC19K1052 as part of the NASA Space Technology Research Institute (STRI) Habitats Optimized for Missions of Exploration (HOME) ‘SmartHab’ Project. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Aeronautics and Space Administration.

### v2

@@
 The cluster assignments and the selected sensors are then utilized in real-time to first diagnose the active failure mode and then to predict the system RUL.
-We validate the complete pipeline of the methodology using a simulated dataset of systems exhibiting two failure modes and on a turbofan degradation dataset from NASA.
-Index Terms:
-Multiple failure modes, Sensor-fusion, prognostics, Sensor Selection, Adaptive Sparse Group Lasso, Mixture of Gaussian regressions, Functional Principal Component Analysis
-publicationid: pubid: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
-I Introduction
+We validate the methodology using a simulated dataset of systems exhibiting two failure modes and on NASA turbofan degradation dataset.
+keywords:
+Multiple Failure Modes , Remaining Useful Life , Sensor Fusion , Functional Principal Component Analysis , Mixture of Gaussian Regression , Adaptive Sparse Group Lasso , Expectation Maximization
+journal: Reliability Engineering & System Safety
+[…]
+[…]
+{highlights}
+We propose a two-stage prognostic framework for multi-sensor systems with multiple failure modes.
+Offline: An unsupervised method labels failure events and ranks sensor importance.
+Online: Sensor features are fused for failure mode diagnosis, then key sensors predict RUL.
+Our method achieves $80\%$ clustering accuracy and $<3\%$ RUL prediction error in simulations.
+In NASA turbofan data, early predictions outperform benchmarks while reducing input by $30\%$.
+1 Introduction
 Prognostic models aim to estimate the remaining useful life (RUL) of critical components and systems.
@@
   • Finally, we provide a two-step sequential approach to leverage this knowledge of failure labels and informative sensors to (1) perform active failure mode diagnosis and (2) subsequent RUL prediction on systems degrading in real-time.
-The paper mentions this as the “online diagnostics and prognostics” step.
+The paper mentions this as the “online diagnosis and prognostics” step.
 The remainder of the paper is outlined as follows.
-In Section II, we review the relevant literature associated with this research area.
+In Section 2, we review the relevant literature associated with this research area.
 In Section III, we outline our proposed multi-failure mode prognostics framework.
-In Section IV, we present our offline sensor selection scheme that utilizes an EM algorithm.
-In Section V, we present our online diagnostics and prognostics method.
+In Section 4, we present our offline sensor selection scheme that utilizes an EM algorithm.
+In Section 5, we present our online diagnostics and prognostics method.
 It discusses a strategy to fuse informative sensors for real-time diagnosis of failure modes and subsequent prediction of RUL.
-Our methodology is demonstrated through two case studies – one using a simulated dataset in section VI and the other NASA turbofan engine dataset in section VII.
-We conclude by summarizing our framework and key findings in Section VIII.
-II Literature Review
+Our methodology is demonstrated through two case studies – one using a simulated dataset in section 6 and the other NASA turbofan engine dataset in section 7.
+We conclude by summarizing our framework and key findings in Section 8.
+2 Literature Review
 Data-driven prognostic models utilize a variety of approaches.
@@
 While the failure labels are unknown, the total number of potential failure modes for the system is assumed to be known.
-III Overview of the Prognostics Framework
+3 Overview of the Prognostics Framework
 This section describes our multiple failure mode prognostics framework, which is summarized in the flowchart in Figure 1. This framework is comprised of two parts:
@@
 Figure 1: Flowchart of the multiple failure mode prognostics framework
-IV Offline Sensor Selection Algorithm
+4 Offline Sensor Selection Algorithm
 We consider a scenario where a system is monitored by $P$ sensors.
@@
 The next two subsections discuss details about CA-FPCA and the penalized Mixture of Gaussian regression model (MGR-ASGL).
-IV-A Feature Extraction: CA-FPCA
+4.1 Feature Extraction: CA-FPCA
 CA-FPCA extends conventional FPCA to handle complex data structures where the dynamics of time-varying signals are affected by one or more additional factors, i.e.
@@
 We then utilize this $\ boldsymbol{x}_{i,p}$’s as predictors for fitting the MGR-ASGL model.
-IV-B Optimization of the MGR-ASGL model
+4.2 Optimization of the MGR-ASGL model
 To perform sensor selection, we seek to model the distribution of $\ln{TTF}$ conditioned on the observed set of features.
@@
 Finally, we provide details for implementing the EM algorithm.
-IV-B1 Mixture of Gaussian Regression (MGR) model
+4.2.1 Mixture of Gaussian Regression (MGR) model
 We assume that the distribution of $y_{i}$ conditioned on the observed features $\boldsymbol{x}_{i,1},\ldots,\boldsymbol{x}_{i,P}$} and $\boldsymbol{Z}_{i}[k]=1$ is Gaussian.
@@
 This marginal distribution is the MGR model.
-$\begin{split}f_{\vartheta}(y_{i}\mid\boldsymbol{x}_{i,1},&\ldots,\boldsymbol{x}_{i,P})=\sum_{k=1}^{K}\pi_{k}\frac{1}{\sqrt{2\pi}\sigma_{k}}\times\\ &\ \exp\left(-\frac{1}{2\sigma_{k}^{2}}\left(y_{i}-\beta_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\beta}_{p,k}\right)^{2}\right)\end{split}$ (3)
-IV-B2 Log-Likelihood
+$f_{\vartheta}(y_{i}\mid\boldsymbol{x}_{i,1},\ldots,\boldsymbol{x}_{i,P})=\sum_{k=1}^{K}\pi_{k}\frac{1}{\sqrt{2\pi}\sigma_{k}}\times\ \exp\left(-\frac{1}{2\sigma_{k}^{2}}\left(y_{i}-\beta_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\beta}_{p,k}\right)^{2}\right)$ (3)
+4.2.2 Log-Likelihood
 […]
@@
 This function is called as the ”incomplete-data log-likelihood” (IDLL), since we do not know the cause of the observed failure in the data.
-$\begin{split}\ell(\boldsymbol{\Theta};&\boldsymbol{Y})=\sum_{i=1}^{N}\ln\sum_{k=1}^{K}\pi_{k}\times\\ &\ \frac{\rho_{k}}{\sqrt{2\pi}}\exp\left(-\frac{1}{2}\left(\rho_{k}y_{i}-\psi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\psi}_{p,k}\right)^{2}\right)\end{split}\vspace*{-1em}$ (4)
+$\ell(\boldsymbol{\Theta};\boldsymbol{Y})=\sum_{i=1}^{N}\ln\sum_{k=1}^{K}\pi_{k}\times\ \frac{\rho_{k}}{\sqrt{2\pi}}\exp\left(-\frac{1}{2}\left(\rho_{k}y_{i}-\psi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\psi}_{p,k}\right)^{2}\right)$ (4)
 Definition 2.
 The “complete-data log-likelihood” (CDLL) given by equation (5) assumes knowledge of the underlying causes of the observed failure in the data.
-$\begin{split}\ell_{C}(\boldsymbol{\Theta};&\boldsymbol{Y},\mathbb{Z})=\sum_{i=1}^{N}\sum_{k=1}^{K}\boldsymbol{Z}_{i}[k]\ln\pi_{k}\left(\frac{\rho_{k}}{\sqrt{2\pi}}\right)\\ &\exp\left(-\frac{1}{2}\left(\rho_{k}y_{i}-\psi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\psi}_{p,k}\right)^{2}\right)\end{split}\vspace*{-1em}$ (5)
+$\ell_{C}(\boldsymbol{\Theta};\boldsymbol{Y},\mathbb{Z})=\sum_{i=1}^{N}\sum_{k=1}^{K}\boldsymbol{Z}_{i}[k]\ln\pi_{k}\left(\frac{\rho_{k}}{\sqrt{2\pi}}\right)\exp\left(-\frac{1}{2}\left(\rho_{k}y_{i}-\psi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\psi}_{p,k}\right)^{2}\right)$ (5)
 The IDLL given in equation (4) is difficult to optimize globally [38].
@@
 Output : Vector $\boldsymbol{Z}$ indicating failure labels, Set of informative sensors $\mathcal{P}_{k}$ for each failure mode $k$
-/* Feature Extraction using CA-FPCA from all $P$ sensors */
-for $p=1$ to $P$ do
-       $\mathcal{I}_{k}^{p}\leftarrow$ K-means clustering on FPC-scores;
-       for $k=1$ to $K$ do
-             $\xi_{i,m,p}(k)\leftarrow$ Use $\mathcal{I}_{k}^{p}$ to obtain CA-FPC-scores (Eq. 2);
-       end for
-      $x_{i,p}\leftarrow$ Obtain standardized CA-FPC-scores;
-       $q_{p}\leftarrow$ Number of CA-FPC-scores retained for sensor $p$;
-end for
-/* MGR-ASGL based Expectation-Maximization (EM) algorithm */
-Initialize $\mathcal{P}_{k}\leftarrow\emptyset$ for all $k$;
-Initialize $\Psi^{0}\leftarrow 0,\Pi^{0}\sim\texttt{Uniform}(K),\boldsymbol{\rho}^{0}\leftarrow\texttt{randint}$;
-$iter\leftarrow 0$;
-while $iter\leq iter_{max}$ do
-       $\gamma_{i}^{iter}\leftarrow$ Use $x_{i,p}$ in Eq. (7) for all $i$ ;
+/* Feature Extraction from all $P$ sensors (CA-FPCA) */
+1 for $p=1$ to $P$ do
+2       $\mathcal{I}_{k}^{p}\leftarrow$ K-means clustering on FPC-scores;
+3       for $k=1$ to $K$ do
+4             $\xi_{i,m,p}(k)\leftarrow$ Use $\mathcal{I}_{k}^{p}$ to obtain CA-FPC-scores (Eq. 2);
+5
+6       end for
+7      $x_{i,p}\leftarrow$ Obtain standardized CA-FPC-scores;
+8       $q_{p}\leftarrow$ Number of CA-FPC-scores retained for sensor $p$;
+9
+10 end for
+/* MGR-ASGL based EM algorithm */
+[…]
+12 while $iter\leq iter_{max}$ do
+       $\gamma_{i}^{iter}\leftarrow$ Use $x_{i,p}$ in Eq. (7) for all $i$; ;
         // E-step
-       $\Pi^{iter}\leftarrow$ Update using Eq. (9) ;
+       $\Pi^{iter}\leftarrow$ Update using Eq. (9); ;
         // M-step for $\Pi$
-       $\boldsymbol{\rho}^{iter},\Psi^{iter}\leftarrow$ Solve Eq. (8) using $\sum_{p=1}^{P}\sqrt{q_{p}}$ and $\gamma_{i}^{iter}$;
-       $iter\leftarrow iter+1$;
-end while
-$\gamma_{i}^{iter_{max}}\leftarrow$ Final responsibilities from Eq. (7);
-$\boldsymbol{Z}_{i}[\operatorname*{arg\,max}_{k}{\gamma_{i}^{iter_{max}}}]\leftarrow 1$ for all $i$;
-$\boldsymbol{Z}_{i}[\sim\operatorname*{arg\,max}_{k}{\gamma_{i}^{iter_{max}}}]\leftarrow 0$ for all $i$;
-foreach $p$ and $k$ do
-       if $\lVert\psi^{iter_{max}}_{p,k}\rVert_{2}\geq\delta$ then
+13       $\boldsymbol{\rho}^{iter},\Psi^{iter}\leftarrow$ Solve Eq. (8) using $\sum_{p=1}^{P}\sqrt{q_{p}}$ and $\gamma_{i}^{iter}$;
+14       $iter\leftarrow iter+1$;
+15
+16 end while
+[…]
+18 $\boldsymbol{Z}_{i}[\operatorname*{arg\,max}_{k}{\gamma_{i}^{iter_{max}}}]\leftarrow 1$ for all $i$;
+19 $\boldsymbol{Z}_{i}[\sim\operatorname*{arg\,max}_{k}{\gamma_{i}^{iter_{max}}}]\leftarrow 0$ for all $i$;
+20 foreach $p$ and $k$ do
+21       if $\lVert\psi^{iter_{max}}_{p,k}\rVert_{2}\geq\delta$ then
              Append $p$ to $\mathcal{P}_{k}$ ;
               // Significant sensors for failure mode $k$
-       end if
-end foreach
+22            ;
+23
+24       end if
+25
+26 end foreach
 Algorithm 1 Offline Sensor Selection Algorithm
@@
 The parameter estimation equations of the EM algorithm are provided in III.B.5.
-IV-B3 Adaptive Sparse Group Lasso (ASGL) penalty
+4.2.3 Adaptive Sparse Group Lasso (ASGL) penalty
 The inclusion of features from non-informative sensors can negatively impact the accuracy of the RUL prediction.
 To address this issue, we penalize the negative-scaled conditional expectation of the CDLL function using the Adaptive Sparse Group Lasso (ASGL) penalty to obtain the following minimization problem given by 6.
-$\begin{split}\min_{\boldsymbol{\Theta}}-\frac{1}{N}\mathbb{E}_{\boldsymbol{\Theta}_{\text{old}}}[\ell_{C}(\boldsymbol{\Theta};&\boldsymbol{Y},\mathbb{Z})|\boldsymbol{Y}]+\lambda\sum_{k=1}^{K}\pi_{k}\sum_{p=1}^{P}\sqrt{q_{p}}\times\\ &\left[(1-\alpha)\left\|\boldsymbol{\psi}_{p,k}\right\|_{2}+\alpha\left\|\boldsymbol{\psi}_{p,k}\right\|_{1}\right]\end{split}$ (6)
+$\min_{\boldsymbol{\Theta}}-\frac{1}{N}\mathbb{E}_{\boldsymbol{\Theta}_{\text{old}}}[\ell_{C}(\boldsymbol{\Theta};\boldsymbol{Y},\mathbb{Z})|\boldsymbol{Y}]+\lambda\sum_{k=1}^{K}\pi_{k}\sum_{p=1}^{P}\sqrt{q_{p}}\times\left[(1-\alpha)\left\|\boldsymbol{\psi}_{p,k}\right\|_{2}+\alpha\left\|\boldsymbol{\psi}_{p,k}\right\|_{1}\right]$ (6)
 ASGL is an extension of lasso [39] that ensures the coefficients for all features of a non-informative group (sensor) are reduced to zero.
@@
 $\lambda>0$ and $\alpha\in[0,1]$ are the hyperparameters for this model.
-IV-B4 Expectation Maximization algorithm
+4.2.4 Expectation Maximization algorithm
 We now describe the EM algorithm used to find a local optimum to the IDLL described.
@@
 M-Step: During the M-step, we update the parameters by solving the optimization problem in (8) with respect to $\boldsymbol{\Theta}$.
-$\begin{split}\min_{\boldsymbol{\Theta}}\frac{1}{N}Q(\boldsymbol{\Theta}\mid&\boldsymbol{\Theta}_{\text{old}})+\lambda\sum_{k=1}^{K}\pi_{k}\sum_{p=1}^{P}\sqrt{q_{p}}\times\\ &\left[(1-\alpha)\left\|\boldsymbol{\psi}_{p,k}\right\|_{2}+\alpha\left\|\boldsymbol{\psi}_{p,k}\right\|_{1}\right]\end{split}\vspace*{-1em}$ (8)
+$\min_{\boldsymbol{\Theta}}\frac{1}{N}Q(\boldsymbol{\Theta}\mid\boldsymbol{\Theta}_{\text{old}})+\lambda\sum_{k=1}^{K}\pi_{k}\sum_{p=1}^{P}\sqrt{q_{p}}\times\left[(1-\alpha)\left\|\boldsymbol{\psi}_{p,k}\right\|_{2}+\alpha\left\|\boldsymbol{\psi}_{p,k}\right\|_{1}\right]$ (8)
 Improvement with respect to $\boldsymbol{\Pi}$: To update the mixing coefficients, we set $\rho_{k}=\rho_{k}^{\text{old}},\boldsymbol{\psi}_{0,k}=\boldsymbol{\psi}_{0,k}^{\text{old }},\boldsymbol{\psi}_{p,k}=\boldsymbol{\psi}_{p,k}^{\text{old }}$.
@@
 The detailed step-wise implementation of the offline sensor selection step is given in algorithm 1.
-V Online diagnosis and Prognostics
+5 Online diagnosis and Prognostics
 This section details the steps of the online part of our multi-failure mode prognostics framework.
@@
 Then, we demonstrate how we utilize MFPCA to i) diagnose the active failure mode of the degrading unit, in real-time, and ii) predict its RUL.
+1
 Input : Smoothed signals from training & test units, Failure labels of training data from Algorithm 1, Set of informative sensors $\mathcal{P}_{k}$ from Algorithm 1
 Output : Active failure mode in the test unit $k^{*}$, Remaining Useful Life (RUL) of the test unit $RUL_{test}$
+2
 /* Diagnostic Model Training */
-$\zeta_{(i),h}\leftarrow$ MFPCA on training units (Eq. 10);
-Train a KNN classifier using $\zeta_{(i),h}$;
+3 $\zeta_{(i),h}\leftarrow$ MFPCA on training units (Eq. 10);
+4 Train a KNN classifier using $\zeta_{(i),h}$;
+5
 /* Diagnostic Model Testing */
-$\zeta_{test,h}\leftarrow$ MFPCA on test unit;
-$k^{*}\leftarrow$ Mode of labels in KNN of $\zeta_{test,h}$;
+6 $\zeta_{test,h}\leftarrow$ MFPCA on test unit;
+7 $k^{*}\leftarrow$ Mode of labels in KNN of $\zeta_{test,h}$;
+8
 /* Prognostic Model Training */
-$\zeta^{k^{*}}_{(i),h}\leftarrow$ MFPCA: selected sensors (Eq. 14);
-$c_{0}^{k^{*}},\dots,c_{h}^{k^{*}}\leftarrow$ Obtain $\operatorname*{arg\,min}$ in Eq. 17;
+9 $\zeta^{k^{*}}_{(i),h}\leftarrow$ MFPCA: selected sensors (Eq. 14);
+10 $c_{0}^{k^{*}},\dots,c_{h}^{k^{*}}\leftarrow$ Obtain $\operatorname*{arg\,min}$ in Eq. 17;
+11
 /* Prognostic Model Testing */
-$\zeta^{k^{*}}_{test,h}\leftarrow$ MFPCA on test unit (Eq. 14);
-$RUL_{test}\leftarrow$ Equation 18;
+12 $\zeta^{k^{*}}_{test,h}\leftarrow$ MFPCA on test unit (Eq. 14);
+13 $RUL_{test}\leftarrow$ Equation 18;
 Algorithm 2 Online Diagnosis and Prognostics
-V-A Sensor Fusion: MFPCA
+5.1 Sensor Fusion: MFPCA
 Without loss of generality, let us redefine the sensor index as $\{1,2,\ldots,\mathcal{P}\}$, where $\mathcal{P}$ is number of sensors we seek to fuse.
@@
 […]
-V-B Failure Mode Diagnosis: K-nearest Neighbors
+5.2 Failure Mode Diagnosis: K-nearest Neighbors
 In order to diagnose the failure mode of a degrading unit at time $t^{*}$, we apply MFPCA on the $t^{*}$ training dataset.
@@
 Given this diagnosis, we now develop a model for predicting the unit RUL.
-V-C RUL Prediction: Weighted Functional Regression
+5.3 RUL Prediction: Weighted Functional Regression
 To predict RUL, we use a weighted time-varying functional regression model that maps multivariate signal features to the $\ln{TTF}$.
 First, we discuss the basics of a time-varying functional regression and then motivate the case for a weighted version of the model
-V-C1 Time-Varying Functional Regression
+5.3.1 Time-Varying Functional Regression
 When predicting RUL, we are often faced with updating the relationship between time-to-failure and the current values of the extracted signal features (FPC scores).
@@
 For variable selection, we include the lasso penalty in the optimization problem.
-$\begin{split}\min_{c_{0}^{k^{*}},c_{1}^{k^{*}},\ldots,c_{H_{k^{*}}}^{k^{*}}}\sum_{i\in\mathcal{I}_{t^{*}}^{k^{*}}}&\left(\sqrt{w_{i}}y_{i}-\sqrt{w_{i}}c_{0}^{k^{*}}-\sum_{h=1}^{H_{k^{*}}}\sqrt{w_{i}}\zeta_{i,h}^{k^{*}}c_{h}^{k^{*}}\right)^{2}\\ &+\lambda\sum_{h=1}^{H_{k^{*}}}\lVert c_{h}^{k^{*}}\rVert\end{split}\vspace*{-0.5em}$ (17)
-V-C2 RUL prediction
+$\min_{c_{0}^{k^{*}},c_{1}^{k^{*}},\ldots,c_{H_{k^{*}}}^{k^{*}}}\sum_{i\in\mathcal{I}_{t^{*}}^{k^{*}}}\left(\sqrt{w_{i}}y_{i}-\sqrt{w_{i}}c_{0}^{k^{*}}-\sum_{h=1}^{H_{k^{*}}}\sqrt{w_{i}}\zeta_{i,h}^{k^{*}}c_{h}^{k^{*}}\right)^{2}+\lambda\sum_{h=1}^{H_{k^{*}}}\lVert c_{h}^{k^{*}}\rVert$ (17)
+5.3.2 RUL prediction
 As we have shown, MFPCA enables the transformation of the functional regression problem to a multiple linear regression problem, which enables us to solve the problem only by estimating $H_{k^{*}}+1$ parameters.
@@
 ${RUL}_{test}=\exp\left(c_{0}^{k^{*}}+\sum_{k=1}^{H_{k^{*}}}\zeta_{test,h}^{k^{*}}c_{k}^{k^{*}}\right)-t^{*}$ (18)
-VI Case Study 1: Simulated Dataset
+6 Case Study 1: Simulated Dataset
 We present a case study using a simulated data set to evaluate the performance of our methodology in terms of its ability to cluster the training data, effectiveness in selecting informative sensors for each failure mode, and accuracy in predicting the RUL.
@@
 Finally, we evaluate the prediction accuracy of our online prognostics approach in section VI-C.
-VI-A Simulation Model
+6.1 Simulation Model
 We utilize a modified version of the simulation model used in [13] to generate degradation signals for a set of engineering systems.
@@
 Finally, for sensor 19, the signal decreases in response to both failure modes.
-Figure 2: Case study 1 data: Signals in sensors 1, 3, 5, 19 for both failure modes 1 and 2 under three different SNRs i.e., [2, 5], [5, 8], and [8, 11]
-VI-B Failure Mode Labeling and Sensor Selection
+Figure 2: Case study 1 data: Signals in sensors 1, 3, 5, 19 (columns 1, 2, 3, 4 respectively) for FM 1 and 2 under different SNRs – [2, 5], [5, 8], and [8, 11] (rows 1, 2, 3 respectively)
+6.2 Failure Mode Labeling and Sensor Selection
 We analyze the ability of the MGR-ASGL model to accurately cluster the degradation signals into one of the two possible failure modes along with selecting the subset of informative sensors.
-We start by utilizing CA-FPCA as mentioned in IV-A to extract features for each sensor separately.
+We start by utilizing CA-FPCA as mentioned in 4.1 to extract features for each sensor separately.
 To determine the number of CA-FPC scores to retain, we use the 95% fraction of variance explained (FVE) criterion.
@@
 We use minimum mean-squared error (minMSE) criterion, which selects the tuning parameters that minimize the mean-squared prediction error averaged over all folds.
-Table I displays the tuning parameter selection using both criteria over all three datasets.
+Table 1 displays the tuning parameter selection using both criteria over all three datasets.
 For each dataset and criterion, we fit the MGR-ASGL model using these tuning parameters.
-TABLE I: Tuning parameter selection for case study 1
+Table 1: Tuning parameter selection for case study 1
 $SNR_{p}^{k}$ [2,5] [5,8] [8,11]
@@
 Furthermore, we have predetermined which sensors are most informative and we would like to determine whether or not the MGR-ASGL selects these sensors.
-The results of this performance evaluation are shown in Table II.
-TABLE II: MGR-ASGL performance for case study 1
+The results of this performance evaluation are shown in Table 2.
+Table 2: MGR-ASGL performance for case study 1
 $SNR_{p}^{k}$ FM Accuracy Sensors Selected
@@
 However, simply reporting the sensors selected does not account for how informative they are.
-Table III reports the four sensors with the highest $l_{2}$-norm of the regression coefficients associated with that sensor.
+Table 3 reports the four sensors with the highest $l_{2}$-norm of the regression coefficients associated with that sensor.
 Except for the [5,8] dataset, at least three of the four sensors with the largest $l_{2}$-norm for each failure mode are elements of the set of informative sensors.
@@
 Given the selected sensors, we want to test the robustness of the online portion of the methodology to predict remaining life.
-TABLE III: $l_{2}$-norms of four most informative sensors for FM1 & FM2
+Table 3: $l_{2}$-norms of four most informative sensors for FM1 & FM2
 $SNR_{p}^{k}$ FM 1 FM 2
@@
 18 0.394 17 0.428
-VI-C RUL Prediction
+6.3 RUL Prediction
 Each of the datasets consist of 80 test systems (40 from each failure mode).
@@
 Since this dataset has less noise, the influence of clustering may be more pronounced compared to the noisier datasets, especially if training samples have been removed for failing before 90% of the monitored system’s lifetime.
-VII Case Study 2: NASA Turbofan Engine Dataset
+7 Case Study 2: NASA Turbofan Engine Dataset
 Figure 4: The system diagram used in Case Study 2 adapted from [44]
@@
 Both sets contain 100 observed engine failures caused by either fan degradation or high-pressure compressor degradation.
-Furthermore, both sets contain data from 21 sensors, which are listed in Table IV.
+Furthermore, both sets contain data from 21 sensors, which are listed in Table 4.
 The signals for these sensors are displayed in Figure 5.
@@
 Therefore, our goal is to use our methodology to predict the remaining life of each testing unit following the final cycle at which their degradation was observed.
-TABLE IV: Sensors used in case study 2 (C-MAPPS-3 [44])
+Table 4: Sensors used in case study 2 (C-MAPPS-3 [44])
 Sensor ID Sensor Name Functionality
@@
 The combination of tuning parameters that minimize the mean squared error across the three folds is $(\lambda_{MSE},\alpha_{MSE})=(0.1089,1)$.
-Table V displays the $l_{2}$-norms of the regression coefficients for each sensor.
+Table 5 displays the $l_{2}$-norms of the regression coefficients for each sensor.
 The sensors selected are those with positive $l_{2}$-norms.
-TABLE V: Sensor Selection for Case Study 2
+Table 5: Sensor Selection for Case Study 2
 Sensor ID FM 1 FM2 Sensor ID FM 1 FM 2
@@
 […]
-VIII Conclusion
+8 Conclusion
 In this paper, we present a methodology for predicting the RUL of systems with multiple failure modes using multiple sensors.

### v3

@@
-Sensor-fusion based Prognostics Framework for Complex Engineering Systems Exhibiting Multiple Failure Modes
+Sensor-fusion based Prognostics for Deep-space Habitats Exhibiting Multiple Unlabeled Failure Modes
 Abstract
-Complex engineering systems are often subject to multiple failure modes.
-Developing a remaining useful life (RUL) prediction model that does not consider the failure mode causing degradation is likely to result in inaccurate predictions.
-However, distinguishing between causes of failure without manually inspecting the system is nontrivial.
-This challenge is increased when the causes of historically observed failures are unknown.
-Sensors, which are useful for monitoring the state-of-health of systems, can also be used for distinguishing between multiple failure modes as the presence of multiple failure modes results in discriminatory behavior of the sensor signals.
-When systems are equipped with multiple sensors, some sensors may exhibit behavior correlated with degradation, while other sensors do not.
-Furthermore, which sensors exhibit this behavior may differ for each failure mode.
-In this paper, we present a simultaneous clustering and sensor selection approach for unlabeled training datasets of systems exhibiting multiple failure modes.
-The cluster assignments and the selected sensors are then utilized in real-time to first diagnose the active failure mode and then to predict the system RUL.
-We validate the methodology using a simulated dataset of systems exhibiting two failure modes and on NASA turbofan degradation dataset.
+Deep-space habitats are complex systems that must operate autonomously over extended durations without ground-based maintenance.
+These systems are vulnerable to multiple, often unknown, failure modes that affect different subsystems and sensors in mode-specific ways.
+Developing accurate remaining useful life (RUL) prognostics is challenging, especially when failure labels are unavailable and sensor relevance varies by failure mode.
+In this paper, we propose an unsupervised prognostics framework that jointly identifies latent failure modes and selects informative sensors using only unlabeled training data.
+The methodology consists of two phases.
+In the offline phase, we model system failure times using a mixture of Gaussian regressions and apply a novel Expectation-Maximization algorithm to cluster degradation trajectories and select mode-specific sensors.
+In the online phase, we extract low-dimensional features from the selected sensors to diagnose the active failure mode and predict RUL using a weighted regression model.
+We demonstrate the effectiveness of our approach on a simulated dataset that reflects deep-space telemetry characteristics and on a real-world engine degradation dataset, showing improved accuracy and interpretability over existing methods.
 keywords:
-Multiple Failure Modes , Remaining Useful Life , Sensor Fusion , Functional Principal Component Analysis , Mixture of Gaussian Regression , Adaptive Sparse Group Lasso , Expectation Maximization
-journal: Reliability Engineering & System Safety
+Deep Space Habitats , Multiple Failure Modes , Remaining Useful Life , Sensor Fusion , Failure Diagnosis , Prognostics
+journal: Acta Astronautica
 […]
 […]
-{highlights}
-We propose a two-stage prognostic framework for multi-sensor systems with multiple failure modes.
-Offline: An unsupervised method labels failure events and ranks sensor importance.
-Online: Sensor features are fused for failure mode diagnosis, then key sensors predict RUL.
-Our method achieves $80\%$ clustering accuracy and $<3\%$ RUL prediction error in simulations.
-In NASA turbofan data, early predictions outperform benchmarks while reducing input by $30\%$.
 1 Introduction
-Prognostic models aim to estimate the remaining useful life (RUL) of critical components and systems.
-They are important because they drive condition-based maintenance decisions that aim to increase equipment availability, reduce downtime, and prevent unexpected failures.
-Generally, prognostic models can be classified into two categories: (1) data-driven models that focus on modeling sensor data and assume that the data is directly (or indirectly) correlated with the underlying physical degradation process, and (2) physics-based models that rely on first principles and require a thorough understanding of the underlying physics driving the degradation process.
-While physics-based approaches are known for their high accuracy, data-driven approaches are growing in popularity, especially for modeling complex systems.
-Several reasons are driving this trend.
-First, complex systems often exhibit complex physics of failure that are difficult to characterize using first principles.
-The second reason is the rise of Machine Learning and Artificial Intelligence models, which when trained properly can be very powerful prognostic tools.
-However, many data-driven prognostics models assume that complex systems exhibit a single predominant failure mode that can be used to accurately predict the system’s RUL.
-In practice, complex engineering systems often exhibit multiple failure modes and thus require more sophisticated modeling approaches to account for the different ways/modes a system can fail.
-In this paper, we focus on developing a prognostic methodology for complex systems monitored by multiple sensors that exhibit multiple failure modes.
-We design an integrated framework comprised of (1) a sensor selection algorithm, (2) a dimensionality reduction and data fusion methodology, and (3) a RUL prediction model.
-Our methodology relies on a historical dataset to simultaneously diagnose observed failure events and select the subset of the most informative sensors associated with each failure mode.
-One of the unique aspects of our methodology relates to the sensor selection model, which uses an unsupervised approach to identify the dominant failure modes.
-Unlike most of the existing works, we assume that the labels of the failure events are unknown a priori.
-This assumption introduces some unique challenges around model estimation.
-These challenges are not trivial, especially when they are considered in conjunction with sensor selection.
-Our methodology is comprised of two parts, an offline component and an online prognostic model.
-In the offline part, we develop a novel algorithm based on the Expectation-Maximization (EM) algorithm that simultaneously labels the failures in the historical training set and selects the optimal sensors for each failure mode.
-In the online part, we utilize real-time data to perform online diagnosis of the predominant failure mode and RUL prediction.
-During the online learning phase, Multivariate Functional Principle Component Analysis (MFPCA) is utilized for feature extraction and dimensionality reduction.
-We then use the nearest neighbor classifier to diagnose the active failure mode in the real-time signals before fitting a weighted functional regression model using the extracted features.
-The main contributions of this paper are threefold:
-  • First, we propose a feature extraction methodology comprised of fusing signals from multiple sensors originating from a set of systems exhibiting multiple failure modes.
-  • Second, we propose an optimization methodology for simultaneous failure mode labeling and sensor selection for any given unlabeled historical dataset.
-We call this as the “offline sensor selection” step.
-  • Finally, we provide a two-step sequential approach to leverage this knowledge of failure labels and informative sensors to (1) perform active failure mode diagnosis and (2) subsequent RUL prediction on systems degrading in real-time.
-The paper mentions this as the “online diagnosis and prognostics” step.
-The remainder of the paper is outlined as follows.
-In Section 2, we review the relevant literature associated with this research area.
-In Section III, we outline our proposed multi-failure mode prognostics framework.
-In Section 4, we present our offline sensor selection scheme that utilizes an EM algorithm.
-In Section 5, we present our online diagnostics and prognostics method.
-It discusses a strategy to fuse informative sensors for real-time diagnosis of failure modes and subsequent prediction of RUL.
-Our methodology is demonstrated through two case studies – one using a simulated dataset in section 6 and the other NASA turbofan engine dataset in section 7.
-We conclude by summarizing our framework and key findings in Section 8.
+A deep-space habitat is a complex engineering system (CES), composed of multiple interacting subsystems that must work together to support human life and mission operations over long durations.
+These subsystems such as the environmental control and life support system, energy and power systems, thermal control, robotic agents, and structural components etc., must function reliably in harsh and isolated environments, where failure of any component can lead to mission-critical consequences [1, 2, 3, 4].
+The complexity of such systems arises not only from the tight integration of these subsystems but also from their heterogeneous origins.
+Components are often developed by different manufacturers and tested independently, making it difficult to anticipate how they will behave once integrated.
+Additionally, the deep-space environment introduces operating conditions such as microgravity, radiation, and thermal extremes that differ significantly from terrestrial test settings.
+These novel conditions can give rise to failure modes that have never been observed during isolated ground-based validation.
+As a result, building a comprehensive library of failure modes becomes essential for predicting degradation behaviors (i.e., prognostics) during the mission.
+However, due to limited diagnostic access and the lack of labeled failure data, this library must often be inferred from unlabeled system observations.
+In addition to the challenge of missing labels, deep-space habitats face another difficulty that limits their prognostic capability.
+These systems are typically equipped with a large number of sensors that generate in situ telemetry data to continuously monitor system health.
+However, analyzing this data presents several issues.
+Not all sensors are informative for predicting remaining useful life (RUL), and many may be redundant or affected by noise.
+Furthermore, in systems with multiple failure modes, sensor informativeness is often mode-specific.
+A sensor that is strongly predictive under one failure mode may be uninformative under another (see Figure 1).
+This variability complicates the task of selecting a reliable subset of sensors for prognostic modeling, especially with unknown failure modes.
+In this paper, we propose a data-driven prognostics framework for deep-space systems that are monitored by many sensors and subject to multiple, unknown failure modes.
+Our methodology consists of two key stages: (1) an offline sensor selection step and (2) an online diagnosis and prognostics step.
+These stages align with two mission stages: during early deployment, the habitat operates under limited supervision to collect historical sensor data for initializing the model; after this setup phase, autonomous health monitoring and prediction begin.
+Figure 1: Example of a deep space habitat with multiple subsytems.
+Health of each subsystem is monitored by a subset of onboard sensors.
+The subset of informative sensors vary for different failure modes.
+Main Contributions.
+A unique aspect of our approach is that we do not assume failure mode labels are known in the historical training data (degradation signals).
+This reflects realistic space mission constraints, where diagnosing past failures requires expert involvement and post-mission analysis.
+To address this, we model failure times using a mixture of Gaussian regressions, where regressors are features extracted from sensor data.
+We develop a novel Expectation-Maximization (EM) algorithm that simultaneously labels failure modes and selects the most informative sensors for each mode.
+In the online phase, we use real-time sensor data to first diagnose the dominant failure mode before predicting RUL.
+We apply Multivariate Functional Principal Component Analysis (MFPCA) to extract low-dimensional, yet informative features from the mode-specific sensors identified in the offline phase.
+We then classify the active failure mode using a nearest-neighbor approach and apply a weighted functional regression model to predict RUL.
+The key technical contributions of this paper can be summarized as:
+  • 1.
+    We develop a feature extraction methodology that fuses multivariate sensor data from systems exhibiting multiple failure modes into compact and informative representations.
+  • 2.
+    We propose a failure-mode-aware sensor selection approach using an Expectation-Maximization algorithm that jointly clusters unlabeled failure events and selects informative sensors for each failure mode.
+  • 3.
+    We present an integrated online framework that uses real-time data to (1) diagnose the active failure mode and (2) predict RUL using a mode-specific regression model.
+  • 4.
+    We demonstrate the effectiveness and robustness of the proposed methodology through two case studies relevant to deep-space habitats: (1) a controlled simulation designed to reflect key telemetry challenges, and (2) a real-world NASA turbofan engine degradation dataset that mimics sensor-driven subsystem degradation in long-duration missions.
+The remainder of the paper is organized as follows.
+Section 2 reviews related work on RUL prediction and multi-failure mode diagnostics.
+Section 3 introduces our proposed prognostics framework.
+Section 4 describes the offline sensor selection method.
+Section 5 details the online diagnosis and RUL prediction approach.
+Section 6 and Section 7 present two case studies—a simulated dataset and the NASA turbofan engine dataset, respectively.
+We conclude the paper in Section 8 with a discussion of future work.
 2 Literature Review
-Data-driven prognostic models utilize a variety of approaches.
-A significant number of these rely on Artificial Intelligence tools, such as neural networks to predict component lifetime [1, 2].
-Some approaches have focused on directly modeling degradation-based signals that are derived from raw sensor data.
-Examples include random coefficient models [3], relevance vector machines [4], stochastic processes such as the Brownian motion process [5], the Gamma process [6], and hidden Markov models [7], and similarity-based prognostic models [8].
-Most of these models have been developed for modeling a single sensor or degradation-based signal.
-Extending these approaches to applications involving multiple sensors is not straightforward.
-Fusion algorithms have been leveraged to model multivariate sensor data.
-According to [9], fusion is usually performed at the decision level or the data level.
-Decision-based fusion involves combining the results of multiple prediction models in an ensemble fashion [10].
-Conversely, data fusion methodologies focus on aggregating data from multiple sensors to develop a single (univariate) health index, which can then be modeled using the conventional methods discussed above.
-Most of the works in this area are characterized by their data fusion approach.
-Examples include principal component analysis (PCA) [11, 12, 13], maximum likelihood [14], deep learning [15], and state space models [16, 17, 18].
-Most data fusion methodologies tend to aggregate all the sensor data.
-In practice, some sensors may not be correlated with the underlying failure process.
-Thus, integrating signals from these sensors into a data fusion algorithm introduces noise and may impact RUL prediction accuracy.
-To address this issue, several works have incorporated sensor selection algorithms.
-For example, [14] and [9] employed adaptive lasso to penalize non-informative sensors.
-[13] combined location-scale regression with group regularization technique to perform sensor selection.
-While these techniques can select an optimal subset of sensors for monitoring, they still assume that system failure is caused by a single predominant failure mode.
-Most complex engineering systems can fail due to multiple failure modes.
-Competing risk models have been a popular approach for modeling multiple failure modes.
-Traditionally, competing risk models rely on statistical approaches.
-For example, [19] uses the Cox proportional hazard model (PHM) [20] with the Weibull baseline hazard rate to develop a mixture of Weibull PHM.
-For an extensive review of statistical competing risk models, we refer the reader to [21].
-Deep learning has also been used to develop competing risk models.
-These models include convolutional neural network (CNN) [22], recurrent neural networks [23], and LSTM [24].
-Recently, [25] developed a deep learning methodology to model complex survival problems, such as recurrent hospitalizations due to kidney failure or pancreas graft.
-For more examples, we refer the reader to [26], which provides a review of machine learning methodologies for survival analysis.
-Competing risk models often output survival/hazard rates that are sometimes difficult to interpret in engineering applications.
-Consequently, many researchers have adopted methods that explicitly model sensor data to predict RUL.
-The literature that focuses explicitly on RUL prediction of systems with multiple failure models can be loosely divided into two groups.
-The first group uses deep learning approaches while the second group utilizes more statistical methods.
-Deep learning models such as [27] and [28] develop feature extraction methodologies based on convolutional neural networks (CNN) for modeling systems with multiple failure modes.
-[29] developed a Gaussian model whose mean is a Bayesian deep neural network that maps sensor data to the RUL.
-In [30], the authors used a support vector machine to diagnose the active failure mode using features extracted from a deep belief network.
-Following diagnosis, a particle filter is employed to predict RUL.
-[31] trained an LSTM-based RUL prediction model for all present failure modes.
-The choice of prediction model is based on a deep CNN-based classifier trained with physics-informed failure labels.
-While deep learning approaches have been successful at predicting RUL, their parameters are often difficult to interpret.
-Additionally, they require significantly larger datasets for training compared to statistical methods.
-In [32], the authors derived a composite health index for each failure mode and used a random coefficients-based polynomial degradation model to predict the system’s RUL.
-In their model training/estimation, the authors assumed that the labels of the failure modes were given.
-In fact, this assumption has been widely used in the literature and is often supported by visual inspection of the data [27, 30, 32] or available domain knowledge [31].
-[33] assumes that a subset of the data is labeled and utilizes a similarity-based approach for classifying the rest of the unlabeled failure modes in the training set.
-Our methodology differs from previous works in that we consider a scenario where the failure labels are completely unknown.
-Furthermore, we assume that for each failure mode, there exists a (not necessarily disjoint) subset of informative sensors that are useful for prognostics.
-Our objective is to design an unsupervised prognostic methodology that synthesizes signals from multiple sensors in a systematic fashion.
-We assume the availability of a historical repository of system failures whose degradation was monitored from the moment an incipient fault occurred up to system failure using multiple sensors.
-The methodology assumes that all systems fail independently of each other.
-In addition, we assume that only one type of failure mode can occur in the entire lifetime of the system.
-While the failure labels are unknown, the total number of potential failure modes for the system is assumed to be known.
+Approaches for RUL prediction of CES are typically categorized as either physics-based or data-driven.
+Physics-based approaches [5, 6] utilize first principles to mathematically model the underlying physics of system degradation.
+For example, [5] used a two-stage nonlinear differential equation to model the rate of air pressure changes inside a space habitat.
+They use an Extended Kalman Filter (EKF) to estimate the parameters of this model and predict RUL by using the same EKF to forecast when the habitat pressure becomes untenable.
+In general, physics-based approaches typically develop state-space models of system degradation and use a Bayesian filter such as the Kalman Filter or the Particle Filter for real-time estimation and prediction of model parameters.
+Several examples of this are highlighted by [7].
+The advantages of physics-based approaches include high prediction accuracy and limited need for data.
+However, their requirement for thorough knowledge of the physics behind system failure renders them difficult to develop, given the growing system complexity.
+Conversely, data-driven approaches attempt to map sensor data to the system RUL using statistical models, machine learning, or hybrid approaches.
+Since sensors are monitoring the system over a prolonged period, they generate continuous, high-volume data called “streaming data.” The approaches developed for single-sensor streaming data include both statistical methods that model univariate degradation signals using exponential models [8, 9], random-coefficient regression [10], Brownian motion [11], Gamma process [12], or hidden Markov model [13] and machine learning methods that use deep learning [14, 15] or relevance vector machine [16].
+However, single-sensor approaches are not scalable to the scenario where multiple sensors monitor systems.
+Under this scenario, sensors produce multivariate streaming data.
+For this type of data, authors have developed sensor fusion algorithms that operate at either the decision level or the data level [17].
+Decision-level fusion consists of systematically combining RUL estimates from different models, whereas data-level fusion synthesizes the multivariate sensor data into a univariate health indicator, which can be modeled using standard univariate techniques.
+Existing data fusion approaches include Principal Component Analysis [18, 19, 20], maximum likelihood methods [21], extreme learning machine [22], deep learning [23, 24], state-space models [25, 26, 27], and logistic regression [28].
+While effective, many data fusion methods rely on aggregating all available sensor signals.
+In practice, not all sensors are informative for a given failure mechanism, and fusing irrelevant signals can introduce noise and reduce prediction accuracy.
+To address this, some studies incorporate sensor selection.
+For instance, adaptive lasso has been used to eliminate uninformative sensors [17, 21], and group regularization techniques have been applied for joint selection and modeling [20].
+However, these approaches typically assume a single dominant failure mode.
+In reality, most CES can fail through multiple failure modes, reflecting the structure and function of their subsystems.
+Competing risk models have been widely used to represent such systems.
+Traditional approaches often rely on statistical models, such as the Cox proportional hazards model with Weibull baselines [29, 30]; a comprehensive review is provided in [31].
+More recently, deep learning has also been used to model competing risks [32].
+For example, both [33] and [34] both employ deep recurrent neural networks for recurrent event survival analysis with competing risks.
+Competing risk models often produce survival or hazard rates that can be difficult to interpret in engineering contexts.
+As a result, many researchers have shifted toward methods that directly model sensor data to predict RUL, though few focus on multiple failure modes.
+It is common for researchers to develop general RUL prediction methodologies and then apply them to datasets with multiple failure modes.
+For example, [35] proposed a correlation and relative entropy feature engineering framework for complex systems and [36] proposed a deep convolutional neural network.
+In both cases, the researchers validated their methodologies on datasets with multiple failure modes.
+However, they do not consider multiple failure modes in an explicit manner.
+Failure to do so could result in highly biased predictions.
+Therefore, the development of RUL prediction methodologies that account for multiple failure modes is growing.
+[37] used a support vector machine for failure mode diagnosis with features extracted from a deep belief network, followed by RUL prediction using a particle filter.
+[38] trained distinct LSTM models for each failure mode.
+Once a physics-informed CNN classifier diagnoses the active failure mode, the appropriate LSTM model is utilized.
+Furthermore, [39] fit random coefficient models to failure mode-specific health indices.
+However, these articles do not incorporate a sensor selection algorithm and they rely on labeling of the failure modes determined either through visual inspection of domain knowledge.
+Consequently, researchers have begun addressing these gaps.
+[40] adopted a semi-supervised graph-based approach for feature extraction from partially labeled failures and used elastic net functional regression for sensor selection and RUL prediction.
+More recently, [41] proposed a time-series clustering algorithm to identify latent failure modes, followed by LSTM-based RUL prediction.
+Furthermore, [42] combined mixture of (log)-location scale regression with deep learning to predict RUL in a scenario with unknown failure modes.
+However, in both papers, sensor selection was limited to excluding sensors with obviously non-informative signals, such as constant or low-variance signals.
+Our methodology differs from prior work by addressing a setting in which failure mode labels are entirely unknown.
+We assume that each failure mode is associated with a (not necessarily disjoint) subset of sensors that are informative for RUL prediction.
+The goal is to develop an unsupervised prognostic framework that systematically fuses multivariate sensor signals for RUL estimation.
+We assume access to a historical repository of system run-to-failure trajectories, where degradation was monitored from the onset of an incipient fault to final failure using multiple sensors.
+Furthermore, we assume that each system fails independently and experiences only one failure mode during its lifetime.
+Finally, while the failure labels are unobserved, we assume that the total number of potential failure modes is known.
+In the following sections, we detail our prognostics framework.
+Table 1: Comparison of Prognostic Methodologies in the Literature
+Prognostic Sensor Sensor Multiple Unlabeled Model Methodology Fusion Selection Failure Modes Failure Modes Interpretability [5, 6, 7] N/A [8, 9, 10, 11, 12, 13] [18, 19, 20, 21, 23] Partial [21, 20, 43] [39, 44, 37, 38] Partial [40] Partial Partial [41, 42] Partial Partial [32, 45, 33, 34, 46] [44, 35, 47, 36] Ours
 3 Overview of the Prognostics Framework
-This section describes our multiple failure mode prognostics framework, which is summarized in the flowchart in Figure 1. This framework is comprised of two parts:
-Offline Sensor Selection: The first part is performed offline and primarily focuses on utilizing a training set of historically observed degradation signals to perform sensor selection.
-Sensor selection is important because not all sensors are informative for prognostics, and the inclusion of non-informative sensors in a prognostics algorithm can compromise the accuracy of RUL predictions.
-When performing sensor selection, it is computationally convenient to represent each sensor by a parsimonious set of features that capture the system dynamics.
-Therefore we further divide this step into the following substeps:
+Figure 2: Overview of the proposed prognostics framework for deep-space habitats
+This section describes our prognostics framework, developed for autonomous health monitoring in deep space habitats.
+A flowchart of the framework is shown in Figure 2.
+The framework consists of two main components: “Offline Sensor Selection” and “Online Diagnosis and Prognostics”.
+I.
+Offline Sensor Selection: This stage occurs shortly after the habitat is deployed, during an initial period where degradation signals are collected with limited ground or crew assistance.
+The data collected during this stage forms a historical dataset, which is used to initialize the prognostic models before autonomous operation begins.
+We assume that the dataset contains multiple failure modes, although the failure modes are not labeled.
+The goal of this step is to (1) label the degradation signals into one of the failure modes, and (2) identify a subset of informative sensors for each failure mode.
+This step consists of two parts discussed below:
   1. 1.
-    Feature Extraction: It involves a variant of the traditional Functional Principal Component Analysis (FPCA) framework to account for multiple failure modes in the unlabeled data.
+    Feature Extraction: The onboard sensors generate high-dimensional, time-varying signals.
+To extract informative features for prognostics, we use a covariate-adjusted functional principal component analysis (CA-FPCA) method.
+CA-FPCA models the variation in each sensor signal while accounting for external covariates.
+In our setting, these covariates represent the underlying failure modes, which are unknown apriori.
+To estimate them, we first perform a K-means clustering on the sensor signals and use the resulting cluster labels as initial covariates.
   2. 2.
-    Optimization: This involves formulating an optimization model to simultaneously label the failure modes and select the informative sensors for each failure mode.
-Feature Extraction: There are many ways to extract parsimonious features from complex and multivariate data.
-In our framework, we choose to use a Functional Principal Component Analysis (FPCA) framework due to its flexibility and extensibility to various types of functional data, allowing us to efficiently capture the underlying patterns and variability in complex, high-dimensional datasets.
-FPCA models the variability of temporal data observations around a mean function.
-However, the presence of multiple failure modes can cause heterogeneity in the magnitude and/or sign of the signals’ rate of change, compromising the effectiveness of FPCA as a prognostic tool.
-Therefore, the first step of our offline sensor selection algorithm is to extract features using a Covariate-Adjusted FPCA (CA-FPCA) approach [34].
-CA-FPCA is an extension of conventional FPCA that allows modeling functional data while accounting for the effects of explanatory variables, known as covariates.
-In our context, the covariates are used to represent the failure modes.
-One of the unique contributions of our approach is the assumption that the failure modes are not known a priori.
-However, CA-FPCA assumes that covariates are observable.
-We note that the way each sensor responds to different failure modes can be highly variable.
-For example, a sensor may be more sensitive to one failure mode than another resulting in heterogeneity in the rate of change of its corresponding sensor signals.
-Furthermore, a sensor signal may increase in magnitude in response to one failure mode, and decrease for another.
-To address the lack of observable covariates, we first perform regular FPCA on the signals from each sensor.
-Then, we perform K-means clustering on the FPC scores for each sensor individually, using the resulting cluster labels as the covariate values to perform CA-FPCA.
-Optimization: After feature extraction the sensor signals are transformed into parsimonious CA-FPC-scores, which serve as predictors for a Mixture of Gaussian (MGR) model used to model the natural logarithm of the time-to-failure, $\ln{TTF}$, as a function of the CA-FPC-scores.
-This model is motivated by the assumption that the functional relationship between the $\ln{TTF}$ and the CA-FPC scores is dependent on the latent failure mode.
-We fit the parameters of the MGR model using the Expectation-Maximization (EM) algorithm.
-In the E-step, we obtain a soft clustering of the ((((CA-FPC-scores, $\ln{TTF})$ tuples into their respective failure modes.
-In the M-step, we incorporate the Adaptive Sparse Group Lasso (ASGL) penalty into the regression model to remove groups of features associated with non-informative sensors.
-From now on, we refer to this model as MGR-ASGL.
-The output of the offline sensor sensor selection step of our methodology are (1) failure mode labels for each sample in the training set and (2) a subset of informative sensors for each failure mode.
-Online diagnosis and Prognostics: The second part of our methodology consists of an online step used to process and analyze data from units operating in the field.
-The online step consists of diagnosing the active failure mode of the fielded unit and then predicting the RUL of the fielded unit.
-In the diagnostic step, features are extracted from the real-time data generated by all selected sensors using Multivariate FPCA (MFPCA), which is an extension of FPCA to multivariate functional data.
-We use MFPCA to generate a parsimonious set of fused signal features from multivariate sensor signals; hereafter referred to as MFPC-scores.
-Then, we classify the active failure mode of the fielded unit by assigning to the unit the failure mode most frequently represented by its K-nearest neighbors.
-Once this step is complete, we recalculate the MFPC scores using only the informative sensors associated with the diagnosed failure mode of the unit.
-In the prognostic step, we fit a functional regression model using the MFPC scores that predict the RUL of the unit.
-In the following sections, we formalize our prognostics framework.
-Figure 1: Flowchart of the multiple failure mode prognostics framework
-4 Offline Sensor Selection Algorithm
-We consider a scenario where a system is monitored by $P$ sensors.
-We assume there exists a historical dataset comprised of $P$-dimensional degradation-based sensor signals from monitoring $N$ similar systems.
+    Optimization: After feature extraction, we obtain a reduced set of features for each sensor.
+These features are called CA-FPC scores.
+We use these scores as predictors in a mixture of Gaussian regression (MGR) model.
+The MGR model estimates the natural logarithm of the time-to-failure, $\ln{TTF}$, as a function of the CA-FPC scores.
+The model assumes that this relationship ($\ln{TTF}$ vs CA-FPC scores) depends on the underlying failure mode.
+We fit the MGR model using the EM algorithm, which results in: (1) optimal failure mode labels for each sample, and (2) an optimal subset of informative sensors for each failure mode.
+II.
+Online Diagnosis and Prognostics: Once, the failure mode labeling and sensor selection are complete, the habitat enters into autonomous mode.
+This stage uses real-time sensor data to (1) identify the active failure mode and (2) predict the RUL:
+  1. 1.
+    Diagnosis: Prior to the diagnosis step, the active failure mode is unknown therefore, all selected sensors are utilized in this step.
+We apply multivariate functional principal component analysis (MFPCA) to signals from all selected sensors to extract a compact set of features, called MFPC scores, that represent the joint behavior of all sensors.
+We classify the active failure mode by finding its K-nearest neighbors and assigning the most common failure mode among them.
+  2. 2.
+    Prognostics: Once the active failure mode is diagnosed, we recalculate the MFPC scores using only the informative sensors for that mode.
+A functional regression model uses these scores to predict the RUL.
+4 Offline Sensor Selection
+We consider a scenario where the deep space habitat is monitored by $P$ different sensors.
+We assume there exists a historical dataset comprised of $P$-dimensional degradation-based sensor signals from monitoring $N$ similar samples.
 We denote $s_{i,p}(t)$ as the signal data from sensor $p\in\{1,...,P\}$ for system $i\in\{1,...,N\}$ observed at time $t\in[0,T]$, where $T=min(\{TTF_{i}\}_{i=1}^{N})$.
@@
 Once the sensor-wise clustering is complete, the next step involves feature extraction by applying CA-FPCA and then regressing these features against the time-to-failure to identify the most informative subset of sensors for each failure mode.
-The next two subsections discuss details about CA-FPCA and the penalized Mixture of Gaussian regression model (MGR-ASGL).
+The next two subsections discuss details about the feature extraction (CA-FPCA), and the optimization (MGR-ASGL) model.
 4.1 Feature Extraction: CA-FPCA
@@
 Using Mercer’s theorem [49], the covariance can be decomposed as follows, $\mathbb{C}_{p}^{k}(t,t^{\prime})=\sum_{m=1}^{\infty}\lambda_{m,p}^{k}\phi_{m,p}^{k}(t)\phi_{m,p}^{k}(t^{\prime})$, where $\{\phi_{m,p}^{k}(t)\}_{m=1}^{\infty}$ are the orthogonal eigenfunctions of $\mathbb{C}_{p}^{k}(t,t^{\prime})$ and $\lambda_{1,p}^{k}\geq...\geq\lambda_{m,p}^{k}\geq...$ are the corresponding eigenvalues.
-By projecting the mean-subtracted signal data onto the eigenfunctions, we can represent the signals by equation (2).
+By projecting the mean-subtracted signal data onto the eigenfunctions, we can represent the signals by Eq (2).
 $s_{i,p}(t)=\mu_{p}^{k}(t)+\sum_{m=1}^{\infty}\xi_{i,m,p}^{k}\phi_{m,p}^{k}(t)+\epsilon_{i,p}(t)\vspace*{-1em}$ (2)
@@
 Since the eigenvalues decrease as $m\to\infty$, we can obtain a low-dimensional representation of $s_{i,p}^{k}(t)$ using the first $q_{p}$ CA-FPC-scores.
-To ensure all features are scaled equally, we use the standardized versions of the CA-FPC-scores: $x_{i,m}^{p}=\frac{\sum_{k=1}^{K}\xi_{i,m,p}^{k}\mathit{\mathbb{I}}(i\in\mathcal{I}_{p}^{k})}{\sqrt{\frac{1}{N-1}\sum_{k=1}^{r_{p}^{*}}N_{p}^{k}\lambda_{m,p}^{k}}}$, where $N_{p}^{k}$ is the number of signals assigned to cluster $k$ for sensor $p$.
+To ensure all features are scaled equally, we standardize the CA-FPC-scores by subtracting the sample mean and dividing by the sample standard deviation.
+Let $x_{i,m}^{p}$ denote the standardized $m$th CA-FPC-score from sensor $p$ of system $i$.
 We extract $\boldsymbol{x}_{i,p}=(x_{i,1}^{p},...,x_{i,q_{p}}^{p})^{T}$ for sensor $p$ of system $i$,.
-We then utilize this $\ boldsymbol{x}_{i,p}$’s as predictors for fitting the MGR-ASGL model.
+We then utilize the $\boldsymbol{x}_{i,p}$’s as predictors for fitting the MGR-ASGL model.
 4.2 Optimization of the MGR-ASGL model
-To perform sensor selection, we seek to model the distribution of $\ln{TTF}$ conditioned on the observed set of features.
-In what follows, we note that values of $\ln{TTF}$ are standardized.
-Our modeling approach assumes that the failure times are conditionally independent given the observed features.
-Let $y_{i}$ denote the random variable corresponding to the $\ln{TTF}$ of the system $i$.
-We consider $\boldsymbol{Z}_{i}$ to denote a “one-hot” random vector of length $K$, where $\boldsymbol{Z}_{i}[k]=1$ if system $i$ failed due to failure mode $k$, and zero otherwise.
-For mathematical brevity, we let $\boldsymbol{Y}=\left(y_{1},\ldots,y_{N}\right)$ denote the observed $\ln{TTF}$ and $\mathbb{Z}=\left(\boldsymbol{Z}_{1},\boldsymbol{Z}_{2},\ldots,\boldsymbol{Z}_{N}\right)$ denote the set of random vectors representing the failure modes for each system.
-In the following subsections, we first assign probability distributions to these random variables to formalize the MGR model.
-Second, we construct the log-likelihood function for the MGR model.
-Third, we provide theoretical intuition on the use of the EM algorithm for model fitting.
-Fourth, we describe the ASGL penalty.
-Finally, we provide details for implementing the EM algorithm.
-4.2.1 Mixture of Gaussian Regression (MGR) model
-We assume that the distribution of $y_{i}$ conditioned on the observed features $\boldsymbol{x}_{i,1},\ldots,\boldsymbol{x}_{i,P}$} and $\boldsymbol{Z}_{i}[k]=1$ is Gaussian.
-If we let $\pi_{1},\ldots,\pi_{K}$, where $\sum_{k=1}^{K}\pi_{k}=1$, denote the mixing coefficients, which represent the probability that $\boldsymbol{Z}_{i}[k]=1$ ($\forall k=\{1,\ldots,K\}$), then equation (3) gives the marginal distribution of $y_{i}$ given only $\{\boldsymbol{x}_{i,1},\ldots,\boldsymbol{x}_{i,p}\}$.
-This marginal distribution is the MGR model.
-$f_{\vartheta}(y_{i}\mid\boldsymbol{x}_{i,1},\ldots,\boldsymbol{x}_{i,P})=\sum_{k=1}^{K}\pi_{k}\frac{1}{\sqrt{2\pi}\sigma_{k}}\times\ \exp\left(-\frac{1}{2\sigma_{k}^{2}}\left(y_{i}-\beta_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\beta}_{p,k}\right)^{2}\right)$ (3)
-4.2.2 Log-Likelihood
-[…]
-Definition 1.
-The log-likelihood function for the MGR model is given by equation (4).
-This function is called as the ”incomplete-data log-likelihood” (IDLL), since we do not know the cause of the observed failure in the data.
-$\ell(\boldsymbol{\Theta};\boldsymbol{Y})=\sum_{i=1}^{N}\ln\sum_{k=1}^{K}\pi_{k}\times\ \frac{\rho_{k}}{\sqrt{2\pi}}\exp\left(-\frac{1}{2}\left(\rho_{k}y_{i}-\psi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\psi}_{p,k}\right)^{2}\right)$ (4)
-Definition 2.
-The “complete-data log-likelihood” (CDLL) given by equation (5) assumes knowledge of the underlying causes of the observed failure in the data.
-$\ell_{C}(\boldsymbol{\Theta};\boldsymbol{Y},\mathbb{Z})=\sum_{i=1}^{N}\sum_{k=1}^{K}\boldsymbol{Z}_{i}[k]\ln\pi_{k}\left(\frac{\rho_{k}}{\sqrt{2\pi}}\right)\exp\left(-\frac{1}{2}\left(\rho_{k}y_{i}-\psi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\psi}_{p,k}\right)^{2}\right)$ (5)
-The IDLL given in equation (4) is difficult to optimize globally [38].
-Instead, we search for a local optimum by using EM algorithm, an iterative algorithm suitable for fitting probability models with latent variables.
-To derive the steps of the EM algorithm, we note that since $\ln(.)$ is a concave function We use Jensen’s inequality to form a lower bound on the IDLL.
-This lower bound is the expectation of the CDLL given in equation (5) and shown in Lemma 1.
-Input : Number of systems $N$, Number of sensors $P$, Unlabeled data from $P$ sensors of all $N$ systems, Maximum iterations $iter_{max}$ for EM algorithm, threshold $\delta$ for regression coefficient’s $l_{2}$-norm
-Output : Vector $\boldsymbol{Z}$ indicating failure labels, Set of informative sensors $\mathcal{P}_{k}$ for each failure mode $k$
-/* Feature Extraction from all $P$ sensors (CA-FPCA) */
-1 for $p=1$ to $P$ do
-2       $\mathcal{I}_{k}^{p}\leftarrow$ K-means clustering on FPC-scores;
-3       for $k=1$ to $K$ do
-4             $\xi_{i,m,p}(k)\leftarrow$ Use $\mathcal{I}_{k}^{p}$ to obtain CA-FPC-scores (Eq. 2);
-5
-6       end for
-7      $x_{i,p}\leftarrow$ Obtain standardized CA-FPC-scores;
-8       $q_{p}\leftarrow$ Number of CA-FPC-scores retained for sensor $p$;
-9
-10 end for
-/* MGR-ASGL based EM algorithm */
-[…]
-12 while $iter\leq iter_{max}$ do
-       $\gamma_{i}^{iter}\leftarrow$ Use $x_{i,p}$ in Eq. (7) for all $i$; ;
-        // E-step
-       $\Pi^{iter}\leftarrow$ Update using Eq. (9); ;
-        // M-step for $\Pi$
-13       $\boldsymbol{\rho}^{iter},\Psi^{iter}\leftarrow$ Solve Eq. (8) using $\sum_{p=1}^{P}\sqrt{q_{p}}$ and $\gamma_{i}^{iter}$;
-14       $iter\leftarrow iter+1$;
-15
-16 end while
-[…]
-18 $\boldsymbol{Z}_{i}[\operatorname*{arg\,max}_{k}{\gamma_{i}^{iter_{max}}}]\leftarrow 1$ for all $i$;
-19 $\boldsymbol{Z}_{i}[\sim\operatorname*{arg\,max}_{k}{\gamma_{i}^{iter_{max}}}]\leftarrow 0$ for all $i$;
-20 foreach $p$ and $k$ do
-21       if $\lVert\psi^{iter_{max}}_{p,k}\rVert_{2}\geq\delta$ then
-             Append $p$ to $\mathcal{P}_{k}$ ;
-              // Significant sensors for failure mode $k$
-22            ;
-23
-24       end if
-25
-26 end foreach
-Algorithm 1 Offline Sensor Selection Algorithm
-Lemma 1.
-The expectation of the CDLL w.r.t distribution $g$ is a lower bound for the IDLL i.e.,
-$\mathbb{E}_{\boldsymbol{g}}\left[\ell_{C}(\boldsymbol{\Theta};\boldsymbol{Y},\mathbb{Z})\right]\leq\ell(\boldsymbol{\Theta};\boldsymbol{Y})$
-Proof.
-Let $g$ be an arbitrary distribution on $\boldsymbol{Z}_{i}[k]$. Then we can re-write equation (4) as follows:
-$\ell(\boldsymbol{\Theta};\boldsymbol{Y})=\sum_{i=1}^{N}\ln\left(\sum_{k=1}^{K}\frac{g\left(\boldsymbol{Z}_{i}[k]\right)p\left(y_{i},\boldsymbol{Z}_{i}[k];\boldsymbol{\Theta}\right)}{g\left(\boldsymbol{Z}_{i}[k]\right)}\right)\vspace*{-1em}$
+Let $Y_{i}=\frac{\ln{TTF}_{i}-(1/N)\sum_{i=1}^{N}\ln{TTF}_{i}}{(1/(N-1))\sum_{i=1}^{N}\big{(}\ln{TTF}_{i}-(1/N)\sum_{i=1}^{N}\ln{TTF}_{i}\big{)}^{2}}$, for $i=1,2…,N$ denote the random variable corresponding to the standardized natural logarithm of the time-to-failure (i.e., $\ln{TTF}$) of system $i$.
+Furthermore, let $\boldsymbol{Z}_{i}$ denote a categorical random vector of size $K$ used to encode the unknown failure mode of system $i$.
+If failure mode $k$ is responsible for the failure of system $i$, which occurs with probability $\pi_{k}$, then $\boldsymbol{Z}_{i}[k]=1$ and the remaining elements are zero.
+In our model, we posit that the relationship between $Y_{i}$ and the observed features $x_{i}$ depends on which failure mode caused the failure of system $i$.
+Thus, we assume that the conditional distribution of $Y_{i}$ given $\boldsymbol{Z}_{i}[k]=1$ is Gaussian with PDF:
+$f_{Y_{i}|\boldsymbol{Z}_{i}}(y_{i}|\boldsymbol{Z}_{i}[k]=1)=\frac{1}{\sqrt{2\pi}\sigma_{k}}\exp\Big{(}-\frac{1}{2\sigma_{k}^{2}}(y_{i}-\beta_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\beta}_{p,k})^{2}\Big{)}$ (3)
+To find the marginal distribution of $Y_{i}$ we use the law of total probability, summing over the joint distribution of $Y_{i}$ and $\boldsymbol{Z}_{i}$ as follows:
+$f_{Y_{i}}(y_{i})=\sum_{k=1}^{K}\pi_{k}\frac{1}{\sqrt{2\pi}\sigma_{k}}\exp\Big{(}-\frac{1}{2\sigma_{k}^{2}}(y_{i}-\beta_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\beta}_{p,k})^{2}\Big{)}$ (4)
+Therefore, the marginal distribution of $Y_{i}$ is a mixture of Gaussian regressions.
+The parameters of this model are $\Theta=\{\pi_{k},\beta_{0,k},\boldsymbol{\beta}_{1,k},\dots,\boldsymbol{\beta}_{P,k},\sigma_{k}\}_{k=1}^{K}$.
+To estimate these parameters, we seek to minimize the negative incomplete-data log-likelihood (IDLL).
+To ensure both scale-invariance of parameter estimation and convexity of the optimization problem in the M-step of our EM algorithm, we let $\varphi_{0,k}=\sigma_{k}^{-1}\beta_{0,k}$, $\boldsymbol{\varphi}_{p,k}=\sigma_{k}^{-1}\boldsymbol{\beta}_{p,k}$, and $\rho_{k}=\sigma_{k}^{-1}$ (see: [41]).
+Thus, the parameters are now $\Theta=\{\boldsymbol{\pi},\boldsymbol{\rho},\boldsymbol{\varphi}\}$, where $\boldsymbol{\pi}=\{\pi_{k}\}_{k=1}^{K}$, $\boldsymbol{\rho}=\{\rho_{k}\}_{k=1}^{K}$, and $\boldsymbol{\varphi}=\{\varphi_{0,k},\boldsymbol{\varphi}_{1,k},\dots,\boldsymbol{\varphi}_{P,k}\}_{k=1}^{K}$.
+Assuming that $Y_{1},\dots,Y_{N}$ are independent, the negative IDLL is computed as follows:
+$-\ell(\Theta|\boldsymbol{Y})=-\sum_{i=1}^{N}\ln\Big{[}\sum_{k=1}^{K}\pi_{k}\rho_{k}\frac{1}{\sqrt{2\pi}}\exp\Big{(}-\frac{1}{2}(y_{i}\rho_{k}-\varphi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,k})^{2}\Big{)}\Big{]}$ (5)
+Estimating $\Theta$ requires minimizing Eq 5.
+However, the logarithm inside the summation makes Eq 5 difficult to optimize globally.
+Instead, we search for a local minimum using the EM algorithm, an iterative algorithm suitable for fitting probability models with latent variables.
+To utilize the EM algorithm, we first construct the negative complete-data log likelihood (CDLL).
+$\begin{split}-\ell_{C}(\Theta|\boldsymbol{Y},\mathbb{Z})&=-\ln\prod_{i=1}^{N}\prod_{k=1}^{K}f_{Y_{i}}(y_{i})^{\boldsymbol{Z}_{i}[k]}=-\sum_{i=1}^{N}\sum_{k=1}^{K}\boldsymbol{Z}_{i}[k]\ln\Big{[}\pi_{k}\rho_{k}\frac{1}{\sqrt{2\pi}}\\ &\times\exp\Big{(}-\frac{1}{2}(y_{i}\rho_{k}-\varphi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,k})^{2}\Big{)}\Big{]}\end{split}$ (6)
+In Appendix A, we show that the expectation of the negative CDLL with respect to the conditional distribution of $\mathbb{Z}$ given $\boldsymbol{Y}$ forms an upper bound for the negative IDLL.
+Therefore, we can obtain a local minimum by minimizing this upper bound.
+The proposed EM algorithm is performed as follows:
+  1. 1.
+    Initialization: Set the initial parameters to $\Theta^{(0)}$ s.t.,
+    $\Theta^{(0)}=\{\pi_{k}^{(0)},\varphi_{0,k}^{(0)},\boldsymbol{\varphi}_{1,k}^{(0)},\dots,\boldsymbol{\varphi}_{P,k}^{(0)},\rho_{k}^{(0)}\}_{k=1}^{K}$ (7)
+  2. 2.
+    E-step: Given $\Theta^{(m)}$, the expectation of $-\ell_{C}(\Theta|\boldsymbol{Y},\mathbb{Z})$ is given as:
+    $\begin{split}Q(\Theta,\Theta^{(m)})&=\mathbb{E}[-\ell_{C}(\Theta|\boldsymbol{Y},\mathbb{Z})|\boldsymbol{Y},\Theta^{(m)}]=-\sum_{i=1}^{N}\sum_{k=1}^{K}\gamma_{i,k}\ln\Big{[}\pi_{k}\rho_{k}\frac{1}{\sqrt{2\pi}}\\ &\times\exp\Big{(}-\frac{1}{2}(y_{i}\rho_{k}-\varphi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,k})^{2}\Big{)}\Big{]}\end{split}$ (8)
     where,
-$\begin{split}&p\left(y_{i},\boldsymbol{Z}_{i}[k];\boldsymbol{\Theta}\right)\\ &=\pi_{k}\cdot\left(\frac{\rho_{k}}{\sqrt{2\pi}}\exp\left(-\frac{1}{2}\left(\rho_{k}y_{i}-\psi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\psi}_{p,k}\right)^{2}\right)\right)\end{split}$
-Since $\ln(.)$ is concave, from Jensen’s inequality we can write that,
-$\ell(\boldsymbol{\Theta};\boldsymbol{Y})\geq\sum_{i=1}^{N}\sum_{k=1}^{K}g\left(\boldsymbol{Z}_{i}[k]\right)\ln{\frac{p\left(y_{i},\boldsymbol{Z}_{i}[k];\boldsymbol{\Theta}\right)}{g\left(\boldsymbol{Z}_{i}[k]\right)}}\vspace*{-1em}$
-and thus,
-$\ell(\boldsymbol{\Theta};\boldsymbol{Y})\geq\mathbb{E}_{\boldsymbol{g}}\left[\ell_{C}(\boldsymbol{\Theta};\boldsymbol{Y},\mathbb{Z})\right]-\sum_{i=1}^{N}\sum_{k=1}^{K}g\left(\boldsymbol{Z}_{i}[k]\right)\ln{g\left(\boldsymbol{Z}_{i}[k]\right)}\vspace*{-1em}$
-The second term on the right side of the inequality is the negative entropy of the $g$ distribution. Since it is positive, the inequality holds. ∎
-For every iteration of the EM algorithm, the E-step calculates the conditional expectation of the CDLL given the observed failure times, which we denote as $\mathbb{E}_{\boldsymbol{\Theta}_{\text{old}}}[\ell_{C}(\boldsymbol{\Theta};\boldsymbol{Y},\mathbb{Z})|\boldsymbol{Y}]$.
-In the M-step, we find the parameters that maximize the conditional expectation (see equation (8)).
-The parameter estimation equations of the EM algorithm are provided in III.B.5.
-4.2.3 Adaptive Sparse Group Lasso (ASGL) penalty
-The inclusion of features from non-informative sensors can negatively impact the accuracy of the RUL prediction.
-To address this issue, we penalize the negative-scaled conditional expectation of the CDLL function using the Adaptive Sparse Group Lasso (ASGL) penalty to obtain the following minimization problem given by 6.
-$\min_{\boldsymbol{\Theta}}-\frac{1}{N}\mathbb{E}_{\boldsymbol{\Theta}_{\text{old}}}[\ell_{C}(\boldsymbol{\Theta};\boldsymbol{Y},\mathbb{Z})|\boldsymbol{Y}]+\lambda\sum_{k=1}^{K}\pi_{k}\sum_{p=1}^{P}\sqrt{q_{p}}\times\left[(1-\alpha)\left\|\boldsymbol{\psi}_{p,k}\right\|_{2}+\alpha\left\|\boldsymbol{\psi}_{p,k}\right\|_{1}\right]$ (6)
-ASGL is an extension of lasso [39] that ensures the coefficients for all features of a non-informative group (sensor) are reduced to zero.
-In the context of regularization, adaptive refers to the penalty accounting for the number of CA-FPC-scores retained for each sensor i.e., $q_{p}$, and the mixing coefficient for each failure mode i.e., $\pi_{k}$.
-While the inclusion of $q_{p}$ ensures a sensor is not retained simply by having more coefficients, $\pi_{k}$ accounts for the class imbalance.
-$\lambda>0$ and $\alpha\in[0,1]$ are the hyperparameters for this model.
-4.2.4 Expectation Maximization algorithm
-We now describe the EM algorithm used to find a local optimum to the IDLL described.
-Given an initial estimate of the model parameters $\boldsymbol{\Theta}_{old}$, we perform the E-step and M-step as follows:
-E-step: The E-step consists of computing the expectation of CDLL in (6) conditioned on $\boldsymbol{Y}$.
-Let $\mathcal{Q}\left(\boldsymbol{\Theta}\mid\boldsymbol{\Theta}_{\text{old}}\right)=-\mathbb{E}_{\boldsymbol{\Theta}_{\text{old}}}\left[\ell_{C}(\boldsymbol{\Theta};\boldsymbol{Y},\mathbb{Z})\mid\boldsymbol{Y}\right]$ denote the negative expectation of the CDLL.
-In order to compute $\mathcal{Q}\left(\boldsymbol{\Theta}\mid\boldsymbol{\Theta}_{\text{old}}\right)$, we need to compute $\gamma_{i,k}=$ $\mathbb{E}_{\boldsymbol{\Theta}_{\text{old }}}\left[Z_{i}[k]\mid\boldsymbol{Y}\right]$, which represents the responsibility of failure mode $k$ to system $i$ and is given by equation (7).
-For any system $i$, we denote the vector of the responsibilites $\gamma_{i,k}$ across all failure modes as $\boldsymbol{\gamma}_{i}$ such that, $\boldsymbol{\gamma}_{i}=\{\gamma_{i,1},...,\gamma_{i,K}\}$.
-$\footnotesize\gamma_{i,k}=\frac{\pi_{k}^{\text{old }}\rho_{k}^{\text{old }}\times\exp\left(-\frac{1}{2}\left(\rho_{k}^{\text{old }}y_{i}-\psi_{0,k}^{\text{old }}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\psi}_{p,k}^{\text{old }}\right)^{2}\right)}{\sum_{l=1}^{K}\pi_{l}^{\text{old }}\rho_{l}^{\text{old }}\exp\left(-\frac{1}{2}\left(\rho_{l}^{\text{old }}y_{i}-\psi_{0,l}^{\text{old }}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\psi}_{p,l}^{\text{old }}\right)^{2}\right)}\vspace*{-0.5em}$ (7)
-M-Step: During the M-step, we update the parameters by solving the optimization problem in (8) with respect to $\boldsymbol{\Theta}$.
-$\min_{\boldsymbol{\Theta}}\frac{1}{N}Q(\boldsymbol{\Theta}\mid\boldsymbol{\Theta}_{\text{old}})+\lambda\sum_{k=1}^{K}\pi_{k}\sum_{p=1}^{P}\sqrt{q_{p}}\times\left[(1-\alpha)\left\|\boldsymbol{\psi}_{p,k}\right\|_{2}+\alpha\left\|\boldsymbol{\psi}_{p,k}\right\|_{1}\right]$ (8)
-Improvement with respect to $\boldsymbol{\Pi}$: To update the mixing coefficients, we set $\rho_{k}=\rho_{k}^{\text{old}},\boldsymbol{\psi}_{0,k}=\boldsymbol{\psi}_{0,k}^{\text{old }},\boldsymbol{\psi}_{p,k}=\boldsymbol{\psi}_{p,k}^{\text{old }}$.
-and solve optimization problem (8).
-We utilize the update strategy from [37].
-We start by identifying a feasible solution $\overline{\boldsymbol{\Pi}}^{\text{new}}=\left(\frac{1}{N}\sum_{i=1}^{N}\gamma_{i,1},\ldots,\frac{1}{N}\sum_{i=1}^{N}\gamma_{i,K}\right)^{T}$ and then update the mixing coefficients as described by equation (9), where $t^{\text{new }}\in(0,1]$ is chosen to be the largest value in $\left\{0.1^{k};k=0,1,2,\ldots\right\}$ such that the objective function does not increase.
-$\boldsymbol{\Pi}^{\text{new}}=\boldsymbol{\Pi}^{\text{old}}+t^{\text{new}}\left(\overline{\boldsymbol{\Pi}}^{\text{new }}-\boldsymbol{\Pi}^{\text{old}}\right)\vspace*{-0.5em}$ (9)
-Improvement with respect to $\boldsymbol{\rho}$ and $\boldsymbol{\Psi}$: For this, we utilize the updated mixing coefficients $\Pi^{\text{new }}$ of the previous step.
-Since the $\rho_{k}$’s and $\boldsymbol{\psi}_{p,k}$’s are independent for each failure mode, the expression in (8) decouples into $K$ distinct optimization problems.
-Once the estimators $\rho_{k}^{\text{new }},\psi_{0,k}^{\text{new }}$, and $\boldsymbol{\psi}_{p,k}^{\text{new }}$ are determined, they are used to update the cluster responsibilities in the E-step.
-After the EM algorithm is completed, we obtain optimal $\boldsymbol{\Pi}$, $\boldsymbol{\rho}$, and, $\boldsymbol{\Psi}$.
-We substitute these optimal values to equation (7) to obtain the responsibilities $\boldsymbol{\gamma}_{i}$’s.
-For each pair of sensor data and failure time, we assign the failure mode with the largest responsibility.
-This results in class labels for the training set.
-To perform sensor selection for failure mode $k$, we compute $\left\|\boldsymbol{\psi}_{p,k}\right\|_{2}$, the $l_{2}$-norm of the regression coefficients for sensor $p$ and failure mode $k$.
-We deem a sensor informative for failure mode $k$ if the $l_{2}$-norm of its regression coefficients is greater than 0.001.
-The detailed step-wise implementation of the offline sensor selection step is given in algorithm 1.
+    $\begin{split}\gamma_{i,k}&=\mathbb{E}(\boldsymbol{Z}_{i}[k]|Y_{i}=y_{i},\Theta^{(m)})=\Pr(\boldsymbol{Z}_{i}[k]=1|Y_{i}=y_{i},\Theta^{(m)})\\ &=\frac{\pi_{k}^{(m)}\rho_{k}^{(m)}\frac{1}{\sqrt{2\pi}}\exp\Big{(}-\frac{1}{2}(y_{i}\rho_{k}^{(m)}-\varphi_{0,k}^{(m)}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,k}^{(m)})^{2}\Big{)}}{\sum_{l=1}^{K}\pi_{l}^{(m)}\rho_{l}^{(m)}\frac{1}{\sqrt{2\pi}}\exp\Big{(}-\frac{1}{2}(y_{i}\rho_{l}^{(m)}-\varphi_{0,l}^{(m)}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,l}^{(m)})^{2}\Big{)}}\end{split}$ (9)
+    is the responsibility that failure mode $k$ takes for the failure of system $i$.
+These responsibilities enable a soft clustering, where we claim system $i$ failed due to failure mode $k$ if $\gamma_{i,k}>\gamma_{i,l}$ for $l\neq k$.
+  3. 3.
+    M-Step: After computing the responsibilities $\gamma_{i,k}$ , we can update our estimate of $\Theta$ by maximizing $Q(\Theta,\Theta^{(m)})$ with respect to $\Theta$.
+To perform sensor selection, we augment $Q(\Theta,\Theta^{(m)})$ with the adaptive sparse group lasso penalty and form the following optimization problem:
+    $\min_{\Theta}\Big{\{}Q(\Theta,\Theta^{(m)})+\lambda\sum_{k=1}^{K}\pi_{k}\Big{(}\alpha\sum_{p=1}^{P}\|\boldsymbol{\varphi}_{p,k}\|_{1}+(1-\alpha)\sum_{p=1}^{P}\sqrt{q_{p}}\|\boldsymbol{\varphi}_{p,k}\|_{2}\Big{)}\Big{\}}$ (10)
+    We refer to the problem in Eq 10 as Mixture of Gaussian Regression–Adaptive Sparse Group Lasso (MGR-ASGL).
+Because $\boldsymbol{\pi}$ is included in the penalty, we follow [51] and use a two-step optimization:
+    • (a)
+      Step 1: Improvement with respect to $\boldsymbol{\pi}$
+      We start by using the responsibilities from the E-step to identify a feasible solution $\bar{\boldsymbol{\pi}}^{(m+1)}=\big{(}\frac{1}{N}\sum_{i=1}^{N}\gamma_{i,1},\dots,\frac{1}{N}\sum_{i=1}^{N}\gamma_{i,K}\big{)}^{T}$.
+Then, we update the mixing coefficients as follows:
+      $\boldsymbol{\pi}^{(m+1)}=\boldsymbol{\pi}^{(m)}+u_{m}(\bar{\boldsymbol{\pi}}^{(m+1)}-\boldsymbol{\pi}^{(m)})$ (11)
+      where $u_{m}\in(0,1]$ is chosen to be the largest value in the sequence $\{1,\frac{1}{2},\frac{1}{4},\dots\}$ such that the objective function does not increase.
+    • (b)
+      Step 2: Improvement with respect to $\boldsymbol{\rho}$ and $\boldsymbol{\varphi}$
+      Given the updated mixing coefficients, we can optimize Eq 10 with respect to the other parameters.
+It can be shown that this problem can be decomposed into solving $K$ individual problems, one for each failure mode.
+Therefore, for $k=1,\dots,K$, we solve
+      $\begin{split}\min_{\rho_{k},\varphi_{0,k},\boldsymbol{\varphi}_{1,k},\dots,\boldsymbol{\varphi}_{P,k}}&\Big{\{}-\sum_{i=1}^{N}\gamma_{i,k}\big{[}\ln\pi_{k}^{(m+1)}+\ln\rho_{k}-\frac{1}{2}(y_{i}\rho_{k}-\varphi_{0,k}\\ &-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,k})^{2}\big{]}\ +\lambda\pi_{k}^{(m+1)}\Big{(}\alpha\sum_{p=1}^{P}\|\boldsymbol{\varphi}_{p,k}\|_{1}\\ &+(1-\alpha)\sum_{p=1}^{P}\sqrt{q_{p}}\|\boldsymbol{\varphi}_{p,k}\|_{2}\Big{)}\Big{\}}\end{split}$ (12)
+      Note that the tuning parameters $\lambda$ and $\alpha$ do not change for different $k$.
+These parameters can be selected using cross-validation.
+When $\lambda$ increases, ASGL drives groups of regression coefficients to zero.
+Since these groups correspond to features extracted from individual sensors, this, in effect, causes the influence of some sensors to vanish.
+Therefore, a sensor is considered significant for failure mode $k$ if $\|\varphi_{p,k}\|_{2}\neq 0$.
+The parameter $\alpha$ allows for some of the coefficients in a group to drop to zero to improve prediction accuracy.
+Finally, incorporating the adaptive parameters $q_{p}$ and $\gamma_{i,k}$ into the penalty ensures that sensor selection is not affected by class imbalance and the dimensionality of the features.
+Let $\boldsymbol{\rho}^{(m+1)}$ and $\boldsymbol{\varphi}^{(m+1)}$ denote the solution to Eq 12.
+Thus, after completing Steps 1 and 2, we have our updated solution $\Theta^{(m+1)}=\{\boldsymbol{\pi}^{(m+1)},\boldsymbol{\rho}^{(m+1)},\boldsymbol{\varphi}^{(m+1)}\}$.
+    Set $\Theta^{(m+1)}\leftarrow\Theta$ and return to (1). Perform successive iterations of (1)(3) until convergence.
 5 Online Diagnosis and Prognostics
-This section details the steps of the online part of our multi-failure mode prognostics framework.
-In this part, a fielded unit is degrading in real-time, and our goal is to predict the time remaining before the unit fails.
-We begin by showing how MFPCA is used to fuse the signals from multiple sensors into a parsimonious set of features.
-Then, we demonstrate how we utilize MFPCA to i) diagnose the active failure mode of the degrading unit, in real-time, and ii) predict its RUL.
-1
-Input : Smoothed signals from training & test units, Failure labels of training data from Algorithm 1, Set of informative sensors $\mathcal{P}_{k}$ from Algorithm 1
-Output : Active failure mode in the test unit $k^{*}$, Remaining Useful Life (RUL) of the test unit $RUL_{test}$
-2
-/* Diagnostic Model Training */
-3 $\zeta_{(i),h}\leftarrow$ MFPCA on training units (Eq. 10);
-4 Train a KNN classifier using $\zeta_{(i),h}$;
-5
-/* Diagnostic Model Testing */
-6 $\zeta_{test,h}\leftarrow$ MFPCA on test unit;
-7 $k^{*}\leftarrow$ Mode of labels in KNN of $\zeta_{test,h}$;
-8
-/* Prognostic Model Training */
-9 $\zeta^{k^{*}}_{(i),h}\leftarrow$ MFPCA: selected sensors (Eq. 14);
-10 $c_{0}^{k^{*}},\dots,c_{h}^{k^{*}}\leftarrow$ Obtain $\operatorname*{arg\,min}$ in Eq. 17;
-11
-/* Prognostic Model Testing */
-12 $\zeta^{k^{*}}_{test,h}\leftarrow$ MFPCA on test unit (Eq. 14);
-13 $RUL_{test}\leftarrow$ Equation 18;
-Algorithm 2 Online Diagnosis and Prognostics
+In this section, a subsystem(s) in deep-space habitat is considered to be degrading in real-time.
+Our goal is to predict the time remaining before the unit fails.
+We begin by showing how to fuse the signals from multiple sensors into a parsimonious set of features.
+Then, we demonstrate how we utilize the fused sensor information to i) diagnose the active failure mode of the degrading unit, in real-time, and ii) predict its RUL.
 5.1 Sensor Fusion: MFPCA
 Without loss of generality, let us redefine the sensor index as $\{1,2,\ldots,\mathcal{P}\}$, where $\mathcal{P}$ is number of sensors we seek to fuse.
-Now assume that we have monitored an online unit (operating in the field) for a duration given by the time interval $[0,t^{*}]$.
+Now assume that we have monitored the degrading subsystem(s) for a duration given by the time interval $[0,t^{*}]$.
 Since we are interested in predicting the remaining lifetime, we only use signals from units with failure times greater than $t^{*}$.
@@
 This method also identifies outliers and assigns them weights of zero to improve signal mean estimation.
-To perform MFPCA, we model the $i$th smoothed multivariate signal evaluated at time $t$ as $\boldsymbol{s}_{i}(t)=\boldsymbol{u}_{i}(t)+\boldsymbol{\epsilon}_{i}(t)$, where $\boldsymbol{u}_{i}(t)\in\mathcal{R}^{\mathcal{P}}$ is a random multivariate function with mean $\boldsymbol{\mu}(t)$ and covariance function $\mathcal{C}(t,t^{\prime})$, $\boldsymbol{\epsilon}_{i}(t)\in\mathcal{R}^{\mathcal{P}}$, and $t\in[0,t^{*}]$ are i.i.d.
-zero-mean error vectors with covariance $\Sigma_{\boldsymbol{\epsilon}}$.
-We assume $\boldsymbol{u}_{i}(t)$ and $\boldsymbol{\epsilon}_{i}(t)$ are independent of each other.
-Using the first $H$ multivariate eigenfunctions, we can represent the multivariate signal by equation 10 where the MFPC-scores $\zeta$’s are computed according to equation 11.
+To perform MFPCA, we model the $i$th smoothed multivariate signal evaluated at time $t\in[0,t^{*}]$ as having mean $\boldsymbol{\mu}(t)\in\mathcal{R}^{\mathcal{P}}$ and $\mathcal{P}\times\mathcal{P}$ block covariance function $\mathcal{C}(t,t^{\prime})$, where $\mathcal{C}_{j,j^{\prime}}(t,t^{\prime})=cov(v_{j}(t),v_{j^{\prime}}(t^{\prime}))$.
+Using Mercer’s theorem [49], we can represent the covariance function as $\mathcal{C}(t,t^{\prime})=\sum_{h=1}^{\infty}\eta_{h}\boldsymbol{\varpi}_{h}(t)\boldsymbol{\varpi}_{h}(t^{\prime})^{T}$, where $(\eta_{h},\boldsymbol{\varpi}_{h}(t))$ are the $h$th eigenvalue-eigenfunction pair and $\eta_{1}\geq\eta_{2}\geq\dots$.
+Using the first $H$ multivariate eigenfunctions, we can represent the multivariate signal by Eq 13 where the MFPC-scores $\zeta$’s are computed according to Eq 14.
 $\boldsymbol{s}_{i}(t)=\boldsymbol{\mu}(t)+\sum_{h=1}^{H}\zeta_{i,h}\boldsymbol{\varpi}_{h}(t)+\boldsymbol{\epsilon}_{i}(t),\vspace*{-1.2em}$ (13)
 […]
+In practice, we perform MFPCA by collecting the $N_{t^{*}}$ observations, each consisting of the $\mathcal{P}$ sensor signals concatenated into a single vector.
+Then we perform traditional PCA on the sample of observations.
+Furthermore, we standardize the extracted MFPC-scores so they have zero mean and unit variance.
 5.2 Failure Mode Diagnosis: K-nearest Neighbors
 In order to diagnose the failure mode of a degrading unit at time $t^{*}$, we apply MFPCA on the $t^{*}$ training dataset.
-Next, $K$ nearest neighbors (KNN) is applied to the extracted MFPC scores to identify the active failure mode.
-Here, the distance metric of the KNN classifier is defined using the Euclidean distance between the MFPC scores.
+Next, $K$ nearest neighbors (KNN) is applied to the standardized MFPC-scores to identify the active failure mode.
+Here, the distance metric of the KNN classifier is defined using the Euclidean distance between the MFPC-scores.
 Given this diagnosis, we now develop a model for predicting the unit RUL.
@@
 5.3.1 Time-Varying Functional Regression
-When predicting RUL, we are often faced with updating the relationship between time-to-failure and the current values of the extracted signal features (FPC scores).
-The predictor trajectories up to the current time are represented by time-varying FPC scores, which are continuously updated as time progresses and are considered to be time-varying predictor variables for the functional regression model.
+When predicting RUL, we are often faced with updating the relationship between time-to-failure and the current values of the extracted signal features (MFPC-scores).
+The predictor trajectories up to the current time are represented by time-varying MFPC-scores, which are continuously updated as time progresses and are considered to be time-varying predictor variables for the functional regression model.
 This is referred to as time-varying functional regression.
@@
 Furthermore, let $\mathcal{P}_{k^{*}}\subset\{1,2,...,\mathcal{P}\}$ denote the set of sensors selected (identified as most informative) for failure mode $k^{*}$.
-To predict RUL, we fit a functional regression model using least squares as shown in equation (12), where $y_{i}=\ln{TTF_{i}}$, $\varphi_{0}^{k^{*}}$ is the bias term, $\boldsymbol{\varphi}^{k^{*}}(t)\in\mathcal{R}^{|\mathcal{P}_{k^{*}}|}$ is the multivariate coefficient function, and $\boldsymbol{s}_{i}^{k^{*}}(t)$ is the multivariate signal consisting only of sensors significant to failure mode $k^{*}$.
+To predict RUL, we fit a functional regression model as shown in Eq (15), where $y_{i}=\ln{TTF_{i}}$, $\varphi_{0}^{k^{*}}$ is the bias term, $\boldsymbol{\varphi}^{k^{*}}(t)\in\mathcal{R}^{|\mathcal{P}_{k^{*}}|}$ is the multivariate coefficient function, and $\boldsymbol{s}_{i}^{k^{*}}(t)$ is the multivariate signal consisting only of sensors significant to failure mode $k^{*}$.
 […]
 We expand the coefficient function using the eigenfunctions obtained from applying MFPCA on the training set $\mathcal{I}_{t^{*}}^{k^{*}}$.
-The regression coefficient function can now be expanded as shown in equation (13), where $H_{k^{*}}$ denotes the number of MFPC-scores that were retained for failure mode $k^{*}$.
+The regression coefficient function can now be expanded as shown in Eq (16), where $H_{k^{*}}$ denotes the number of MFPC-scores that were retained for failure mode $k^{*}$.
 $\boldsymbol{\varphi}(t)=\sum_{h=1}^{H_{k^{*}}}c_{h}^{k^{*}}\boldsymbol{\varpi}_{k}^{k^{*}}(t)\vspace*{-0.5em}$ (16)
-Likewise, we expand the multivariate signals using the eigenfunctions as shown in equation (14).
+Likewise, we expand the multivariate signals using the eigenfunctions as shown in Eq (17).
 $\boldsymbol{s}_{i}^{k^{*}}(t)=\boldsymbol{\mu}^{k^{*}}(t)+\sum_{h=1}^{H_{k^{*}}}\zeta_{i,h}^{k^{*}}\boldsymbol{\varpi}_{h}^{k^{*}}(t)\vspace*{-0.5em}$ (17)
-Since the eigenfunctions are orthogonal, substituting equations (13) and (14) into equation (12) yields the following optimization problem,
+Since the eigenfunctions are orthogonal, substituting Eqs (16) and (17) into Eq (15) yields the following optimization problem,
 $\min_{c_{0}^{k^{*}},c_{1}^{k^{*}},\ldots,c_{H_{k^{*}}}^{k^{*}}}\sum_{i\in\mathcal{I}_{t^{*}}^{k^{*}}}\left(y_{i}-c_{0}^{k^{*}}-\sum_{h=1}^{H_{k^{*}}}\zeta_{i,h}^{k^{*}}c_{h}^{k^{*}}\right)^{2}\vspace*{-1em}$ (18)
@@
 Weighted Version: We note that the offline sensor selection algorithm may generate some outliers, which can impact the accuracy of the regression problem.
-These outliers become more apparent after identifying the active failure mode and the MFPC scores to estimate the time-varying functional regression model.
-To address this issue, we use weighted least squares to estimate the time-varying functional regression mode.
+These outliers become more apparent after identifying the active failure mode and the MFPC-scores to estimate the time-varying functional regression model.
+To address this issue, we use weighted least squares to estimate the time-varying functional regression model.
 To derive the weights, we first measure the Euclidean distance between the MFPC score corresponding to a unit $i$ and the centroid of the cluster of the active failure mode.
@@
 If the unit’s failure is misclassified, $w_{i}$ would be large, and thus, unit $i$ is assigned a smaller weight in the regression problem.
-Next, We regress the $\ln{TTF}$ on the MFPC scores using a weighted linear regression shown in equation (17).
-The solution to this problem can be found using weighted least squares regression [43].
+Next, We regress the standardized $\ln{TTF}$ on the MFPC-scores using a penalized weighted linear regression shown in Eq (20).
 For variable selection, we include the lasso penalty in the optimization problem.
+To ensure a regression coefficient represents the significance of their respective predictor to the prediction of the failure time, Eq (20) should be read such that $\zeta_{i,h}^{k^{*}}$ is standardized.
 $\min_{c_{0}^{k^{*}},c_{1}^{k^{*}},\ldots,c_{H_{k^{*}}}^{k^{*}}}\sum_{i\in\mathcal{I}_{t^{*}}^{k^{*}}}\left(\sqrt{w_{i}}y_{i}-\sqrt{w_{i}}c_{0}^{k^{*}}-\sum_{h=1}^{H_{k^{*}}}\sqrt{w_{i}}\zeta_{i,h}^{k^{*}}c_{h}^{k^{*}}\right)^{2}+\lambda\sum_{h=1}^{H_{k^{*}}}\lVert c_{h}^{k^{*}}\rVert$ (20)
 5.3.2 RUL prediction
-As we have shown, MFPCA enables the transformation of the functional regression problem to a multiple linear regression problem, which enables us to solve the problem only by estimating $H_{k^{*}}+1$ parameters.
-After fitting the regression coefficients, we extract the MFPC scores for the test signal $\zeta_{test,1}^{k^{*}},\ldots,\zeta_{test,H_{k^{*}}}^{k^{*}}$.
-These scores are then used to predict the RUL at time $t^{*}$ using equation (18).
-${RUL}_{test}=\exp\left(c_{0}^{k^{*}}+\sum_{k=1}^{H_{k^{*}}}\zeta_{test,h}^{k^{*}}c_{k}^{k^{*}}\right)-t^{*}$ (18)
+As we have shown, MFPCA enables the transformation of the functional regression problem into a multiple linear regression problem, which allows us to solve the problem by estimating only $H_{k^{*}}+1$ parameters.
+After fitting the regression coefficients, we extract the MFPC-scores for the test signal $\zeta_{test,1}^{k^{*}},\ldots,\zeta_{test,H_{k^{*}}}^{k^{*}}$.
+These MFPC scores are standardized using the sample mean and sample standard deviation of the MFPC scores from the training set.
+These scores are then used to predict the RUL at time $t^{*}$ using Eq (21).
+$\begin{split}{RUL}_{test}=\exp&\bigg{(}\Big{[}c_{0}^{k^{*}}+\sum_{k=1}^{H_{k^{*}}}\zeta_{test,h}^{k^{*}}c_{k}^{k^{*}}\Big{]}\times\\ &\Big{[}(1/(N-1))\sum_{i=1}^{N}\big{(}\ln{TTF}_{i}-(1/N)\sum_{i=1}^{N}\ln{TTF}_{i}\big{)}^{2}\Big{]}\\ &+(1/N)\sum_{i=1}^{N}\ln{TTF_{i}}\bigg{)}-t^{*}\end{split}$ (21)
 6 Case Study 1: Simulated Dataset
-We present a case study using a simulated data set to evaluate the performance of our methodology in terms of its ability to cluster the training data, effectiveness in selecting informative sensors for each failure mode, and accuracy in predicting the RUL.
-Furthermore, we test the robustness of our methodology to varying degrees of signal noise.
-In section VI-A, we describe the simulation model.
-In section VI-B, we display the simulated data and evaluate the performance of the MGR-ASGL model at clustering and sensor selection.
-Finally, we evaluate the prediction accuracy of our online prognostics approach in section VI-C.
+We present a simulation study that mimics the operational and diagnostic challenges of a deep-space habitat.
+In such systems, critical subsystems like life support, power generation, etc., are monitored by numerous sensors, and degradation can occur through multiple unknown failure modes.
+Due to the scarcity of in-mission failure data and the infeasibility of performing full-scale failure testing in space, simulation-based validation is pursued in this case study.
+This study enables us to systematically evaluate our methodology’s ability to: (1) identify latent failure modes in unlabeled multivariate sensor data, (2) select informative sensors specific to each mode, and (3) accurately predict remaining useful life (RUL).
+To emulate the complex telemetry environment of deep-space habitats, we use a modified version of the simulation model from [20] to generate degradation signals for engineered systems monitored by 20 sensors.
+Each system is assumed to experience one of two possible failure modes, FM 1 or FM 2, with only a subset of sensors being informative for degradation under each mode.
+This setup reflects the sensor redundancy and noise similar to onboard telemetry in a deep space habitat.
+We generate data for 200 systems per failure mode, using 160 systems for training and 40 for testing.
+To evaluate the robustness of our methodology, we simulate varying signal-to-noise ratios (SNRs) to mimic differences in sensor quality and environmental interference expected in long-duration missions.
+The resulting dataset includes multivariate sensor signals and the true time-to-failure (TTF), used for assessing prediction accuracy.
 6.1 Simulation Model
-We utilize a modified version of the simulation model used in [13] to generate degradation signals for a set of engineering systems.
-We assume that each system is monitored by 20 sensors and exhibits one of two possible failure modes that we refer to as failure mode 1 (FM 1) and failure mode 2 (FM 2).
-We generate data for 200 systems for each failure mode.
-Out of the 200 systems, we use 160 systems for model training and 40 systems for testing model performance.
-The simulated data consists of the sensor signals for each system and the corresponding failure time, which is utilized later as the ground truth to evaluate prediction accuracy.
-The signals of the underlying degradation process for system $i$ exhibiting failure mode $k$ are generated according to equation (19).
+The signals of the underlying degradation process for system $i$ exhibiting failure mode $k$ are generated according to Eq (22).
 The signal is parameterized by the random coefficient $\theta_{i}^{k}\sim N\left(\mu_{k},0.1^{2}\right)$, which is used to control the degradation rate.
@@
 $s_{i}^{k}(t)=-\frac{\theta_{i}^{k}}{\ln t}$ (22)
-This paragraph describes the procedure for simulating degradation signals for each sensor.
-Here we consider a practical scenario where only a subset of sensors is correlated with the underlying degradation process associated with a given failure mode.
+We consider a scenario where only a subset of sensors is correlated with the underlying degradation process associated with a given failure mode.
 These sensors are the informative sensors and their signal trends reflect the severity and progression of the degradation process.
@@
 Out of the 20 sensors, we set $\mathcal{J}_{1}=\{5,12,16,19\}$ as the set of informative sensors for FM 1 and $\mathcal{J}_{2}=\{3,7,9,19\}$ as the set of informative sensors for FM 2.
-The $p^{th}$ sensor signal for the $i^{th}$ system exhibiting failure mode $k$ is generated using equation (20).
+The $p^{th}$ sensor signal for the $i^{th}$ system exhibiting failure mode $k$ is generated using Eq (23).
 $s_{i,p}^{k}(t)=-\frac{\theta_{i,p}^{k}}{\ln t}+\epsilon_{i,p}^{k}(t)$ (23)
-In this equation, $\epsilon_{i,p}^{k}(t)\sim N\left(0,\left(\sigma_{k}^{p}\right)^{2}\right)$ is a white noise process parameterized by $\sigma_{k}^{p}=\mu_{k}/SNR_{p}^{k}$, where $SNR_{p}^{k}$ is the signal-to-noise for sensor $p$ and failure mode $k$.
+In this Eq, $\epsilon_{i,p}^{k}(t)\sim N\left(0,\left(\sigma_{k}^{p}\right)^{2}\right)$ is a white noise process parameterized by $\sigma_{k}^{p}=\mu_{k}/SNR_{p}^{k}$, where $SNR_{p}^{k}$ is the signal-to-noise for sensor $p$ and failure mode $k$.
 To account for the correlation between the signal trends of the informative sensors and the underlying degradation process, we generate $\theta_{i,p}^{k}$ from the following conditional distribution $\theta_{i,p}^{k}\mid\theta_{i}^{k}\sim N\left(\mu_{k}\left(1-\sqrt{1-\rho_{p}^{k}}\right),(0.1)^{2}\right)$ such that the correlation, $\rho_{p}^{k}$, is uniformly sampled from the interval $[0.80,0.99]$.
@@
 For each dataset and criterion, we fit the MGR-ASGL model using these tuning parameters.
-Table 1: Tuning parameter selection for case study 1
-$SNR_{p}^{k}$ [2,5] [5,8] [8,11]
-$\lambda$ 0.0466 0.0258 0.0258
-$\alpha$ 1 0 0.25
 We then evaluate the performance of the MGR-ASGL model on the three datasets.
@@
 The results of this performance evaluation are shown in Table 3.
+Table 2: Tuning parameter selection for case study 1
+$SNR_{p}^{k}$ [2,5] [5,8] [8,11]
+$\lambda$ 0.0466 0.0258 0.0258
+$\alpha$ 1 0 0.25
 Table 3: MGR-ASGL performance for case study 1
@@
 We compute the “Estimated Life” as the current operating time plus the predicted RUL.
-A relative error is then computed using equation (21).
+A relative error is then computed using Eq (24).
 […]
@@
 Figure 5: The system diagram used in Case Study 2 adapted from [55]
-To further validate our methodology, we evaluate its performance on one of the Commercial Modular Aero-Propulsion System Simulation (C-MAPSS) datasets [44].
-C-MAPSS is a simulation module that generates turbofan engine degradation datasets.
+To further validate our methodology in a realistic operational context, we evaluate its performance on a benchmark dataset that closely mirrors subsystem degradation in deep-space habitats.
+Specifically, we apply our approach to the NASA’s Commercial Modular Aero-Propulsion System Simulation (C-MAPSS) dataset [55].
+C-MAPSS simulates degradation in the components of turbofan engines that share telemetry and fault characteristics with propulsion and environmental control systems of deep space habitats.
+C-MAPSS involves multivariate sensor monitoring from healthy operation to system failure, with no explicit failure mode labels and only partial degradation histories for test systems.
 Figure 5 displays a diagram of this turbofan engine.
@@
 Therefore, our goal is to use our methodology to predict the remaining life of each testing unit following the final cycle at which their degradation was observed.
+Our first step is to perform traditional FPCA on the signals for each sensor.
+We obtain initial failure mode labels for the training set by applying K-means clustering to all FPC-scores.
+Then, for each sensor, we apply K-means clustering to the FPC-scores associated with that sensor to obtain the covariate information needed to perform CA-FPCA.
+Following feature extraction, we use 3-fold cross-validation to determine $(\lambda,\alpha)$.
+The range of values for these tuning parameters are identical to that of the simulation case study.
+The combination of tuning parameters that minimize the mean squared error across the three folds is $(\lambda_{MSE},\alpha_{MSE})=(0.1089,1)$.
+Table 6 displays the $l_{2}$-norms of the regression coefficients for each sensor.
+The sensors selected are those with positive $l_{2}$-norms.
 Table 5: Sensors used in case study 2 (C-MAPPS-3 [55])
@@
 21 W32 LPT coolant bleed
-Figure 5: Raw sensor data used for offline step in case study 2. The data is inherently unlabeled i.e., failure modes are not known.
-Our first step is to perform traditional FPCA on the signals for each sensor.
-We obtain initial failure mode labels for the training set by applying K-means clustering to all FPC-scores.
-Then, for each sensor, we apply K-means clustering to the FPC-scores associated with that sensor to obtain the covariate information needed to perform CA-FPCA.
-Following feature extraction, we use 3-fold cross-validation to determine $(\lambda,\alpha)$.
-The range of values for these tuning parameters are identical to that of the simulation case study.
-The combination of tuning parameters that minimize the mean squared error across the three folds is $(\lambda_{MSE},\alpha_{MSE})=(0.1089,1)$.
-Table 5 displays the $l_{2}$-norms of the regression coefficients for each sensor.
-The sensors selected are those with positive $l_{2}$-norms.
-Table 5: Sensor Selection for Case Study 2
-Sensor ID FM 1 FM2 Sensor ID FM 1 FM 2
-2 0.266 0 12 0 1.148
-3 0.584 0 13 0.90 0
-4 0.897 0.033 14 0.148 0.119
-7 0 0.877 15 0.501 0
-8 0.117 0 17 0.230 0.042
-9 0.545 0.291 20 0 0.573
-11 0 0 21 0.116 0.318
 In addition to performing sensor selection, the MGR-ASGL model provides cluster labels for each observed failure.
@@
 Finally, we perform Weighted Time-varying Functional Regression to regress $lnTTF$ on the MFPC-scores using Leave-one-out cross-validation to select the tuning parameter.
-Benchmarking: To analyze our methodology, we compare its performance to that of [32] (Chehade et al.
-2018), whose methodology assumes knowledge of the failure modes but does not perform sensor selection (outside of removing the blank sensors), and [33] (Li et al.
-2023), whose methodology assumes knowledge of a fraction of the training set and does perform sensor selection.
+Figure 6: Raw sensor data used for offline step in case study 2. The data is inherently unlabeled i.e., failure modes are not known.
+7.1 Comparison to Baselines
+To analyze our methodology, we compare its performance to that of [39] Chehade et al.
+(2018), whose methodology assumes knowledge of the failure modes but does not perform sensor selection (outside of removing the blank sensors), and [40] Li et al.
+(2023), whose methodology assumes knowledge of a fraction of the training set and does perform sensor selection.
 Figure 7 displays the relative prediction error averaged over all test systems with RUL at most 100, 80, 60, 40, and 20 cycles.
@@
 Furthermore, we were capable of achieving more accurate RUL prediction than both methods from earlier in the systems’ lives despite not having information about the actual failure mode.
+Table 6: Sensor Selection for Case Study 2
+Sensor ID FM 1 FM2 Sensor ID FM 1 FM 2
+2 0.266 0 12 0 1.148
+3 0.584 0 13 0.90 0
+4 0.897 0.033 14 0.148 0.119
+7 0 0.877 15 0.501 0
+8 0.117 0 17 0.230 0.042
+9 0.545 0.291 20 0 0.573
+11 0 0 21 0.116 0.318
 […]
@@
 8 Conclusion
-In this paper, we present a methodology for predicting the RUL of systems with multiple failure modes using multiple sensors.
-This methodology consists of two phases: an offline phase and an online phase.
-Assuming that failure modes are mutually exclusive, the offline phase consists of a framework for diagnosing the observed failures in the training set and selecting an optimal subset of sensors for monitoring.
-In this framework, we utilize CA-FPCA to capture the multiple failure modes exhibited by the sensor signals enabling the extraction of informative features.
-These features are used by our proposed EM algorithm to train the MGR-ASGL model which simultaneously diagnoses the training set failures and selects an optimal subset of informative sensors for each failure mode.
-In the online phase, we use MFPCA to fuse all the signals of the selected sensors into a parsimonious set of features to diagnose the failure mode for an actively degrading system.
-Then MFPCA is re-used to fuse the signals from informative sensors of the diagnosed failure mode for RUL prediction.
-RUL prediction is done using a time-varying weighted functional regression model.
-To assess the performance of our methodology, we implemented it on two datasets – a) a simulated dataset and b) the NASA turbofan engine degradation dataset.
-For the simulated dataset, we tested the robustness of our methodology to various degrees of signal noise.
-We demonstrated that with a sufficiently high $\lambda$ value, the clustering matches the actual clusters at an accuracy of about 80%.
-Furthermore, the most significant sensors occupied 3 of the top 4 slots with regard to $l_{2}$-norm of regression coefficients.
-Finally, the methodology showed high accuracy with an average relative prediction error of no larger than 3, irrespective of the percentage of lifetime the prediction was performed.
-In our real-world data case study, results show that our model makes early predictions more accurately than the two benchmarks while achieving comparable accuracy in later predictions.
-To achieve this, we identified 10 sensors to monitor the first failure mode while identifying 8 sensors to monitor the second failure mode.
-Our methodology removed one sensor completely.
-The region in which our methodology performed worst relative to the two baselines was when the RUL was between 25 and 50.
-A possibility for improvement would be to relax the assumption of a linear relationship between the $lnTTF$ and the sensor signals.
+In this paper, we presented a prognostics methodology tailored to the unique challenges of deep-space habitats.
+Complex engineering systems such as deep space habitats, are monitored by a large number of sensors and are susceptible to multiple failure modes that are often unlabeled.
+Our approach addresses the need for autonomous and accurate remaining useful life prediction in environments where expert diagnosis and ground-based maintenance are not feasible.
+The methodology consists of two phases.
+In the offline phase, we extract failure-mode-specific features using component-aligned functional principal component analysis.
+These features are used by a novel expectation maximization algorithm to train a regression model that simultaneously clusters systems by failure mode and selects the most informative sensors for each mode.
+In the online phase, we use multivariate functional principal component analysis to reduce the dimensionality of real-time sensor data, diagnose the active failure mode, and predict remaining useful life using a time-varying weighted functional regression model.
+We validated our methodology using two case studies.
+The first is a simulated dataset that reflects key properties of deep-space habitats, including mode-specific degradation behavior, redundant sensors, and variable signal-to-noise ratios.
+Our method achieved strong performance in clustering failure modes, identifying informative sensors, and predicting remaining useful life with low relative error.
+The second case study applied our method to the C-MAPSS dataset, which we used to test real-world feasibility under partial degradation observation and unlabeled failure conditions.
+Our model outperformed baseline approaches in early-life prediction and produced interpretable results by selecting a minimal and relevant set of sensors.
+This work is based on the assumption of a linear relationship between the log of time-to-failure (TTF) and the sensor signals.
+A possibility for improvement would be to relax this assumption.
 However, embedding a variable selection technique within a nonlinear regression model is an open topic in the literature.
-Therefore, the extension of this work to a nonlinear setting is a topic for future work.
+Future work will focus on extending the methodology to nonlinear models to improve prediction accuracy and capture complex degradation trends in autonomous mission scenarios.
 Acknowledgments
@@
 Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Aeronautics and Space Administration.
+Appendix A Log Likelihood
+Definition 1.
+The log-likelihood function for the MGR model is given by equation (25).
+This function is called as the ”incomplete-data log-likelihood” (IDLL), since we do not know the cause of the observed failure in the data.
+$\ell(\boldsymbol{\Theta};\boldsymbol{Y})=\sum_{i=1}^{N}\ln\sum_{k=1}^{K}\pi_{k}\times\ \frac{\rho_{k}}{\sqrt{2\pi}}\exp\left(-\frac{1}{2}\left(\rho_{k}y_{i}-\psi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\psi}_{p,k}\right)^{2}\right)$ (25)
+Definition 2.
+The “complete-data log-likelihood” (CDLL) given by equation (26) assumes knowledge of the underlying causes of the observed failure in the data.
+$\begin{split}\ell_{C}(\boldsymbol{\Theta};\boldsymbol{Y},\mathbb{Z})&=\sum_{i=1}^{N}\sum_{k=1}^{K}\boldsymbol{Z}_{i}[k]\ln\pi_{k}\left(\frac{\rho_{k}}{\sqrt{2\pi}}\right)\times\\ &\exp\left(-\frac{1}{2}\left(\rho_{k}y_{i}-\psi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\psi}_{p,k}\right)^{2}\right)\end{split}$ (26)
+The IDLL given in equation (25) is difficult to optimize globally [56].
+Instead, we search for a local optimum by using EM algorithm, an iterative algorithm suitable for fitting probability models with latent variables.
+To derive the steps of the EM algorithm, we note that since $\ln(.)$ is a concave function We use Jensen’s inequality to form a lower bound on the IDLL.
+This lower bound is the expectation of the CDLL given in equation (26) and shown in Lemma 1.
+Lemma 1.
+The expectation of the CDLL w.r.t distribution $g$ is a lower bound for the IDLL i.e.,
+$\mathbb{E}_{\boldsymbol{g}}\left[\ell_{C}(\boldsymbol{\Theta};\boldsymbol{Y},\mathbb{Z})\right]\leq\ell(\boldsymbol{\Theta};\boldsymbol{Y})$
+Proof.
+Let $g$ be an arbitrary distribution on $\boldsymbol{Z}_{i}[k]$. Then we can re-write equation (25) as follows:
+$\ell(\boldsymbol{\Theta};\boldsymbol{Y})=\sum_{i=1}^{N}\ln\left(\sum_{k=1}^{K}\frac{g\left(\boldsymbol{Z}_{i}[k]\right)p\left(y_{i},\boldsymbol{Z}_{i}[k];\boldsymbol{\Theta}\right)}{g\left(\boldsymbol{Z}_{i}[k]\right)}\right)\vspace*{-1em}$
+where,
+$\begin{split}&p\left(y_{i},\boldsymbol{Z}_{i}[k];\boldsymbol{\Theta}\right)\\ &=\pi_{k}\cdot\left(\frac{\rho_{k}}{\sqrt{2\pi}}\exp\left(-\frac{1}{2}\left(\rho_{k}y_{i}-\psi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\psi}_{p,k}\right)^{2}\right)\right)\end{split}$
+Since $\ln(.)$ is concave, from Jensen’s inequality we can write that,
+$\ell(\boldsymbol{\Theta};\boldsymbol{Y})\geq\sum_{i=1}^{N}\sum_{k=1}^{K}g\left(\boldsymbol{Z}_{i}[k]\right)\ln{\frac{p\left(y_{i},\boldsymbol{Z}_{i}[k];\boldsymbol{\Theta}\right)}{g\left(\boldsymbol{Z}_{i}[k]\right)}}\vspace*{-1em}$
+and thus,
+$\ell(\boldsymbol{\Theta};\boldsymbol{Y})\geq\mathbb{E}_{\boldsymbol{g}}\left[\ell_{C}(\boldsymbol{\Theta};\boldsymbol{Y},\mathbb{Z})\right]-\sum_{i=1}^{N}\sum_{k=1}^{K}g\left(\boldsymbol{Z}_{i}[k]\right)\ln{g\left(\boldsymbol{Z}_{i}[k]\right)}\vspace*{-1em}$
+The second term on the right side of the inequality is the negative entropy of the $g$ distribution. Since it is positive, the inequality holds. ∎

### v4

@@
-Sensor-fusion based Prognostics for Deep-space Habitats Exhibiting Multiple Unlabeled Failure Modes
+
 Abstract
@@
 The methodology consists of two phases.
-In the offline phase, we model system failure times using a mixture of Gaussian regressions and apply a novel Expectation-Maximization algorithm to cluster degradation trajectories and select mode-specific sensors.
+In the offline phase, we model system failure times using a mixture of Gaussian regressions and apply an Expectation-Maximization algorithm to cluster degradation trajectories and select mode-specific sensors.
 In the online phase, we extract low-dimensional features from the selected sensors to diagnose the active failure mode and predict RUL using a weighted regression model.
 We demonstrate the effectiveness of our approach on a simulated dataset that reflects deep-space telemetry characteristics and on a real-world engine degradation dataset, showing improved accuracy and interpretability over existing methods.
-keywords:
-Deep Space Habitats , Multiple Failure Modes , Remaining Useful Life , Sensor Fusion , Failure Diagnosis , Prognostics
-journal: Acta Astronautica
-[…]
-[…]
+{IEEEkeywords}
+Deep-space Habitats, Failure Diagnosis, Multiple Failure Modes, Prognostics, Remaining Useful Life, Sensor Fusion
 1 INTRODUCTION
-A deep-space habitat is a complex engineering system (CES), composed of multiple interacting subsystems that must work together to support human life and mission operations over long durations.
-These subsystems such as the environmental control and life support system, energy and power systems, thermal control, robotic agents, and structural components etc., must function reliably in harsh and isolated environments, where failure of any component can lead to mission-critical consequences [1, 2, 3, 4].
+Deep-space habitat is a complex engineering system (CES), composed of multiple interacting subsystems that must work together to support human life and mission operations over long durations.
+These subsystems such as the environmental control and life support system, energy and power systems, thermal control, robotic agents, and structural components etc., must function reliably in harsh and isolated environments, where failure of any component can lead to mission-critical consequences [1, 2].
 The complexity of such systems arises not only from the tight integration of these subsystems but also from their heterogeneous origins.
@@
 These stages align with two mission stages: during early deployment, the habitat operates under limited supervision to collect historical sensor data for initializing the model; after this setup phase, autonomous health monitoring and prediction begin.
-Figure 1: Example of a deep space habitat with multiple subsytems.
-Health of each subsystem is monitored by a subset of onboard sensors.
-The subset of informative sensors vary for different failure modes.
 Main Contributions.
@@
 The key technical contributions of this paper can be summarized as:
-  • 1.
   • We develop a feature extraction methodology that fuses multivariate sensor data from systems exhibiting multiple failure modes into compact and informative representations.
-  • 2.
   • We propose a failure-mode-aware sensor selection approach using an Expectation-Maximization algorithm that jointly clusters unlabeled failure events and selects informative sensors for each failure mode.
-  • 3.
   • We present an integrated online framework that uses real-time data to (1) diagnose the active failure mode and (2) predict RUL using a mode-specific regression model.
-  • 4.
   • We demonstrate the effectiveness and robustness of the proposed methodology through two case studies relevant to deep-space habitats: (1) a controlled simulation designed to reflect key telemetry challenges, and (2) a real-world NASA turbofan engine degradation dataset that mimics sensor-driven subsystem degradation in long-duration missions.
@@
 We conclude the paper in Section 8 with a discussion of future work.
+Figure 1: Example of a deep-space habitat with multiple subsytems.
+Health of each subsystem is monitored by a subset of onboard sensors.
+The subset of informative sensors vary for different failure modes.
 2 LITERATURE REVIEW
@@
 However, these approaches typically assume a single dominant failure mode.
-In reality, most CES can fail through multiple failure modes, reflecting the structure and function of their subsystems.
+In reality, most CES experience multiple failure modes, reflecting the structure and function of their subsystems.
 Competing risk models have been widely used to represent such systems.
@@
 As a result, many researchers have shifted toward methods that directly model sensor data to predict RUL, though few focus on multiple failure modes.
-It is common for researchers to develop general RUL prediction methodologies and then apply them to datasets with multiple failure modes.
+Often, researchers develop general RUL prediction methodologies and then apply them to datasets with multiple failure modes.
 For example, [33] proposed a correlation and relative entropy feature engineering framework for complex systems and [34] proposed a deep convolutional neural network.
@@
 Furthermore, [37] fit random coefficient models to failure mode-specific health indices.
-However, these articles do not incorporate a sensor selection algorithm and they rely on labeling of the failure modes determined either through visual inspection of domain knowledge.
+However, these articles do not incorporate a sensor selection algorithm and they rely on labeling of the failure modes determined either through visual inspection or domain knowledge.
 Consequently, researchers have begun addressing these gaps.
@@
 3 OVERVIEW OF THE PROGNOSTICS FRAMEWORK
-Figure 2: Overview of the proposed prognostics framework for deep-space habitats
 This section describes our prognostics framework, developed for autonomous health monitoring in deep space habitats.
@@
 The framework consists of two main components: “Offline Sensor Selection” and “Online Diagnosis and Prognostics”.
+Figure 2: Overview of the proposed prognostics framework for deep-space habitats
 I.
@@
   1. 1.
-    Diagnosis: Prior to the diagnosis step, the active failure mode is unknown therefore, all selected sensors are utilized in this step.
-We apply multivariate functional principal component analysis (MFPCA) to signals from all selected sensors to extract a compact set of features, called MFPC scores, that represent the joint behavior of all sensors.
-We classify the active failure mode by finding its K-nearest neighbors and assigning the most common failure mode among them.
+    Diagnosis: Prior to the diagnosis step, the active failure mode is unknown.
+Therefore, all selected sensors are utilized in this step.
+First we apply multivariate functional principal component analysis (MFPCA) to signals from all selected sensors to extract a compact set of features, called MFPC scores, that represent the joint behavior of all sensors.
+Then we classify the active failure mode by finding its K-nearest neighbors and assigning the most common failure mode among them.
   2. 2.
@@
 4 OFFLINE SENSOR SELECTION
-We consider a scenario where the deep space habitat is monitored by $P$ different sensors.
-We assume there exists a historical dataset comprised of $P$-dimensional degradation-based sensor signals from monitoring $N$ similar samples.
+We consider a scenario where the deep-space habitat is monitored by $P$ different sensors.
+We assume there exists a historical dataset comprised of $P$-dimensional degradation-based sensor signals from monitoring $N$ similar systems.
 We denote $s_{i,p}(t)$ as the signal data from sensor $p\in\{1,...,P\}$ for system $i\in\{1,...,N\}$ observed at time $t\in[0,T]$, where $T=min(\{TTF_{i}\}_{i=1}^{N})$.
 By truncating all signals to the minimum failure time, we ensure that all signals in the training set have the same time domain and that they are all included when modeling.
-For this framework, we assume that the total number of failure modes, $K$ is known, but that the actual failure modes in the training dataset are unknown.
+For this framework, we assume that the total number of failure modes, $K$ is known, but that the actual failure labels in the training dataset are unknown.
 This poses a challenge for CA-FPCA, which requires observed covariates for computation.
@@
 For a given system $i$, signals from cluster $k$ of sensor $p$ are modeled as:
-$s_{i,p}(t)=v_{i,p}^{k}(t)+\epsilon_{i,p}(t),i\in\mathcal{I}_{k}^{p}\vspace*{-1em}$ (1)
+$s_{i,p}(t)=v_{i,p}^{k}(t)+\epsilon_{i,p}(t),i\in\mathcal{I}_{k}^{p}\vskip-10.00002pt$ (1)
 where $v_{i,p}^{k}(t)$ is a smooth random function, and $\epsilon_{i,p}(t)$ are assumed to be independently and identically distributed (i.i.d.) errors with mean zero and variance $\varsigma_{p}^{2}$.
@@
 The mean and covariance functions of $v_{i,p}^{k}(t)$ are given by $\mu_{p}^{k}(t)$ and $\mathbb{C}_{p}^{k}(t,t^{\prime})$, respectively.
-Using Mercer’s theorem [49], the covariance can be decomposed as follows, $\mathbb{C}_{p}^{k}(t,t^{\prime})=\sum_{m=1}^{\infty}\lambda_{m,p}^{k}\phi_{m,p}^{k}(t)\phi_{m,p}^{k}(t^{\prime})$, where $\{\phi_{m,p}^{k}(t)\}_{m=1}^{\infty}$ are the orthogonal eigenfunctions of $\mathbb{C}_{p}^{k}(t,t^{\prime})$ and $\lambda_{1,p}^{k}\geq...\geq\lambda_{m,p}^{k}\geq...$ are the corresponding eigenvalues.
+Using Mercer’s theorem [47], the covariance can be decomposed as follows, $\mathbb{C}_{p}^{k}(t,t^{\prime})=\sum_{m=1}^{\infty}\lambda_{m,p}^{k}\phi_{m,p}^{k}(t)\phi_{m,p}^{k}(t^{\prime})$, where $\{\phi_{m,p}^{k}(t)\}_{m=1}^{\infty}$ are the orthogonal eigenfunctions of $\mathbb{C}_{p}^{k}(t,t^{\prime})$ and $\lambda_{1,p}^{k}\geq...\geq\lambda_{m,p}^{k}\geq$...
+are the corresponding eigenvalues.
 By projecting the mean-subtracted signal data onto the eigenfunctions, we can represent the signals by Eq (2).
-$s_{i,p}(t)=\mu_{p}^{k}(t)+\sum_{m=1}^{\infty}\xi_{i,m,p}^{k}\phi_{m,p}^{k}(t)+\epsilon_{i,p}(t)\vspace*{-1em}$ (2)
+$s_{i,p}(t)=\mu_{p}^{k}(t)+\sum_{m=1}^{\infty}\xi_{i,m,p}^{k}\phi_{m,p}^{k}(t)+\epsilon_{i,p}(t)\vskip-10.00002pt$ (2)
 where $\xi_{i,m,p}^{k}$ is the $m$th CA-FPC-score with mean 0 and variance $\lambda_{m,p}^{k}$.
@@
 4.2 Optimization of the MGR-ASGL model
-Let $Y_{i}=\frac{\ln{TTF}_{i}-(1/N)\sum_{i=1}^{N}\ln{TTF}_{i}}{(1/(N-1))\sum_{i=1}^{N}\big{(}\ln{TTF}_{i}-(1/N)\sum_{i=1}^{N}\ln{TTF}_{i}\big{)}^{2}}$, for $i=1,2…,N$ denote the random variable corresponding to the standardized natural logarithm of the time-to-failure (i.e., $\ln{TTF}$) of system $i$.
+Let $Y_{i}=\frac{\ln{TTF}_{i}-(1/N)\sum_{i=1}^{N}\ln{TTF}_{i}}{(1/(N-1))\sum_{i=1}^{N}\big(\ln{TTF}_{i}-(1/N)\sum_{i=1}^{N}\ln{TTF}_{i}\big)^{2}}$, for i=1,2,Ni=1,2…,N denote the random variable corresponding to the standardized natural logarithm of the time-to-failure (i.e., $\ln{TTF}$) of system $i$.
 Furthermore, let $\boldsymbol{Z}_{i}$ denote a categorical random vector of size $K$ used to encode the unknown failure mode of system $i$.
@@
 Thus, we assume that the conditional distribution of $Y_{i}$ given $\boldsymbol{Z}_{i}[k]=1$ is Gaussian with PDF:
-$f_{Y_{i}|\boldsymbol{Z}_{i}}(y_{i}|\boldsymbol{Z}_{i}[k]=1)=\frac{1}{\sqrt{2\pi}\sigma_{k}}\exp\Big{(}-\frac{1}{2\sigma_{k}^{2}}(y_{i}-\beta_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\beta}_{p,k})^{2}\Big{)}$ (3)
+$\begin{split}f_{Y_{i}|\boldsymbol{Z}_{i}}(y_{i}|\boldsymbol{Z}_{i}[k]=1)=\frac{1}{\sqrt{2\pi}\sigma_{k}}&\exp\Big(-\frac{1}{2\sigma_{k}^{2}}(y_{i}-\beta_{0,k}\\ &-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\beta}_{p,k})^{2}\Big)\end{split}$ (3)
 To find the marginal distribution of $Y_{i}$ we use the law of total probability, summing over the joint distribution of $Y_{i}$ and $\boldsymbol{Z}_{i}$ as follows:
-$f_{Y_{i}}(y_{i})=\sum_{k=1}^{K}\pi_{k}\frac{1}{\sqrt{2\pi}\sigma_{k}}\exp\Big{(}-\frac{1}{2\sigma_{k}^{2}}(y_{i}-\beta_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\beta}_{p,k})^{2}\Big{)}$ (4)
+$\begin{split}f_{Y_{i}}(y_{i})=\sum_{k=1}^{K}\pi_{k}\frac{1}{\sqrt{2\pi}\sigma_{k}}&\exp\Big(-\frac{1}{2\sigma_{k}^{2}}(y_{i}-\beta_{0,k}\\ &-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\beta}_{p,k})^{2}\Big)\end{split}$ (4)
 Therefore, the marginal distribution of $Y_{i}$ is a mixture of Gaussian regressions.
@@
 Assuming that $Y_{1},\dots,Y_{N}$ are independent, the negative IDLL is computed as follows:
-$-\ell(\Theta|\boldsymbol{Y})=-\sum_{i=1}^{N}\ln\Big{[}\sum_{k=1}^{K}\pi_{k}\rho_{k}\frac{1}{\sqrt{2\pi}}\exp\Big{(}-\frac{1}{2}(y_{i}\rho_{k}-\varphi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,k})^{2}\Big{)}\Big{]}$ (5)
+$\begin{split}-\ell(\Theta|\boldsymbol{Y})=-\sum_{i=1}^{N}\ln\Big[\sum_{k=1}^{K}\pi_{k}&\rho_{k}\frac{1}{\sqrt{2\pi}}\exp\Big(-\frac{1}{2}(y_{i}\rho_{k}\\ &-\varphi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,k})^{2}\Big)\Big]\end{split}$ (5)
 Estimating $\Theta$ requires minimizing Eq 5.
@@
 To utilize the EM algorithm, we first construct the negative complete-data log likelihood (CDLL).
-$\begin{split}-\ell_{C}(\Theta|\boldsymbol{Y},\mathbb{Z})&=-\ln\prod_{i=1}^{N}\prod_{k=1}^{K}f_{Y_{i}}(y_{i})^{\boldsymbol{Z}_{i}[k]}=-\sum_{i=1}^{N}\sum_{k=1}^{K}\boldsymbol{Z}_{i}[k]\ln\Big{[}\pi_{k}\rho_{k}\frac{1}{\sqrt{2\pi}}\\ &\times\exp\Big{(}-\frac{1}{2}(y_{i}\rho_{k}-\varphi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,k})^{2}\Big{)}\Big{]}\end{split}$ (6)
+$\begin{split}-\ell_{C}(\Theta|\boldsymbol{Y},\mathbb{Z})&=-\ln\bigg(\prod_{i=1}^{N}\prod_{k=1}^{K}f_{Y_{i}}(y_{i})^{\boldsymbol{Z}_{i}[k]}\bigg)\\ &=-\sum_{i=1}^{N}\sum_{k=1}^{K}\boldsymbol{Z}_{i}[k]\ln\Big[\pi_{k}\rho_{k}\frac{1}{\sqrt{2\pi}}\exp\Big(\\ &-\frac{1}{2}(y_{i}\rho_{k}-\varphi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,k})^{2}\Big)\Big]\end{split}$ (6)
 In Appendix A, we show that the expectation of the negative CDLL with respect to the conditional distribution of $\mathbb{Z}$ given $\boldsymbol{Y}$ forms an upper bound for the negative IDLL.
@@
   2. 2.
-    E-step: Given $\Theta^{(m)}$, the expectation of $-\ell_{C}(\Theta|\boldsymbol{Y},\mathbb{Z})$ is given as:
-    $\begin{split}Q(\Theta,\Theta^{(m)})&=\mathbb{E}[-\ell_{C}(\Theta|\boldsymbol{Y},\mathbb{Z})|\boldsymbol{Y},\Theta^{(m)}]=-\sum_{i=1}^{N}\sum_{k=1}^{K}\gamma_{i,k}\ln\Big{[}\pi_{k}\rho_{k}\frac{1}{\sqrt{2\pi}}\\ &\times\exp\Big{(}-\frac{1}{2}(y_{i}\rho_{k}-\varphi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,k})^{2}\Big{)}\Big{]}\end{split}$ (8)
+    E-step: Given $\Theta^{(m)}$, the expectation of $\ell_{C}(\Theta|\boldsymbol{Y},\mathbb{Z})$ is given as:
+    $\begin{split}Q(\Theta,\Theta^{(m)})&=\mathbb{E}[-\ell_{C}(\Theta|\boldsymbol{Y},\mathbb{Z})|\boldsymbol{Y},\Theta^{(m)}]\\ &=-\sum_{i=1}^{N}\sum_{k=1}^{K}\gamma_{i,k}\ln\Big[\pi_{k}\rho_{k}\frac{1}{\sqrt{2\pi}}\exp\Big(\\ &-\frac{1}{2}(y_{i}\rho_{k}-\varphi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,k})^{2}\Big)\Big]\end{split}$ (8)
     where,
-    $\begin{split}\gamma_{i,k}&=\mathbb{E}(\boldsymbol{Z}_{i}[k]|Y_{i}=y_{i},\Theta^{(m)})=\Pr(\boldsymbol{Z}_{i}[k]=1|Y_{i}=y_{i},\Theta^{(m)})\\ &=\frac{\pi_{k}^{(m)}\rho_{k}^{(m)}\frac{1}{\sqrt{2\pi}}\exp\Big{(}-\frac{1}{2}(y_{i}\rho_{k}^{(m)}-\varphi_{0,k}^{(m)}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,k}^{(m)})^{2}\Big{)}}{\sum_{l=1}^{K}\pi_{l}^{(m)}\rho_{l}^{(m)}\frac{1}{\sqrt{2\pi}}\exp\Big{(}-\frac{1}{2}(y_{i}\rho_{l}^{(m)}-\varphi_{0,l}^{(m)}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,l}^{(m)})^{2}\Big{)}}\end{split}$ (9)
+    $\begin{split}\gamma_{i,k}&:=\mathbb{E}\!\left(\boldsymbol{Z}_{i}[k]\mid Y_{i}=y_{i},\Theta^{(m)}\right)\\ &=\Pr\!\left(\boldsymbol{Z}_{i}[k]=1\mid Y_{i}=y_{i},\Theta^{(m)}\right)\\ &=\footnotesize\frac{\begin{aligned} &\pi_{k}^{(m)}\rho_{k}^{(m)}\frac{1}{\sqrt{2\pi}}\exp\!\Big(-\frac{1}{2}\big(y_{i}\rho_{k}^{(m)}-\varphi_{0,k}^{(m)}\\ &\hskip 34.14322pt-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,k}^{(m)}\big)^{2}\Big)\end{aligned}}{\begin{aligned} &\sum_{l=1}^{K}\pi_{l}^{(m)}\rho_{l}^{(m)}\frac{1}{\sqrt{2\pi}}\exp\!\Big(-\frac{1}{2}\big(y_{i}\rho_{l}^{(m)}-\varphi_{0,l}^{(m)}\\ &\hskip 34.14322pt-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,l}^{(m)}\big)^{2}\Big)\end{aligned}}\end{split}$ (9)
     is the responsibility that failure mode $k$ takes for the failure of system $i$.
@@
 To perform sensor selection, we augment $Q(\Theta,\Theta^{(m)})$ with the adaptive sparse group lasso penalty and form the following optimization problem:
-    $\min_{\Theta}\Big{\{}Q(\Theta,\Theta^{(m)})+\lambda\sum_{k=1}^{K}\pi_{k}\Big{(}\alpha\sum_{p=1}^{P}\|\boldsymbol{\varphi}_{p,k}\|_{1}+(1-\alpha)\sum_{p=1}^{P}\sqrt{q_{p}}\|\boldsymbol{\varphi}_{p,k}\|_{2}\Big{)}\Big{\}}$ (10)
+    $\begin{split}\min_{\Theta}\Big\{Q(\Theta,\Theta^{(m)})&+\lambda\sum_{k=1}^{K}\pi_{k}\Big(\alpha\sum_{p=1}^{P}\|\boldsymbol{\varphi}_{p,k}\|_{1}\\ &+(1-\alpha)\sum_{p=1}^{P}\sqrt{q_{p}}\|\boldsymbol{\varphi}_{p,k}\|_{2}\Big)\Big\}\end{split}$ (10)
     We refer to the problem in Eq 10 as Mixture of Gaussian Regression–Adaptive Sparse Group Lasso (MGR-ASGL).
 Because $\boldsymbol{\pi}$ is included in the penalty, we follow [49] and use a two-step optimization:
-    • (a)
-      Step 1: Improvement with respect to $\boldsymbol{\pi}$
+    • Step 1: Improvement w.r.t. $\boldsymbol{\pi}$
       We start by using the responsibilities from the E-step to identify a feasible solution $\bar{\boldsymbol{\pi}}^{(m+1)}=\big(\frac{1}{N}\sum_{i=1}^{N}\gamma_{i,1},\dots,\frac{1}{N}\sum_{i=1}^{N}\gamma_{i,K}\big)^{T}$.
@@
       where $u_{m}\in(0,1]$ is chosen to be the largest value in the sequence $\{1,\frac{1}{2},\frac{1}{4},\dots\}$ such that the objective function does not increase.
-    • (b)
-      Step 2: Improvement with respect to $\boldsymbol{\rho}$ and $\boldsymbol{\varphi}$
+    • Step 2: Improvement w.r.t. $\boldsymbol{\rho}$ and $\boldsymbol{\varphi}$
       Given the updated mixing coefficients, we can optimize Eq 10 with respect to the other parameters.
@@
 Therefore, for $k=1,\dots,K$, we solve
-      $\begin{split}\min_{\rho_{k},\varphi_{0,k},\boldsymbol{\varphi}_{1,k},\dots,\boldsymbol{\varphi}_{P,k}}&\Big{\{}-\sum_{i=1}^{N}\gamma_{i,k}\big{[}\ln\pi_{k}^{(m+1)}+\ln\rho_{k}-\frac{1}{2}(y_{i}\rho_{k}-\varphi_{0,k}\\ &-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,k})^{2}\big{]}\ +\lambda\pi_{k}^{(m+1)}\Big{(}\alpha\sum_{p=1}^{P}\|\boldsymbol{\varphi}_{p,k}\|_{1}\\ &+(1-\alpha)\sum_{p=1}^{P}\sqrt{q_{p}}\|\boldsymbol{\varphi}_{p,k}\|_{2}\Big{)}\Big{\}}\end{split}$ (12)
+      $\begin{split}\min\Big\{&-\sum_{i=1}^{N}\gamma_{i,k}\big[\ln\pi_{k}^{(m+1)}+\ln\rho_{k}\\ &-\frac{1}{2}(y_{i}\rho_{k}-\varphi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,k})^{2}\big]\ \\ &+\lambda\pi_{k}^{(m+1)}\Big(\alpha\sum_{p=1}^{P}\|\boldsymbol{\varphi}_{p,k}\|_{1}\\ &+(1-\alpha)\sum_{p=1}^{P}\sqrt{q_{p}}\|\boldsymbol{\varphi}_{p,k}\|_{2}\Big)\Big\}\end{split}$ (12)
       Note that the tuning parameters $\lambda$ and $\alpha$ do not change for different $k$.
@@
 5 ONLINE DIAGNOSIS AND PROGNOSTICS
-In this section, a subsystem(s) in deep-space habitat is considered to be degrading in real-time.
-Our goal is to predict the time remaining before the unit fails.
+In this section, a system within a deep-space habitat is considered to be degrading in real-time.
+Our goal is to predict the RUL of the system.
 We begin by showing how to fuse the signals from multiple sensors into a parsimonious set of features.
-Then, we demonstrate how we utilize the fused sensor information to i) diagnose the active failure mode of the degrading unit, in real-time, and ii) predict its RUL.
+Then, we demonstrate how we utilize the fused sensor information to i) diagnose the active failure mode of the degrading system in real-time, and ii) predict its RUL.
 5.1 Sensor Fusion: MFPCA
 Without loss of generality, let us redefine the sensor index as $\{1,2,\ldots,\mathcal{P}\}$, where $\mathcal{P}$ is number of sensors we seek to fuse.
-Now assume that we have monitored the degrading subsystem(s) for a duration given by the time interval $[0,t^{*}]$.
-Since we are interested in predicting the remaining lifetime, we only use signals from units with failure times greater than $t^{*}$.
-We define this set as the “$t^{*}-$training” set, whose indices we denote by the set $\mathcal{I}_{t^{*}}$, which has cardinality $N_{t^{*}}$.
+Now assume that we have monitored the degrading system for a duration given by the time interval $0,t^{*}]$.
+Since we are interested in RUL, we only use signals from training systems with failure times greater than $t^{*}$.
+We define this set as the $t^{*}$-training” set, whose indices we denote by the set $\mathcal{I}_{t^{*}}$, which has cardinality $N_{t^{*}}$.
 Furthermore, we smooth the degradation signals using the ’rloess’ method as described in [50].
@@
 Using the first $H$ multivariate eigenfunctions, we can represent the multivariate signal by Eq 13 where the MFPC-scores $\zeta$’s are computed according to Eq 14.
-$\boldsymbol{s}_{i}(t)=\boldsymbol{\mu}(t)+\sum_{h=1}^{H}\zeta_{i,h}\boldsymbol{\varpi}_{h}(t)+\boldsymbol{\epsilon}_{i}(t),\vspace*{-1.2em}$ (13)
-[…]
+$\boldsymbol{s}_{i}(t)=\boldsymbol{\mu}(t)+\sum_{h=1}^{H}\zeta_{i,h}\boldsymbol{\varpi}_{h}(t)+\boldsymbol{\epsilon}_{i}(t),\vskip-11.99998pt$ (13)
+$\zeta_{i,h}=\int_{0}^{t^{*}}(\boldsymbol{s}_{i}(t)-\boldsymbol{\mu}(t))^{T}\boldsymbol{\varpi}_{h}(t)dt\vskip-5.0pt$ (14)
 In practice, we perform MFPCA by collecting the $N_{t^{*}}$ observations, each consisting of the $\mathcal{P}$ sensor signals concatenated into a single vector.
@@
 5.2 Failure Mode Diagnosis: K-nearest Neighbors
-In order to diagnose the failure mode of a degrading unit at time $t^{*}$, we apply MFPCA on the $t^{*}$ training dataset.
+In order to diagnose the failure mode of a degrading system at time $t^{*}$, we apply MFPCA on the $t^{*}$ training dataset.
 Next, K nearest neighbors (KNN) is applied to the standardized MFPC-scores to identify the active failure mode.
 Here, the distance metric of the KNN classifier is defined using the Euclidean distance between the MFPC-scores.
-Given this diagnosis, we now develop a model for predicting the unit RUL.
+Given this diagnosis, we now develop a model for predicting system RUL.
 5.3 RUL Prediction: Weighted Functional Regression
@@
 It was first proposed by [51] and has been used for prognostics in [52, 18].
-In our problem setting, let $k^{*}$ denote the diagnosed failure mode and $\mathcal{I}_{t^{*}}^{k^{*}}\subset\mathcal{I}_{t^{*}}$ denote the set of indices for training units that failed due to failure mode $k^{*}$, but survived up to time $t^{*}$.
+In our problem setting, let $k^{*}$ denote the diagnosed failure mode and $\mathcal{I}_{t^{*}}^{k^{*}}\subset\mathcal{I}_{t^{*}}$ denote the set of indices for training systems that failed due to failure mode $k^{*}$, but survived up to time $t^{*}$.
 Furthermore, let $\mathcal{P}_{k^{*}}\subset\{1,2,...,\mathcal{P}\}$ denote the set of sensors selected (identified as most informative) for failure mode $k^{*}$.
 To predict RUL, we fit a functional regression model as shown in Eq (15), where $y_{i}=\ln{TTF_{i}}$, $\varphi_{0}^{k^{*}}$ is the bias term, $\boldsymbol{\varphi}^{k^{*}}(t)\in\mathcal{R}^{|\mathcal{P}_{k^{*}}|}$ is the multivariate coefficient function, and $\boldsymbol{s}_{i}^{k^{*}}(t)$ is the multivariate signal consisting only of sensors significant to failure mode $k^{*}$.
-[…]
+$\min_{\varphi_{0}^{k^{*}},\boldsymbol{\varphi}^{k^{*}}(t)}\sum_{i\in\mathcal{I}_{t^{*}}^{k^{*}}}\left(y_{i}-\varphi_{0}-\int_{0}^{t^{*}}\boldsymbol{\varphi}^{k^{*}}(t)^{T}\boldsymbol{s}_{i}^{k^{*}}(t)dt\right)^{2}\vskip-4.62497pt$ (15)
 We expand the coefficient function using the eigenfunctions obtained from applying MFPCA on the training set $\mathcal{I}_{t^{*}}^{k^{*}}$.
 The regression coefficient function can now be expanded as shown in Eq (16), where $H_{k^{*}}$ denotes the number of MFPC-scores that were retained for failure mode $k^{*}$.
-$\boldsymbol{\varphi}(t)=\sum_{h=1}^{H_{k^{*}}}c_{h}^{k^{*}}\boldsymbol{\varpi}_{k}^{k^{*}}(t)\vspace*{-0.5em}$ (16)
+$\boldsymbol{\varphi}(t)=\sum_{h=1}^{H_{k^{*}}}c_{h}^{k^{*}}\boldsymbol{\varpi}_{k}^{k^{*}}(t)\vskip-5.0pt$ (16)
 Likewise, we expand the multivariate signals using the eigenfunctions as shown in Eq (17).
-$\boldsymbol{s}_{i}^{k^{*}}(t)=\boldsymbol{\mu}^{k^{*}}(t)+\sum_{h=1}^{H_{k^{*}}}\zeta_{i,h}^{k^{*}}\boldsymbol{\varpi}_{h}^{k^{*}}(t)\vspace*{-0.5em}$ (17)
+$\boldsymbol{s}_{i}^{k^{*}}(t)=\boldsymbol{\mu}^{k^{*}}(t)+\sum_{h=1}^{H_{k^{*}}}\zeta_{i,h}^{k^{*}}\boldsymbol{\varpi}_{h}^{k^{*}}(t)\vskip-5.0pt$ (17)
 Since the eigenfunctions are orthogonal, substituting Eqs (16) and (17) into Eq (15) yields the following optimization problem,
-$\min_{c_{0}^{k^{*}},c_{1}^{k^{*}},\ldots,c_{H_{k^{*}}}^{k^{*}}}\sum_{i\in\mathcal{I}_{t^{*}}^{k^{*}}}\left(y_{i}-c_{0}^{k^{*}}-\sum_{h=1}^{H_{k^{*}}}\zeta_{i,h}^{k^{*}}c_{h}^{k^{*}}\right)^{2}\vspace*{-1em}$ (18)
+$\min_{c_{0}^{k^{*}},c_{1}^{k^{*}},\ldots,c_{H_{k^{*}}}^{k^{*}}}\sum_{i\in\mathcal{I}_{t^{*}}^{k^{*}}}\left(y_{i}-c_{0}^{k^{*}}-\sum_{h=1}^{H_{k^{*}}}\zeta_{i,h}^{k^{*}}c_{h}^{k^{*}}\right)^{2}\vskip-10.00002pt$ (18)
 where,
-$c_{0}^{k^{*}}=\varphi_{0}^{k^{*}}+\int_{0}^{t^{*}}\sum_{h=1}^{H_{k^{*}}}c_{h}^{k^{*}}\boldsymbol{\varpi}_{h}^{k^{*}}(t)^{T}\boldsymbol{\mu}^{k^{*}}(t)\vspace*{-0.5em}$ (19)
-Weighted Version: We note that the offline sensor selection algorithm may generate some outliers, which can impact the accuracy of the regression problem.
-These outliers become more apparent after identifying the active failure mode and the MFPC-scores to estimate the time-varying functional regression model.
-To address this issue, we use weighted least squares to estimate the time-varying functional regression model.
-To derive the weights, we first measure the Euclidean distance between the MFPC score corresponding to a unit $i$ and the centroid of the cluster of the active failure mode.
-We then use the reciprocal of that distance to define the corresponding weight for the unit, $w_{i}$, of that unit.
-If the unit’s failure is misclassified, $w_{i}$ would be large, and thus, unit $i$ is assigned a smaller weight in the regression problem.
+$c_{0}^{k^{*}}=\varphi_{0}^{k^{*}}+\int_{0}^{t^{*}}\sum_{h=1}^{H_{k^{*}}}c_{h}^{k^{*}}\boldsymbol{\varpi}_{h}^{k^{*}}(t)^{T}\boldsymbol{\mu}^{k^{*}}(t)\vskip-5.0pt$ (19)
+Weighted Version: The offline sensor selection algorithm is unlikely to cluster failures perfectly.
+Therefore, when fitting the regression model, it is possible that sample systems that failed due to other failure modes are included.
+These systems may present themselves as outliers, compromising RUL prediction accuracy.
+To address this issue, we use weighted least squares to fit the time-varying functional regression model.
+To derive the weights, we first measure the Euclidean distance between the MFPC score of each training system to the centroid of the cluster of the diagnosed failure mode.
+Then we calculate the reciprocal of that distance to define the corresponding weights for each system in the training set.
+If a training system’s failure mode is misclassified, its corresponding weight $w_{i}$ would be large, and thus, training system $i$ would be assigned a smaller weight in the regression problem.
 Next, we regress the standardized $\ln{TTF}$ on the MFPC-scores using a penalized weighted linear regression shown in Eq (20).
@@
 To ensure a regression coefficient represents the significance of their respective predictor to the prediction of the failure time, Eq (20) should be read such that $\zeta_{i,h}^{k^{*}}$ is standardized.
-$\min_{c_{0}^{k^{*}},c_{1}^{k^{*}},\ldots,c_{H_{k^{*}}}^{k^{*}}}\sum_{i\in\mathcal{I}_{t^{*}}^{k^{*}}}\left(\sqrt{w_{i}}y_{i}-\sqrt{w_{i}}c_{0}^{k^{*}}-\sum_{h=1}^{H_{k^{*}}}\sqrt{w_{i}}\zeta_{i,h}^{k^{*}}c_{h}^{k^{*}}\right)^{2}+\lambda\sum_{h=1}^{H_{k^{*}}}\lVert c_{h}^{k^{*}}\rVert$ (20)
+$\begin{split}\min_{c_{0}^{k^{*}},c_{1}^{k^{*}},\ldots,c_{H_{k^{*}}}^{k^{*}}}&\sum_{i\in\mathcal{I}_{t^{*}}^{k^{*}}}\left(\sqrt{w_{i}}y_{i}-\sqrt{w_{i}}c_{0}^{k^{*}}-\sum_{h=1}^{H_{k^{*}}}\sqrt{w_{i}}\zeta_{i,h}^{k^{*}}c_{h}^{k^{*}}\right)^{2}\\ &+\lambda\sum_{h=1}^{H_{k^{*}}}\lVert c_{h}^{k^{*}}\rVert\end{split}$ (20)
 5.3.2 RUL prediction
@@
 These scores are then used to predict the RUL at time $t^{*}$ using Eq (21).
-$\begin{split}{RUL}_{test}=\exp&\bigg{(}\Big{[}c_{0}^{k^{*}}+\sum_{k=1}^{H_{k^{*}}}\zeta_{test,h}^{k^{*}}c_{k}^{k^{*}}\Big{]}\times\\ &\Big{[}(1/(N-1))\sum_{i=1}^{N}\big{(}\ln{TTF}_{i}-(1/N)\sum_{i=1}^{N}\ln{TTF}_{i}\big{)}^{2}\Big{]}\\ &+(1/N)\sum_{i=1}^{N}\ln{TTF_{i}}\bigg{)}-t^{*}\end{split}$ (21)
+$\begin{split}{RUL}_{test}=\exp&\bigg(\Big[c_{0}^{k^{*}}+\sum_{k=1}^{H_{k^{*}}}\zeta_{test,h}^{k^{*}}c_{k}^{k^{*}}\Big]\times\Big[(1/(N-1))\\ &\sum_{i=1}^{N}\big(\ln{TTF}_{i}-(1/N)\sum_{i=1}^{N}\ln{TTF}_{i}\big)^{2}\Big]\\ &+(1/N)\sum_{i=1}^{N}\ln{TTF_{i}}\bigg)-t^{*}\end{split}$ (21)
 6 CASE STUDY 1: SIMULATED DATASET
@@
 Due to the scarcity of in-mission failure data and the infeasibility of performing full-scale failure testing in space, simulation-based validation is pursued in this case study.
-This study enables us to systematically evaluate our methodology’s ability to: (1) identify latent failure modes in unlabeled multivariate sensor data, (2) select informative sensors specific to each mode, and (3) accurately predict remaining useful life (RUL).
+This study enables us to systematically evaluate our methodology’s ability to: (1) identify latent failure modes in unlabeled multivariate sensor data, (2) select informative sensors specific to each mode, and (3) accurately predict RUL.
 To emulate the complex telemetry environment of deep-space habitats, we use a modified version of the simulation model from [18] to generate degradation signals for engineered systems monitored by 20 sensors.
@@
 $s_{i,p}^{k}(t)=-\frac{\theta_{i,p}^{k}}{\ln t}+\epsilon_{i,p}^{k}(t)$ (23)
-In this Eq, $\epsilon_{i,p}^{k}(t)\sim N\left(0,\left(\sigma_{k}^{p}\right)^{2}\right)$ is a white noise process parameterized by $\sigma_{k}^{p}=\mu_{k}/SNR_{p}^{k}$, where $SNR_{p}^{k}$ is the signal-to-noise for sensor $p$ and failure mode $k$.
+In this equation, $\epsilon_{i,p}^{k}(t)\sim N\left(0,\left(\sigma_{k}^{p}\right)^{2}\right)$ is a white noise process parameterized by $\sigma_{k}^{p}=\mu_{k}/SNR_{p}^{k}$, where $SNR_{p}^{k}$ is the SNR for sensor $p$ and failure mode $k$.
 To account for the correlation between the signal trends of the informative sensors and the underlying degradation process, we generate $\theta_{i,p}^{k}$ from the following conditional distribution $\theta_{i,p}^{k}\mid\theta_{i}^{k}\sim N\left(\mu_{k}\left(1-\sqrt{1-\rho_{p}^{k}}\right),(0.1)^{2}\right)$ such that the correlation, $\rho_{p}^{k}$, is uniformly sampled from the interval $[0.80,0.99]$.
@@
 Finally, we randomly assign a sign to $\theta_{i,p}^{k}$ to allow for signals that increase or decrease in response to degradation.
-In this simulation, we seek to test the robustness of our methodology to various levels of signal noise.
-Therefore, we generate three datasets, each with a unique interval from which to sample $SNR_{p}^{k}$.
-For $p\in\mathcal{J}_{k}$, $SNR_{p}^{k}$ is sampled uniformly from $[2,5]$, $[5,8]$, and $[8,11]$ for datasets 1, 2, and 3, respectively.
-For $p\in\mathcal{N}_{k}$, $SNR_{p}^{k}$ is sampled uniformly from $[1,3]$ for all three datasets.
-Figure 3 displays samples of our sensor signals in a 3$\times$4 grid.
-Each row displays signals from four sensors (1, 3, 5, 19) for a particular signal-to-noise ratio.
-Sensor 1 is not informative for either failure mode so no discernible trend is present.
-Sensors 3 and 5 are informative for FM 2 and FM 1, respectively.
-For these sensors, one signal trends downward due to degradation while the other signal appears to resemble random noise.
-Finally, for sensor 19, the signal decreases in response to both failure modes.
-Figure 3: Case study 1 data: Signals in sensors 1, 3, 5, 19 (columns 1, 2, 3, 4 respectively) for FM 1 and 2 under different SNRs – [2, 5], [5, 8], and [8, 11] (rows 1, 2, 3 respectively)
+[…]
 6.2 Failure Mode Labeling and Sensor Selection
@@
 The results of this performance evaluation are shown in Table 3.
+For all scenarios, the initial failure mode labels were determined randomly.
+Table II shows that for the lowest and the highest SNRs, the clustering algorithm was capable of achieving at worst 78.8% accuracy for a particular failure mode.
+The exception is the dataset [5,8], where the algorithm has low accuracy at clustering.
+This is due to the tuning parameter combination resulting in minimal coefficient shrinkage that enabled the nonsignificant sensors to influence cluster assignments.
+A slightly larger $\lambda$ value would decrease this error.
+For all three models, several sensors that are deemed ”noninformative” were selected.
+This can be attributed to some of these sensors having high values of $\rho_{p}^{k}$ relative to the specified range of [0.1,0.6].
+However, simply reporting the sensors selected does not account for how informative they are.
+Table 4 reports the four sensors with the highest $l_{2}$-norm of the regression coefficients associated with that sensor.
+Except for the [5,8] dataset, at least three of the four sensors with the largest $l_{2}$-norm for each failure mode are elements of the set of informative sensors.
+This indicates that the MGR-ASGL model results in the informative sensors having larger regression coefficients (in magnitude) than the noninformative sensors.
+Given the selected sensors, we want to test the robustness of the online portion of the methodology to predict remaining life.
+6.3 RUL Prediction
+Each of the datasets consist of 80 test systems (40 from each failure mode).
+For each test system, we observe degradation up until the following life percentiles (in retrospect): 10%, 20%, 30%, 40%, 50%, 60%, 70%, 80%, and 90%.
+We smooth the training and the test signals using a bandwidth parameter of 0.5.
+Next, we perform MFPCA selecting the number using only the training systems that survived up to the observation time $t^{*}$ corresponding to each life percentile.
+For a given test signal, we calculate its MFPC scores.
+The active failure mode is diagnosed by being assigned to the most frequent failure mode present in its closest $0.1N_{t^{*}}$ neighbors.
+Once the active failure mode is identified, the MFPC scores of the test signals are updated and recomputed using the subset of training signals associated with the active failure mode.
+Next, we fit the weighted regression model and predict the RUL using the scores of the test signal.
+Prediction accurately is calculated retrospectively.
+We compute the “Estimated Life” as the current operating time plus the predicted RUL.
+A relative error is then computed using Eq (24).
+ $\text{ Relative Error }=\frac{\mid\text{Estimated Life}-\text{Actual Life}\mid}{\text{Actual Life}}\times 100\%$ (24)
+Figure 3: Case study 1 data: Signals in sensors 1, 3, 5, 19 (columns 1, 2, 3, 4 respectively) for FM 1 and 2 under different SNRs – [2, 5], [5, 8], and [8, 11] (rows 1, 2, 3 respectively)
+The performance of our methodology on predicting RUL is shown in Figure 4.
+In this figure, the relative errors have a generally decreasing trend for all datasets, with the [8,11] dataset tending to have smaller relative errors on average.
+Furthermore, the variability appears to be much larger when making predictions early in the system’s life rather than later.
+Interestingly, the relative error increased for the [8,11] dataset when making predictions at 90% of the system’s life.
+This increase may be attributed to the increased influence of incorrectly clustered training samples affecting the RUL prediction.
+Since this dataset has less noise, the influence of clustering may be more pronounced compared to the noisier datasets, especially if training samples have been removed for failing before 90% of the monitored system’s lifetime.
+Figure 4: Prediction error (in $\%$) for all test systems of case study 1 under three different SNRs i.e., [2, 5], [5, 8], and [8, 11]
+Figure 5: The system diagram used in Case Study 2 adapted from [53]
+Figure 6: Raw sensor data used for offline step in case study 2. The data is inherently unlabeled i.e., failure modes are not known.
+Figure 7: Average Prediction error ($\%$) in case study 2 with comparison to [38], and [37]
+Figure 8: Prediction error ($\%$) for different RUL intervals in case study 2 with comparison to [37]
 Table 2: Tuning parameter selection for case study 1
@@
 2 0.788 1:3,5:20
-For all scenarios, the initial failure mode labels were determined randomly.
-Table II shows that for the lowest and the highest signal-to-noise ratios, the clustering algorithm was capable of achieving at worst 78.8% accuracy for a particular failure mode.
-The exception is the dataset [5,8], where the algorithm has low accuracy at clustering.
-This is due to the tuning parameter combination resulting in minimal coefficient shrinkage that enabled the nonsignificant sensors to influence cluster assignments.
-A slightly larger $\lambda$ value would decrease this error.
-For all three models, several sensors that are deemed ”noninformative” were selected.
-This can be attributed to some of these sensors having high values of $\rho_{p}^{k}$ relative to the specified range of [0.1,0.6].
-However, simply reporting the sensors selected does not account for how informative they are.
-Table 4 reports the four sensors with the highest $l_{2}$-norm of the regression coefficients associated with that sensor.
-Except for the [5,8] dataset, at least three of the four sensors with the largest $l_{2}$-norm for each failure mode are elements of the set of informative sensors.
-This indicates that the MGR-ASGL model results in the informative sensors having larger regression coefficients (in magnitude) than the noninformative sensors.
-Given the selected sensors, we want to test the robustness of the online portion of the methodology to predict remaining life.
 Table 4: $l_{2}$-norms of four most informative sensors for FM1 & FM2
@@
 18 0.394 17 0.428
-6.3 RUL Prediction
-Each of the datasets consist of 80 test systems (40 from each failure mode).
-For each test system, we observe degradation up until the following life percentiles (in retrospect): 10%, 20%, 30%, 40%, 50%, 60%, 70%, 80%, and 90%.
-We smooth the training and the test signals using a bandwidth parameter of 0.5.
-Next, we perform MFPCA selecting the number using only the training systems that survived up to the observation time $t^{*}$ corresponding to each life percentile.
-For a given test signal, we calculate its MFPC scores.
-The active failure mode is diagnosed by being assigned to the most frequent failure mode present in its closest $0.1N_{t^{*}}$ neighbors.
-Once the active failure mode is identified, the MFPC scores of the test signals are updated and recomputed using the subset of training signals associated with the active failure mode.
-Next, we fit the weighted regression model and predict the RUL using the scores of the test signal.
-Prediction accurately is calculated retrospectively.
-We compute the “Estimated Life” as the current operating time plus the predicted RUL.
-A relative error is then computed using Eq (24).
-[…]
-[…]
-The performance of our methodology on predicting RUL is shown in Figure 4.
-In this figure, the relative errors have a generally decreasing trend for all datasets, with the [8,11] dataset tending to have smaller relative errors on average.
-Furthermore, the variability appears to be much larger when making predictions early in the system’s life rather than later.
-Interestingly, the relative error increased for the [8,11] dataset when making predictions at 90% of the system’s life.
-This increase may be attributed to the increased influence of incorrectly clustered training samples affecting the RUL prediction.
-Since this dataset has less noise, the influence of clustering may be more pronounced compared to the noisier datasets, especially if training samples have been removed for failing before 90% of the monitored system’s lifetime.
 7 CASE STUDY 2: NASA TURBOFAN ENGINE DATASET
-Figure 5: The system diagram used in Case Study 2 adapted from [55]
 To further validate our methodology in a realistic operational context, we evaluate its performance on a benchmark dataset that closely mirrors subsystem degradation in deep-space habitats.
-Specifically, we apply our approach to the NASA’s Commercial Modular Aero-Propulsion System Simulation (C-MAPSS) dataset [55].
+Specifically, we apply our approach to NASA’s Commercial Modular Aero-Propulsion System Simulation (C-MAPSS) dataset [53].
 C-MAPSS simulates degradation in the components of turbofan engines that share telemetry and fault characteristics with propulsion and environmental control systems of deep space habitats.
@@
 While the dataset provides the potential causes of failure, it does not provide the failure labels.
-Furthermore, the test set consists of systems whose degradation is only partially observed along with their corresponding failure times.
+Furthermore, the test set consists of systems for which degradation is only partially observed, together with their corresponding failure times.
 Therefore, our goal is to use our methodology to predict the remaining life of each testing unit following the final cycle at which their degradation was observed.
@@
 The sensors selected are those with positive $l_{2}$-norms.
+In addition to performing sensor selection, the MGR-ASGL model provides cluster labels for each observed failure.
+These labels are utilized for the online prediction portion of the methodology.
+For the online prediction step, we first smooth the degradation signals using 5-fold cross-validation to select the bandwidth parameter.
+After smoothing, we perform MFPCA on signals from the sensors selected across both failure modes to build our classifier.
+Then, we classify the failure mode based on the $0.1N_{t^{*}}$ nearest neighbors.
+Given this classification, we perform MFPCA on signals from the sensors selected for the classified failure mode.
+Finally, we perform Weighted Time-varying Functional Regression to regress $lnTTF$ on the MFPC-scores using Leave-one-out cross-validation to select the tuning parameter.
+7.1 Comparison to Baselines
+To analyze our methodology, we compare its performance to that of [37] Chehade et al.
+(2018), whose methodology assumes knowledge of the failure modes but does not perform sensor selection (outside of removing the blank sensors), and [38] Li et al.
+(2023), whose methodology assumes knowledge of a fraction of the training set and does perform sensor selection.
+Figure 7 displays the relative prediction error averaged over all test systems with RUL at most 100, 80, 60, 40, and 20 cycles.
+From this figure, our methodology becomes more accurate as the RUL decreases.
+Furthermore, our methodology is more accurate than both when making longer predictions.
+For test units whose RUL is at most 20, our accuracy is between that of [37] and [38].
+Only at RUL at most 40, does our methodology perform worse than both.
+To further analyze our methodology, we refer to Figure 8, which displays boxplots showing the distribution of the relative prediction error for test units with remaining life in the following intervals: [0,25],[26,50],[51,75],[76,100],[101,125],[126,150][0,25],[26,50],[51,75],[76,100],[101,125],[126,150].
+The results indicate that our methodology achieves a lower median relative error than [37] in all intervals while maintaining comparable variability.
+However, the variability for systems with RUL between 25 and 50 is higher for our methodology than for [37].
+This indicates a targeted area for improvement of our methodology.
+However, the general trend is that sensor selection resulted in a more accurate RUL prediction than no sensor selection.
+Furthermore, we were capable of achieving more accurate RUL prediction than both baselines despite not having information about the actual failure mode.
 Table 5: Sensors used in case study 2 (C-MAPPS-3 [53])
@@
 21 W32 LPT coolant bleed
-In addition to performing sensor selection, the MGR-ASGL model provides cluster labels for each observed failure.
-These labels are utilized for the online prediction portion of the methodology.
-For the online prediction step, we first smooth the degradation signals using 5-fold cross-validation to select the bandwidth parameter.
-After smoothing, we perform MFPCA on signals from the sensors selected across both failure modes to build our classifier.
-Then, we classify the failure mode based on the $0.1N_{t^{*}}$ nearest neighbors.
-Given this classification, we perform MFPCA on signals from the sensors selected for the classified failure mode.
-Finally, we perform Weighted Time-varying Functional Regression to regress $lnTTF$ on the MFPC-scores using Leave-one-out cross-validation to select the tuning parameter.
-Figure 6: Raw sensor data used for offline step in case study 2. The data is inherently unlabeled i.e., failure modes are not known.
-7.1 Comparison to Baselines
-To analyze our methodology, we compare its performance to that of [39] Chehade et al.
-(2018), whose methodology assumes knowledge of the failure modes but does not perform sensor selection (outside of removing the blank sensors), and [40] Li et al.
-(2023), whose methodology assumes knowledge of a fraction of the training set and does perform sensor selection.
-Figure 7 displays the relative prediction error averaged over all test systems with RUL at most 100, 80, 60, 40, and 20 cycles.
-From this figure, our methodology becomes more accurate as the RUL decreases.
-Furthermore, our methodology is more accurate than both when making longer predictions.
-For test units whose RUL is at most 20, our accuracy is between that of [39] and [40].
-Only at RUL at most 40, does our methodology perform worse than both.
-To further analyze our methodology, we refer to Figure 8, which displays boxplots showing the distribution of the relative prediction error for test units with remaining life in the following intervals: $[0,25],[26,50],[51,75],[76,100],[101,125],[126,150]$.
-The results indicate that our methodology achieves a lower median relative error than [39] in all intervals while maintaining comparable variability.
-However, the variability for systems with RUL between 25 and 50 is higher for our methodology than for [39].
-This indicates a targeted area for improvement of our methodology.
-However, the general trend is that sensor selection resulted in a more accurate RUL prediction than no sensor selection.
-Furthermore, we were capable of achieving more accurate RUL prediction than both methods from earlier in the systems’ lives despite not having information about the actual failure mode.
 Table 6: Sensor Selection for Case Study 2
@@
 11 0 0 21 0.116 0.318
-[…]
-[…]
 8 CONCLUSION
@@
 Future work will focus on extending the methodology to nonlinear models to improve prediction accuracy and capture complex degradation trends in autonomous mission scenarios.
-Acknowledgments
-This effort is supported by NASA under grant number 80NSSC19K1052 as part of the NASA Space Technology Research Institute (STRI) Habitats Optimized for Missions of Exploration (HOME) ‘SmartHab’ Project.
-Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Aeronautics and Space Administration.
 Appendix A Log Likelihood
@@
 This function is called as the ”incomplete-data log-likelihood” (IDLL), since we do not know the cause of the observed failure in the data.
-$\ell(\boldsymbol{\Theta};\boldsymbol{Y})=\sum_{i=1}^{N}\ln\sum_{k=1}^{K}\pi_{k}\times\ \frac{\rho_{k}}{\sqrt{2\pi}}\exp\left(-\frac{1}{2}\left(\rho_{k}y_{i}-\psi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\psi}_{p,k}\right)^{2}\right)$ (25)
+$\begin{split}\ell(\boldsymbol{\Theta};\boldsymbol{Y})=\sum_{i=1}^{N}\ln\sum_{k=1}^{K}\pi_{k}\frac{\rho_{k}}{\sqrt{2\pi}}&\exp(-\frac{1}{2}(\rho_{k}y_{i}\\ &-\psi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\psi}_{p,k})^{2})\end{split}$ (25)
 Definition 2.
@@
 Let $g$ be an arbitrary distribution on $\boldsymbol{Z}_{i}[k]$. Then we can re-write equation (25) as follows:
-$\ell(\boldsymbol{\Theta};\boldsymbol{Y})=\sum_{i=1}^{N}\ln\left(\sum_{k=1}^{K}\frac{g\left(\boldsymbol{Z}_{i}[k]\right)p\left(y_{i},\boldsymbol{Z}_{i}[k];\boldsymbol{\Theta}\right)}{g\left(\boldsymbol{Z}_{i}[k]\right)}\right)\vspace*{-1em}$
+$\ell(\boldsymbol{\Theta};\boldsymbol{Y})=\sum_{i=1}^{N}\ln\left(\sum_{k=1}^{K}\frac{g\left(\boldsymbol{Z}_{i}[k]\right)p\left(y_{i},\boldsymbol{Z}_{i}[k];\boldsymbol{\Theta}\right)}{g\left(\boldsymbol{Z}_{i}[k]\right)}\right)\vskip-10.00002pt$
 where,
@@
 Since $\ln(.)$ is concave, from Jensen’s inequality we can write that,
-$\ell(\boldsymbol{\Theta};\boldsymbol{Y})\geq\sum_{i=1}^{N}\sum_{k=1}^{K}g\left(\boldsymbol{Z}_{i}[k]\right)\ln{\frac{p\left(y_{i},\boldsymbol{Z}_{i}[k];\boldsymbol{\Theta}\right)}{g\left(\boldsymbol{Z}_{i}[k]\right)}}\vspace*{-1em}$
+$\ell(\boldsymbol{\Theta};\boldsymbol{Y})\geq\sum_{i=1}^{N}\sum_{k=1}^{K}g\left(\boldsymbol{Z}_{i}[k]\right)\ln{\frac{p\left(y_{i},\boldsymbol{Z}_{i}[k];\boldsymbol{\Theta}\right)}{g\left(\boldsymbol{Z}_{i}[k]\right)}}\vskip-10.00002pt$
 and thus,
-$\ell(\boldsymbol{\Theta};\boldsymbol{Y})\geq\mathbb{E}_{\boldsymbol{g}}\left[\ell_{C}(\boldsymbol{\Theta};\boldsymbol{Y},\mathbb{Z})\right]-\sum_{i=1}^{N}\sum_{k=1}^{K}g\left(\boldsymbol{Z}_{i}[k]\right)\ln{g\left(\boldsymbol{Z}_{i}[k]\right)}\vspace*{-1em}$
+$\ell(\boldsymbol{\Theta};\boldsymbol{Y})\geq\mathbb{E}_{\boldsymbol{g}}\left[\ell_{C}(\boldsymbol{\Theta};\boldsymbol{Y},\mathbb{Z})\right]-\sum_{i=1}^{N}\sum_{k=1}^{K}g\left(\boldsymbol{Z}_{i}[k]\right)\ln{g\left(\boldsymbol{Z}_{i}[k]\right)}\vskip-10.00002pt$
 The second term on the right side of the inequality is the negative entropy of the $g$ distribution. Since it is positive, the inequality holds. ∎